### v1

# PACEE: Supporting Children’s Personal Emotion Education through Parent-AI Collaboration

Abstract.
Emotion education is a crucial lesson for children aged 3 to 6. However, existing technologies primarily focus on promoting emotion education from the child’s perspective, often neglecting the central role of parents in guiding early childhood emotion development. In this work, we conducted co-design sessions with five experienced kindergarten teachers and five parents to identify parental challenges and the roles that AI can play in family emotion education. Guided by these insights, we developed PACEE, an assistant for supporting parent-AI collaborative emotion education. PACEE enables parents to engage in emotional dialogues about common scenarios, with multiple forms of support provided by generative AI. It combines insights from parents and AI to model children’s emotional states and collaboratively delivers personalized, parent-mediated guidance. In a user study involving 16 families, we found that PACEE significantly enhances parent-child engagement, encourages more in-depth emotional communication, and improves the parental experience. Our findings advance emotion coaching theory in both family settings and LLM-assisted contexts, offering valuable insights for designing AI-supported, parent-centered family education systems.
Parent-child Interaction, Emotion Education, Large Language Models
journalyear: 2025copyright: ccconference: ; ; $c$s: Human-centered computing Collaborative interaction
Figure 1. PACEE enables parents to engage in emotion education with their children through interactive adventure game. PACEE consists three key modules. (1) Parent-Child Adventure Game. In each adventure game, PACEE provides parents with real-time guidance on emotional dialogues, addressing the challenge of limited parental skills ($C_{1}$). (2) Parental Feedback Report. It provides parents with review and suggestions after each adventure game, addressing the challenge of insufficient feedback ($C_{2}$). (3) Child Emotional Modeling. It creates an emotional profile for each child by combining two sources: long-term understanding provided by parents and AI’s analysis of parent-child dialogue data. The emotional profile is then used to offer personalized coaching support to parents. It also enhances parents’ understanding of their child and effectively addressing the challenge of lack of understanding of children’s emotional states ($C_{3}$). :
1. INTRODUCTION
Emotion education plays a critical role in the developmental years of children aged 3 to 6. According to developmental psychology, during this period, children begin to understand various emotions — such as anger, fear, and sadness — and learn to express and regulate these feelings (bartlett1982emotional; saarni1999development; wulandari2021development).
Recent studies have explored AI-based systems that help children recognize, express, and manage their emotions. However, these systems mainly target children aged 8 and above, as they rely on independent conversations with LLM-based agents — for example, expressing emotions through dialogue (seo2024chacha; tang2024emoeden; shen2025easel). Such designs are not well-suited for younger children aged 3 to 6, since at this stage, parents play a foundational role in shaping children’s emotional development by guiding their early emotional learning at home (fadlillah2022analysis; alegre2011parenting). Unlike older children, those aged 3 to 6 require parental support to engage with technology (toran2024parent; morawska2023managing) and rely on their parents for emotional security and co-regulation (corwin2012secure; brock2016interparental).
However, providing effective emotion education can be challenging for parents, as it requires knowledge of child psychology and strong communication skills — expertise most parents lack (sanders1999triple; saarni1999development; cole1994development). In practice, parents often encounter difficulties like misreading children’s ambiguous emotional signals (castro2015parents), or failing to support emotional regulation through dialogue (meng2020effects).
To gain a more systematic understanding of these difficulties, we conducted interviews with five kindergarten teachers and five parents and identified three key challenges: ($C_{1}$) limited parental expertise in child psychology and communication techniques; ($C_{2}$) insufficient feedback on the effectiveness of emotion education for children; ($C_{3}$) a lack of comprehensive understanding of children’s emotional states. Furthermore, both parents and teachers expressed a clear need for AI-driven tools to facilitate emotion education at home.
We developed PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education for children aged 3 to 6. PACEE enables parents to engage in emotion education with their children through interactive adventure games. PACEE consists of three key modules, each addressing one of the aforementioned challenges: (1) Parent-Child Adventure Game. In each adventure game, parents and their children engage in five-step emotional dialogues based on Gottman’s Emotion Coaching Theory (gable1999heart). The AI system provides parents with real-time guidance, such as actionable suggestions and progress tracking, thus solving $C_{1}$. (2) Parental Feedback Report. This module provides parents with feedback after each adventure game. The feedback includes a comprehensive review of past parent-child conversations and actionable suggestions for future interactions, effectively addressing $C_{2}$. (3) Child Emotional Modeling. This module creates an emotional profile for each child by combining two views: long-term understanding provided by parents and AI’s analysis of parent-child dialogue data in the adventure game. The emotional profile is then used to offer tailored guidance to parents during the adventure game. After each adventure game, it also provides parents with a comparison between their own understanding of their child’s emotional states and AI observations, thereby improving parents’ understanding of their child and effectively addressing $C_{3}$.
To evaluate PACEE’s effectiveness in supporting family emotion education, we conducted a user study with 16 pairs of parent and child. Compared to a baseline (LLM-assisted) method, PACEE significantly fosters meaningful parent-child interactions and in-depth emotion education. Furthermore, with PACEE, children show greater engagement and richer emotional expression during conversations. Meanwhile, parents (1) improve their expertise and confidence in communication, (2) adopt more effective approaches in emotional education, and (3) gain a better understanding of their children’s emotional states.
The contributions of our research are as follows:
  • Through a formative study, we identified three key challenges in family emotion education, AI’s role in supporting this process, and principles for AI-assisted emotion education from experts’ perspectives.
  • We present PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education. PACEE collaboratively builds children’s emotional profiles with parents and provides personalized, parent-mediated guidance.
  • Through a user study with 16 families, we validated PACEE’s effectiveness in enhancing family emotion education and advanced emotion coaching theory in both family and LLM-assisted contexts, offering valuable insights for designing AI-supported, parent-centered family education systems.
2. RELATED WORK
2.1. Emotion Coaching Techniques and Theories
Emotion coaching is a supportive parenting technique that helps children understand and regulate emotions and has been widely studied in parenting and child development (fu2022social; havighurst2021tuning). A key framework widely applied in HCI (havighurst2021tuning) is the “Five Steps of Emotion Coaching” framework by John Gottman (gottman2005five), which includes (1) attending to children’s emotional signals, (2) viewing emotional moments as opportunities for connection and learning, (3) helping children identify and articulate emotions, (4) expressing empathy and validating emotions, and (5) providing guidance and boundaries while collaboratively exploring solutions. This model is often complemented by principles from positive psychology (baker2017positive; fan2018emostory), such as the PERMA model (seligman2010flourish). PERMA model outlines five elements of well-being (Positive Emotion, Engagement, Relationships, Meaning, and Achievement) and has been applied in HCI areas such as well-being-centered design (kovich2023application), persuasive technologies (nebrida2018m), and digital mental health (chisale2022perma).
Current HCI implementations of emotion coaching primarily target older children or specific clinical populations. Research has explored using AI to enhance emotional expression and feedback for school-aged children and adolescents (seo2024chacha; kim2018can; ibrahim2025designing). Meanwhile, other research focuses on providing specialized learning support for children with clinical needs, such as autism (tang2024emoeden; ghiotti2023prototyping; sharma2016promoting) or trauma (grassl2024coding). However, insufficient attention is given to children aged 3 to 6, highlighting a critical gap in HCI-driven emotional education.
Compared to older children, 3- to 6-year-olds experience emotions more immediately, intensely, and situationally while having limited verbal and self-regulation skills (sun2024exploring). These traits make parental guidance central for fostering emotional competence (gottman1996parental; hightower2019exploring). However, many parents face barriers such as time constraints and a lack of knowledge in addressing their children’s emotional needs (ahmadpour2023understanding; quan2025parents; BROWN2022126). Existing tools frequently overlook these developmental traits, assuming children can engage with AI independently and failing to empower parents in their crucial coaching role (sharma2025robot).
To address this gap, we propose a parent-centered, theory-grounded approach designed to empower parents as the primary emotion coaches for their 3- to 6-year-old children.
2.2. Parent-AI Collaborative Coaching System
The introduction of artificial intelligence (AI) has opened new possibilities for children’s emotional development in modern family education. Based on how AI participates in parent-child interactions, existing research can be categorized into three types: AI-centered, parent-centered, and parent-AI collaborative systems.
AI-centered systems position AI as a “teaching tool” that interacts directly with children, offering guidance to help children (wang2024chaitok; wang2022don; karunaratna2024evaluation). For example, systems may generate personalized learning content (tang2024emoeden), support children in learning math language via story-based interaction (xu2023mathkingdom), co-create stories to boost vocabulary and feedback quality (zhang2024mathemyths), or generate context-based prompts to aid language expression (seo2024chacha). These systems emphasize direct AI-child interaction.
Parent-centered systems emphasize parents as primary educators, with AI as a supportive tool that does not interact directly with the child. These studies often use interviews or co-design methods to explore parents’ needs and challenges (dumaru2023after; cingel2017parents; nikkhah2021helping; hiniker2016not; mazmanian2017okay; wang2021protection). For example, research shows how parents co-use digital media with children (yu2024parent) and welcome daily prompts for reflection (ibrahim2025designing). However, few systems support real-world implementation, especially regarding sustained engagement and feedback mechanisms.
Parent-AI collaborative systems integrate the strengths of both parties through a division of labor. Parents can act as information providers, content configurators, facilitators, and evaluators (slovak2016scaffolding; chen2020new; yarosh2013almost). Meanwhile, AI is used to facilitate communication and storytelling (choi2024aacesstalk; zhang2022storybuddy), structure collaborative tasks and routines (jonas2023supporting; xiao2012supporting), and integrate smart devices into daily life (theofanopoulou2022exploring; chan2017wakey). These systems establish a three-way interaction model among the parent, child, and AI.
Despite these developments, most systems still focus on school-aged children and involve direct AI-child interaction, overlooking the parents’ central role in early childhood. To address this, we propose a parent-centered, AI-supported approach, which aims to enhance, rather than replace, parental engagement.
2.3. Modeling Children’s Emotional States
Personalized instruction is crucial in modern education and children’s emotional development (bernacki2021systematic; dewey2024democracy). Effective personalization requires modeling a child’s emotional state to enable targeted interventions. We examine the challenges of such modeling through two aspects: content design and assessment methods.
Regarding content design, a classic assessment for children’s emotional development is the Test of Emotion Comprehension (TEC) (pons2004emotion; albanese2006children). TEC identifies nine key elements for evaluating children’s emotional competence, which have also been widely acknowledged in prior research, including: emotion recognition (bullock1985further), emotional expression (barden1980children), emotion external cause understanding (harris1989young), emotion–belief link (fonagy1997relationship), emotion reminder effect (harris1983children), emotional masking (gardner1988japanese), mixed emotional understanding (arsenio1995children), emotional morality (harter1989developmental), and emotion regulation (altshuler1989developmental).
Regarding assessment methods, traditional approaches like parent-reported or teacher-reported questionnaires offer high ecological validity but suffer from subjectivity and inefficiency (denham1997parental; rothbart2004temperament). In contrast, self-report instruments like TEIQue-CF (mavroveli2008investigation) provide direct insight into children’s self-perceived emotional competence. However, for young children aged 3 to 6, limited cognitive and expressive abilities (wellman2004scaling) can lead to unreliable results. More recently, emerging AI systems using chatbots or games show promise for automated observation (santos2020therapist; nicolaidou2022gamified). However, these systems often struggle to interpret subtle and developmentally specific cues, especially when the data types available are limited due to privacy concerns (calvo2010affect).
To address the limitations of existing assessment methods, we propose a human-AI collaborative framework that integrates the real-world understanding of parents with the multimodal analysis of AI, enabling more comprehensive emotional modeling.
3. STUDY PROCEDURE
Figure 2. Our 3-phased study procedure.
This study was approved by the Institutional Review Board, and we obtained consent from all participating parents and children. We conducted three study phases to explore, develop, and evaluate an AI assistant for supporting family emotion education (Figure 2).
Phase 1: Formative study. To gain deeper insights about potential needs and challenges in family emotion education, we conducted three formative studies, including: (A) an interview session with 5 experienced kindergarten teachers, (B) an interview session with 5 parents of children aged 3-6, and (C) an observational experiment focusing on emotion education involving 5 children. Thus, we identified three key challenges in family emotion education ($C_{1}$-$C_{3}$).
Phase 2: Iterative design and implementation. To mitigate these challenges, we developed three core modules of PACEE: Parent-Child Adventure Game ($C_{1}$), Parental Feedback Report ($C_{2}$), and Child Emotional Modeling ($C_{3}$). PACEE follows Gottman’s Five Steps of Emotion Coaching framework, the PERMA model, and the design details advised from formative study interviews. Section 5 describes the design and implementation details.
Phase 3: Evaluation study. To evaluate the system, we conducted a within-subject study with 16 pairs of parents and children (3-6 years). We designed a Baseline condition with a generic LLM chatbot. Section 6 explains our deployment setup, and Section 7 reports the findings.
4. FORMATIVE STUDY
To gain a deeper understanding of family emotion education and explore how AI can assist in this process, we conducted three formative studies: an interview session with 5 experienced kindergarten teachers (denoted as K1-K5), an interview session with 5 parents of children aged 3-6 (denoted as G1-G5), and an observational experiment focusing on emotion education involving 5 children (denoted as O1-O5).
4.1. Procedure and Analysis
4.1.1. Need-finding Interview with Experienced Kindergarten Teachers
We interviewed 5 experienced kindergarten teachers (3 female, 2 male; $M=9.40$ years of experience, $SD=6.34$) recruited via social media, each receiving a \$10 stipend. Each interview session lasted about one hour and was conducted through video conferencing, moderated by two researchers.
First, we interviewed the teachers on several key topics, including common negative emotions in children aged 3-6, triggers, and teachers’ strategies. Next, we explored the potential role of AI in supporting emotion education. Specifically, we introduced the design objectives and scenarios of parents providing regular emotional guidance to their children at home while the AI offers additional support throughout the process. Further, we co-designed an AI-assisted system based on participants’ visions and expertise, using conceptual design images to illustrate our design objectives (see figure 3(b)).
4.1.2. Need-finding Interview with Parents of Children Aged 3-6
We also interviewed 5 parents (4 mothers, 1 father; $M=41.0$ years old, $SD=3.16$) of children aged 3-6, recruited via social media, with \$10 stipends. Each hour-long video session was moderated by two researchers.
First, we interviewed the parents about their observations of children’s emotional behaviors, challenges in guiding emotions, and understanding of emotional regulation. Next, our design scenarios explored the parents’ perspectives on AI-assisted tools, preferred features, and concerns.
Together with the kindergarten teachers’ sessions, these interview sessions were audio-recorded, transcribed, and analyzed using a coding process by three HCI experts.
4.1.3. Observational Experiment on Children’s Emotion Education
To additionally understand the context of how parents and children communicate and interact during emotion education conversations, we observed emotion education interactions in 5 families recruited via social media, with \$10 compensation. The participants included four mothers, one father ($M=38.2$ years old, $SD=3.49$), and their children (4 boys, 1 girl; $M=4.8$ years old, $SD=1.17$). Each family received \$10 or an equivalent gift as compensation.
Conducted in a lab over 30 minutes, parents guided their children based on a predefined script adapted from Gottman’s Five Steps of Emotion Coaching theory (gable1999heart). Parents followed a script and selected one of three negative emotion scenarios provided. Sessions were video-recorded, analyzed, and annotated by three HCI experts. Finally, we conducted a brief interview to ask parents about their coaching experience and difficulties.
(a) Observational experiment settings
(b) A conceptual design image in the interview session
Figure 3. Photos taken in the formative studies, including: (a) the setup of an observational experiment, where parents guide their children following a predefined script, and (b) a conceptual design image presented during the interview session to illustrate our design objectives.
4.2. Findings
4.2.1. Key Challenges in Family Emotion Education
Through interviews with 5 kindergarten teachers (K1-K5) and 5 parents (G1-G5), we identified common scenarios where children aged 3-6 struggle with negative emotions. These emotions are often difficult to manage and require repeated guidance for early childhood. The findings serve as initial data for our system’s scenario database (see Table 4 in Appendix B).
Additionally, parental interviews revealed current practices and challenges in emotional communication and guidance. Combined with insights from observational experiments, we categorized the key challenges in family emotion education into three dimensions:
$C_{1}$
. Limited parental expertise. All five interviewed parents (5/5) acknowledged lacking professional-level skills in child emotion education, which manifests in three key aspects. First, parents demonstrated limited psychological and educational knowledge and struggled to adapt general theoretical guidelines to specific situations. For example, G2 remarked: “I feel like I don’t have access to expert advice or reliable sources, and the examples in books don’t always apply to my kid’s situation.” We also observed similar patterns in the observational study, as parents also reported that they “don’t often apply every step before”(O3) and “don’t know what to say” about certain steps (O1). Second, parents reported insufficient communication techniques, emphasizing the need for “concrete, easy-to-implement language and techniques” (G1, G2, G5) in daily practice. Third, parents experience occasional impatience during emotional communication with their children. G3 observed that “Sometimes my husband gets impatient when he and our child disagree, and it makes our child hesitant to open up and share feelings with him”, illustrating how parental emotional instability can negatively impact communication effectiveness.
$C_{2}$
. Insufficient feedback. Parents frequently lack objective, professional feedback when guiding children’s emotional development. As G3 noted, “When he has strong emotions, I explain and comfort him in the moment, but I can’t tell if he really understood or whether my approach was effective”. In the observational experiment, parents reported similar uncertainties (O4: “I just followed the guidelines step-by-step, but I’m not sure each step worked. Next time this happens, he might forget what we discussed today.”).
$C_{3}$
. A lack of understanding of children’s emotional states. Most parents reported difficulties in accurately recognizing their children’s emotional states. This stems partially from limited awareness of school experiences, as G4 noted “not knowing what happens at kindergarten since children don’t always share things at school.” Parents may struggle to identify the root causes of emotions (G1: “Once he cried because he was just too tired, but I initially thought it was because of something that happened at kindergarten.”). Additionally, parents acknowledged their subjective biases when assessing their children (G5: “We naturally think our child is the best, so this bias colors our evaluations. We need something objective to show how our child compares with peers.”).
4.2.2. Roles and Design Principles of AI in Family Emotion Education
During the interview sessions, parents expressed their vision for AI’s potential roles in emotional guidance, while experts highlighted the cognitive and emotional characteristics of young children that should guide system design.
All five participating parents expressed a preference for maintaining their primary role in emotional guidance while leveraging AI to supplement their limitations. They also envisioned specific roles for AI, including serving as a dialogue facilitator, an expert resource, and a generative tool. 4/5 parents anticipated AI’s involvement as a dialogue facilitator. For instance, G3 suggested that AI could dynamically adjust its level of participation based on the parent’s engagement: “When parents take the lead, AI should speak less; when parents struggle, AI can step in to enrich the conversation and reduce communication barriers.” G2 further highlighted AI’s potential to assist less experienced parents: “If parents are unsure how to respond, AI could intervene. It could also detect and correct misguided parental guidance in real time.
All five parents expected AI to serve as an expert resource. By integrating relevant knowledge into conversations, as G4 noted: “Providing real-time pedagogical insights or case studies, such as developmental benchmarks for 5-6-year-olds.” Additionally, it could offer post-conversation reports to improve parents’ long-term educational skills and provide objective assessments of children’s behavior to counteract parental biases, with G4 adding: “I hope AI can help me observe my child’s development more objectively and comprehensively.
Parents also proposed practical applications for AI functioning as a generative tool. AI could reduce communication burdens by “generating images or videos to clarify scenarios for children” or “suggesting real-time conversation strategies”. It could also produce post-interaction reports with performance feedback, as G5 mentioned: “It could analyze which guidance techniques I applied successfully and how well I executed them.
The interview sessions with kindergarten teachers highlighted key cognitive characteristics of 3-6-year-old children, informing the design of AI-assisted emotion education systems. First, the system design must align with children’s developmental stage. Considering children’s tendency toward concrete and visual thinking, the system should incorporate visual aids, such as pictorial representations, to help them better grasp emotional concepts. Additionally, given their limited language skills, the interface should guide parents on how to use age-appropriate language during emotional coaching. Second, the system should provide personalized instruction to accommodate individual differences in emotional and cognitive development. Experts emphasized that developmental abilities vary significantly among young children. Thus, it is crucial to consider each child’s specific cognitive traits and emotional states to deliver tailored learning strategies. An AI system could effectively track behavioral patterns and provide developmental benchmarks, enabling personalized adjustments to the learning process.
It is important to note that these design principles enable us to delve deeper into some of the main features of the core modules, which address three primary challenges in AI-assisted family emotion education. Since these principles are crucial design details that enhance the system’s functionality and user experience, they are not evaluated separately in subsequent studies.
5. DESIGN AND IMPLEMENTATION OF PACEE
5.1. User Interface and Design of System Modules
Based on the insights from the formative studies, we designed and implemented PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education. PACEE consists of 3 modules—Parent-Child Adventure Game, Parental Feedback Report, Child Emotional Modeling— each addressing a key challenge in family emotion education.
Figure 4. Main screens of PACEE. PACEE consists of 3 modules: (a) Parent-Child Adventure Game. Each adventure game facilitates parent-guided emotional conversations centered around a specific emotional scenario, promoting meaningful parent-child interactions. (b) Parental Feedback Report, which provides feedback on the adventure game and practical recommendations for parents. (c) Child Emotional Modeling collaboratively develops a child’s emotional profile with parental input. Parents contribute long-term observational data about their child, gathered through interactions with an AI interviewer.
5.1.1. Parent-Child Adventure Game
To address the challenge of $C_{1}$ limited parental expertise, we designed the Parent-Child Adventure Game module. Parents and children can choose scenarios from the system’s scenario database, which was initialized using formative study data to engage in conversations (e.g., performing on stage). During these dialogues, parents are guided through five steps derived from Gottman’s Five Steps of Emotion Coaching (gable1999heart). We made specific adaptations for children aged 3 to 6 based on formative study findings: (1) These children often have limited memory recall, so if they cannot remember relevant experiences, they can create a virtual scenario with their parents during the first step; (2) These children respond well to positive reinforcement, so the fourth step includes reflection on both negative and positive emotions, and the fifth step focuses on finding positive solutions. PACEE’s dialogue ultimately includes the following five stages (referred to as the FiveStepsFive\ Steps):
$S_{1}$
. Help the child recall an experience or describe a virtual scenario.
$S_{2}$
. Help the child label emotions and reason.
$S_{3}$
. Express empathy to children.
$S_{4}$
. Help the child reflect on positive and negative emotions.
$S_{5}$
. Help the child set boundaries and find positive solutions.
Within this module, PACEE offers a range of support for parents. As a dialogue facilitator, PACEE can join the conversation when prompted by the parent, which helps reduce the communication burden. As an expert, PACEE provides professional advice for each of the 5 steps, considering the child’s emotional profile. It also periodically updates real-time advice, offering communication strategies suited to the cognitive characteristics of 3 to 6-year-old children. As a tool, PACEE can generate images of the current scenario for parents to show to their children, catering to children’s preference for visual thinking. Additionally, PACEE monitors the dialogue’s progress and stage, providing useful references for parents to decide when to progress to the next stage. After completing a scenario, PACEE generates rewards like stars and medals for the children, supporting their preference for encouragement-based learning.
Our design also aligns with the PERMA model (seligman2011flourish), where PACEE encompasses Positive Emotion, Engagement, Relationship, and Achievement, contributing to a holistic approach to emotion education.
5.1.2. Parental Feedback Report
To address the challenge of $C_{2}$ insufficient feedback, we designed the Parental Feedback Report module. PPACEE evaluates parents’ performance during the Parent-Child Adventure Game, and offers actionable guidance for future interactions. Over extended use, parents have the opportunity to unlock achievement badges, enhancing their sense of accomplishment.
For the evaluation of parental performance, PACEE analyzes real-time audio of parent-child conversations based on Emotional Intelligence Theory (goleman1996emotional) and Attachment Theory (howe2012attachment). This analysis provides insights into how parents effectively navigate the emotional education process and establish secure attachments with their children through responsive and empathetic communication.
For future suggestions, PACEE offers detailed, personalized guidance tailored to individual differences between parents and children. Additionally, when formulating suggestions, PACEE considers the cognitive characteristics of children aged 3-6 to ensure the recommendations are practical and implementable.
5.1.3. Child Emotional Modeling
To address the challenge of $C_{3}$ a lack of understanding of children’s emotional states, we designed the Child Emotional Modeling module. This module combines insights from both parents and AI to overcome the limitations of singular perspectives, ensuring that the system provides personalized guidance in emotional and cognitive development.
On the AI interviewer page, PACEE conducts semi-structured interviews with parents about their children’s emotional states, capturing long-term insights from the parents. Parents can respond verbally and have the option to skip questions if they find them challenging to answer.
PACEE presents a comparative analysis of children’s emotional states on the child modeling page based on both parents’ and AI’s perceptions. The AI’s understanding is derived from objective, expert-level analysis of children’s speech during the Parent-Child Adventure Game. We draw inspiration from the Test of Emotion Comprehension (pons2004emotion; albanese2006children), and model children’s emotions across the following three dimensions:
  • Emotional understanding, including emotion recognition, emotion elicitors, emotion-related beliefs, the impact of memory on emotions, and mixed emotions.
  • Emotional expression, including emotional expression and emotional masking.
  • Emotional regulation, including emotion regulation abilities and moral emotions.
5.2. LLM-based Agent Workflow Design
Figure 5. Agent workflow of PACEE.
In this section, we introduce the agent workflow design of PACEE (see figure 5). In the Parent-Child Adventure Game, PACEE applies rapid and reflective reasoning to real-time dialogue information. Inspired by how human systems of “thinking fast and slow” (kahneman2011thinking) can be applied to generative agents (christakopoulou2024agents; lin2023swiftsage), we implement rapid and intuitive reasoning to generate real-time coaching support for parents, while employing reflective and deliberative reasoning steps to understand the emotional characteristics of the child progressively.
PACEE transcribes the real-time parent-child dialogues and monitors explicit transitions across the FiveStepsFive\ Steps by mouse clicks. The dialogue analyzer uses this information to assess the completion level of the current dialogue stage and update the node status in the dialogue graph.
The content generator module, combining dialogue information and the child’s emotional profile, applies rapid reasoning to provide timely support for parents: (1) phase advice updated at each stage, (2) real-time advice updated every half minute, (3) agent dialogue converted into voice mode, and (4) scenarios converted into image mode by text-to-image models. In addition, the content generator module integrates all dialogue information from the latest Parent-Child Adventure Game and the child’s emotional profile to generate the Parental Feedback Report for the parents.
PACEE also applies reflective reasoning within the parent-child dialogue to gradually build the child’s emotional profile based on real-time dialogue information. At certain intervals (empirically set to every five dialogue turns), the cognitive analyzer is tasked with multi-step reasoning to update the child’s emotional profile (see Section 5.3).
5.3. Modeling Children’s Emotional Profiles
Figure 6. Mechanism of PACEE modeling children’s emotional profiles collaboratively with parents. It functions in two scenarios: (1) individual parent interviews conducted by an AI Interviewer and (2) AI-facilitated parent-child emotion dialogues monitored by a Dialog Agent. The Cognitive Analyzer processes inputs from both scenarios, analyzing and synthesizing emotional profile data. The resulting Child’s Emotional Profile includes three sub-dimensions: emotion understanding, emotion expression, and emotion regulation. Additionally, it compares the parent’s understanding of the child’s emotions with observations made by the AI, offering a nuanced perspective on the child’s emotional development.
In the Child Emotional Modeling module, PACEE enables parents to contribute their long-term understanding and empathetic observations of their children in the physical world. To facilitate this, PACEE employs an AI interviewer to simulate a semi-structured interview (scenario 1). The process begins by selecting a question from a fixed list, and based on the parent’s responses, the interviewer decides whether to ask follow-up questions, adjusting the phrasing as necessary. If no further probing is required, the interviewer proceeds to the next question on the list. This list is generated with reference to the Test of Emotion Comprehension (pons2004emotion; albanese2006children).
The AI provides detailed and objective analyses of parent-child interactions (scenario 2), addressing limitations such as human attentional constraints and subjective biases.
Conversation data from both scenarios are analyzed by the cognitive analyzer. Initially, the profile extractor identifies incremental updates across the dimensions of emotional understanding, emotional expression, and emotional regulation. Subsequently, the profile integrator merges these incremental updates into the existing profile, minimizing redundancy. Furthermore, the profile integrator attempts to compare and organize entries from the AI and parents within the same dimension, presenting potential cognitive gaps to the parents.
5.4. Implementation
We implemented the core agent workflow in Python on a Flask server. The PACEE interface, written in Vue, interacts with the server by exchanging messages via WebSocket. Children’s profiles and dialogue graph information are stored in separate JSON files.
For most text generation tasks, we leveraged the GPT-4o API to balance speed and output quality. To minimize latency in agent dialogues and real-time advice delivery, the faster GPT-4o-mini model was adopted. Speech-to-text processing utilized the TINGWU API, with audio sliced and compressed on the frontend before transcription. Text-to-speech conversion relied on Alibaba Cloud’s streaming synthesis API, while image generation integrated the Stable Diffusion XL (SDXL) model.
6. USER EVALUATION
We conducted a within-subject study to evaluate how PACEE supports family emotion education collaboratively with parents. We aimed to answer the following questions:
RQ1. How does PACEE enhance family emotion education?
RQ2. How do parents collaborate with PACEE to develop children’s emotional profiles and offer guidance to children?
RQ3. How do parents perceive PACEE in addressing their challenges in family emotion education?
We included a baseline condition to compare the effectiveness of PACEE in supporting family emotion education with a generic LLM chatbot. Each pair of parent and child participants was tasked with two common emotional scenarios from PACEE’s scenario database, one under each condition.
6.1. Participants
We recruited 16 pairs of parents and children from social media platforms. Parent participants (P1-16) included four fathers and twelve mothers, with an average age of 37.69 years ($SD=3.31$). Child participants (C1-16) included seven girls and nine boys, with ages ranging from 3 to 6 ($M=5.06,SD=0.66$).
We carefully took ethical measures for participants in our user study. First, we obtained verbal consent from parent and child participants, confirming their willingness to participate. Second, we informed the participants that they had the right to withdraw from the experiment at any stage without facing any consequences. Third, parents signed a written consent form before the experiment, which detailed the complete process, the types of information we collected (video data, audio data, screen activity), and the anonymization methods. All 16 parent participants agreed to sign the consent form. Lastly, we conducted remote real-time manual reviews on the server during the experiment. These reviews included examining text and images to ensure their content appropriateness and to prevent displaying any harmful content to the children.
6.2. Baseline Condition
To compare PACEE with generic LLM chatbots, we designed a Baseline condition. In this condition, parents are provided with a brief guide to FiveStepsFive\ Steps and a standalone LLM (see Figure 7). They are required to guide children through the assigned scenario by following the FiveStepsFive\ Steps and can request assistance from the LLM via audio when needed. We do not limit the number of times parents can use the LLM during the experiment. For the LLM in the Baseline, we chose the Doubao model  [^1] https://www.doubao.com/chat/ to ensure a realistic and robust comparison. Our selection was guided by three reasons: (1) As one of the most popular free AI chatbots in Chinese families (newsDoubao1; newsDoubao2; yu2025suitability), it provides an ecologically valid baseline that participants might encounter in daily lives. (2) It has strong support for real-time Chinese audio conversations, which aligns with PACEE’s hands-free interaction flow. (3) To ensure our comparison focused on the interaction framework rather than model capability, we selected a model with competitive performance in the context of our study. In the Chinese context, Doubao-Seed-1.6-thinking-250715 model outperforms the ChatGPT-4o-latest model in the SuperCLUE Benchmark  [^2] https://www.superclueai.com (xu2023superclue).
Figure 7. Illustration of Baseline condition. In the Baseline condition, participants followed a brief guide to the FiveStepsFive\ Steps and an LLM-chatbot attached to Doubao model.
6.3. Evaluation Design
We randomly chose two scenarios from our system’s scenario database (see Table 4), which was initialized during the formative study: (1) not winning first place (referred to as Not First), where children often encounter anger and disappointment, and (2) performing on stage (referred to as Up Stage), where children often encounter anxiety and fear.
As illustrated in Figure 8, we used a counterbalanced design where each parent and child pair was randomly assigned to one of four groups. These groups were determined by the order of using the two conditions (Baseline vs. PACEE) and learning about the two emotional scenarios (Not First vs. Up Stage).
Figure 8. Design of the evaluation study. We followed a counterbalanced design where each pair of parent and child was randomly assigned to one of four groups.
6.4. Procedure
Each pair of parent and child went through a 2-hour study in our lab (see figure 8). We provided a touch-screen computer for each family and recorded the entire evaluation process.
Introduction session. Before each emotion education session, we introduced the parent participants to the designated system. We demonstrated how to interact with each system and provided tips to avoid misuse. We also provided parents with a practice session using several example scenarios to familiarize them with the procedure. In addition, we informed parents that the session required joint participation from both parents and children, and it was not intended for the child to engage solely with the AI.
Emotion Education Session Using PACEE or Baseline. In each emotion education session, we asked parents to freely use the system to learn about the specified emotional scenario with their children. Considering children aged 3 to 6 may be shy around strangers, the researcher left the room to allow parents and children to interact freely. We monitored the session via live video feed and informed parents they could contact us anytime if they encountered any issues.
Post-Questionnaire, Parent Interview, and Child Drawing. Parent participants completed a post-questionnaire to evaluate their educational experience after completing all education sessions. Then, one researcher interviewed the parents about their user experience and perceptions, while another researcher asked the child participants to use emoji stickers to represent their feelings during conversations, the scenarios discussed, and the image of PACEE. Finally, the children drew a picture to depict the scenarios discussed with PACEE.
(a) Evaluation experiment settings using PACEE.
(b) Children create drawings after the experiment.
Figure 9. Photos taken in the evaluation studies, including: (a) the setup of an evaluation experiment session, where parents and child engage in emotion education using PACEE, and (b) a drawing created by a child after the experiment session, depicting the Not First scenario discussed with PACEE.
6.5. Data Analysis
For qualitative analysis, we transcribed all parent interviews. Five HCI researchers with qualitative coding experience conducted a thematic analysis (braun2006using; terry2017thematic; braun2019reflecting). The process began with a data familiarization phase, where all researchers went through the transcripts to gain a comprehensive understanding of the content. The leading author then generated initial code themes, which were reviewed and refined with feedback from another researcher. The team then collaboratively developed a shared codebook by coding data from one participant (P1), with discussions for consensus. The remaining data were then independently coded, and disagreements were resolved through discussion. Finally, the team manually and iteratively clustered the codes to identify recurring and meaningful patterns.
Details of the statistical analysis are presented in Appendix D.
7. FINDINGS
7.1. RQ1. How Does PACEE Enhance Family Emotion Education?
7.1.1. Increased Parent-Child Interaction
Figure 10. User interaction data from the PACEE and Baseline systems during the teaching sessions. (a-d) represent data on four dimensions: interaction duration, number of dialogue turns, word count, and the count of emotion-related words. (e) Reflect the interaction duration in Five Steps. The x-axis indicates participant identity for (a-d) and teaching steps for (e), with error bars representing standard deviation. Significance analysis: (a, e) uses the Paired Sample t-test, while (b-d) are based on the Wilcoxon signed-rank test. Significance results are marked with asterisks (***: p ¡ .001, **: p ¡ .01, *: p ¡ .05).
To compare user engagement and interaction patterns between the PACEE and Baseline systems, we conducted descriptive statistics and significance testing across four key dimensions: interaction duration, number of dialogue turns, word count, and the count of emotion-related words. The results show that PACEE consistently outperforms the Baseline, with statistically significant advantages across all four areas.
For PACEE, the average interaction duration was 892.37 seconds ($SD=310.96$), with an average of 60 dialogue turns ($SD=46.979$). In contrast, the average interaction duration for the Baseline system was 560.17 seconds ($SD=194.61$), with 38 average dialogue turns ($SD=25.3$). Considering that both systems were matched in workload and participants were given flexibility in session completion, the results suggest that PACEE leads to higher user engagement, with significant differences in interaction duration ($p<0.01$) and dialogue turns ($p<0.05$). This difference was evident in both parents and children. Parents using PACEE produced an average of 1,597 words ($SD=678.56$), compared to 994 words ($SD=468.36$) with the Baseline ($p<0.05$). Children expressed an average of 370 words ($SD=240.43$) with PACEE and 165 words ($SD=102.25$) with the Baseline ($p<0.001$), as shown in Figure 10.
To further examine conversation quality alongside user engagement, we compared the number of emotion-related words used by parents and children. Under PACEE, parents used an average of 111 emotion-related words (53 positive, 58 negative), while children used 19 (8 positive, 11 negative). In contrast, under the Baseline, parents used an average of 70 emotion-related words (31 positive, 39 negative), and children used 10 (5 positive, 5 negative). Both parents ($p<0.05$) and children ($p<0.01$) are more willing to express their emotions using PACEE. These findings suggest that PACEE fosters more meaningful emotional exchanges, which are essential for supporting children’s early development of language and emotional competence (brushe2025causal; reschke2024links).
To gain a more intuitive understanding of participants’ engagement across the FiveStepsFive\ Steps of the teaching session, we calculated the average time spent at each step for both systems, as shown in Figure 10 (e). Participants’ engagement was generally higher in PACEE at all five steps, particularly in $S_{1}$ ($p<0.05$) and $S_{4}$ ($p<0.05$). Considering the limited memory capacity of preschool children (scarf2013have; roman2014assessment), the higher participation time in $S_{1}$ suggests that PACEE effectively supported participants in recalling scene details (PACEE $M=183.38$ vsvs Baseline $M=123.50$; $p<0.05$). Meanwhile, in $S_{4}$, PACEE participants engaged in significantly more reflection on positive and negative emotions (PACEE $M=163.86$ vsvs Baseline $M=85.33$; $p<0.05$).
7.1.2. Sustained Participation Over Time
To investigate the potential influence of the novelty effect, we conducted a temporal analysis by dividing each interaction session within the PACEE system into two equal halves by time (see figure 11). Our analysis revealed no statistical differences between the first and second halves of the session regarding user dialogue turns, number of expression words, or number of emotion-related words over time. This stability suggests that the high level of user participation is a consistent feature within PACEE.
Figure 11. Temporal analysis of user interaction data within the PACEE system, comparing interaction during the first and second temporal halves of the PACEE teaching sessions across three metrics: (a) number of dialogue turns, (b) user word count, and (c) the count of emotion-related words. Error bars represent standard deviation. No significant differences were found between the two periods based on Wilcoxon signed-rank tests.
7.1.3. Improved Parental Experience and Children’s Engagement Respectively
Figure 12. Post-questionnaire data on parental experience, including metrics from the NASA-TLX and the Parenting Sense of Competence Scale. We applied the Wilcoxon signed-rank tests to analyze the significance of each metric. Statistically significant ratings are marked with an asterisk (*).
Figure 12 shows parent participants’ responses to the post-questionnaire on their parental experience. Compared to Baseline, PACEE improves parents’ immediate interaction experience, with higher satisfaction ($p<0.05$), reduced effort ($p<0.01$), and less frustration ($p<0.05$). However, intrinsic psychological factors (e.g., parenting confidence or effort with the child) show no significant change during the study, likely because these factors need long-term exposure and reinforcement to yield measurable differences (bandura1997self; sandler2011long).
In addition, we analyzed children’s post-study drawings to explore their perceptions of PACEE. Thirteen of the 16 children were able to accurately reconstruct key interaction scenes, such as the stage, parents, and audience, demonstrating strong memory retention. Their illustrations often conveyed positive emotions through visual cues, including smiley faces, hand-holding gestures, and a dominance of warm color palettes. For example, C1’s drawing depicted the stage with smiling characters rendered in warm hues. Similarly, C5’s drawing was colorful, featuring balloons, smiles, and hand-holding gestures.
Beyond scene recollection, the drawings revealed nuanced perceptions of PACEE itself. Thirteen of the 16 children portrayed PACEE as a friendly and supportive figure. Common representations included anthropomorphic characters with open arms and smiling faces (C1, C8, C12, C16), friendly cartoon animals like cats with smiles (C5, C6, C10), or simplified smiley-face icons (C3, C7, C13). Personalized elements, such as C9’s drawing of PACEE resting on the child’s own head, further implied emotional attachment. In contrast, three of the 16 children (C2, C4, C11) depicted PACEE as a machine-like figure with neutral or absent facial expressions. These children also had difficulty conveying scene details and emotions in their drawings. These findings suggest that perceiving PACEE as impersonal may correlate with lower engagement.
Figure 13. Drawings created by children (C1, C5, C10) after the experiment.
7.2. RQ2. How Do Parents Collaborate with PACEE to Develop Children’s Emotional Profiles and Offer Guidance to Children?
7.2.1. Parents and PACEE Contribute Different Factors in Construction of Children’s Emotional Profiles
According to experts’ annotations, parents and PACEE play distinct roles in constructing children’s emotional profiles. Regarding data sources, PACEE derives its data from the content of parent-child communication. At the same time, parents tend to provide information based on firsthand observations of events, third-party feedback, school progress, and comparisons to their children’s childhood experiences. Furthermore, Table 5 (see Appendix E) illustrates the cognition-level themes in children’s emotional profiles identified in our evaluation studies. We found that parents prefer to extract insights from various life events, focusing on experiential knowledge of their children’s habits. In contrast, PACEE evaluates developmental capabilities by analyzing parent-child interactions, emphasizing expert assessments of abilities and developmental characteristics.
7.2.2. PACEE Provides New Insights on Children’s Profile for Parental Reflection
During the interviews, parents acknowledged that PACEE offers a more objective assessment of children’s emotional capacities. Some parents realized they had underestimated their child’s emotional understanding after reviewing the AI’s evaluation. As P2 noted, “Sometimes I didn’t realize how well my child could understand emotions, but the AI, speaking with the tone of an expert, showed me that he is truly growing.” On the other hand, some parents discovered they had overestimated their child’s abilities. As P4 shared, “Parents often view their children through a biased lens, believing they’re performing well. However, the AI objectively highlighted certain issues, and despite its evaluation diverging from my own perspective, I found the AI more rational and unbiased.” Additionally, AI can help parents understand the underlying reasons behind their child’s emotional responses. As P8 explained, “When my child participated in an English competition, he was very reluctant to go on stage. I assumed he was nervous and anxious because he was afraid of the stage, but the AI suggested that the anxiety might have stemmed from the absence of his mother’s support.
In a broader view, parents express opinions on AI’s advantages in modeling children’s emotional profiles, including more comprehensive summaries, more objective perspectives, and more educational insights. Also, AI has potential disadvantages such as lacking long-term data, lacking specificity, and making incorrect inferences. These factors can lead to cognitive differences between parents and AI regarding children’s emotional states. We further inquired how parents perceive these differences. 14/16 parents considered these differences between humans and AI as valuable new insights. They acknowledged that the individual suggestions made by AI are derived from broader patterns observed in children of similar age, allowing for more long-term and holistic reflections on a child’s emotional tendencies. As P16 explained, “I can only see the emotional issues my child is currently experiencing and how he copes with them, I have no experience dealing with newly emerging emotions.” In contrast, AI offers a more rational point of view that can help parents break free from fixed thought patterns. P8 remarked, “Because our communication with the child is usually smooth, we’ve fallen into a comfortable but fixed mode of interaction. While this is pleasant for daily life, it might hinder development in certain areas. The involvement of AI could reduce this limitation.
However, 2/16 parents expressed concerns about the accuracy of AI-generated insights, viewing discrepancies from their own understanding as potential system errors. P15 commented, “The AI detected a sense of compromise in my child during the conversation, but I doubt he has the capacity to express such emotion yet. The term ’compromise’ seems imprecise.” Some parents also believed that AI could not detect when a child was lying, whereas their familiarity with their child enabled them to make more accurate judgments. As P11 noted, “AI can’t tell whether what the child says is sincere or not —it records everything regardless of accuracy. So if the child lies, it might offer the parent an incorrect solution.
7.2.3. Parents Appreciate Open and Positive Advice from PACEE
Parent participants received a total of 513 updates of real-time advice during the studies. Two experienced HCI researchers reviewed system logs and video recordings to determine whether the parents adopted the suggestions. They then iteratively developed classification criteria for the accepted advice, selecting the most prominent type in cases where advice fell into multiple categories. Table 6 (see Appendix E) summarizes the types of accepted real-time advice. We found that the types of advice most frequently adopted by parents were collaborative problem-solving and open-ended questioning. Parents also mentioned the usefulness of open-ended questioning in the interview, as P2 stated, “Real-time advice helps me realize some issues I hadn’t thought about, such as prompting my child to provide more details. Sometimes I think it’s okay when a child says one word, but the advice reminds me to ask for more details.” Furthermore, advice on scenario simulation, concrete suggestions, positive encouragement, and empathy and acceptance also helped some parents.
7.2.4. Parents Accept PACEE in Collaborative Guidance, but with Concerns
Parents shared their perspectives on integrating PACEE into the emotional guidance process during the interviews. Participants consistently viewed themselves as the primary agents in their children’s education. Nine obviously reported making the greatest contribution during the experimental sessions, whereas only P10 regarded the PACEE as having had a larger impact. Furthermore, P5, P8, and P13 explained that their main role was to evaluate and select which of the PACEE recommendations to implement, while P6, P9, P10, and P13 credited the role of the system in providing the specialized professional expertise necessary to complete the tasks.
In addition, parents expressed a positive attitude toward incorporating PACEE into parent–child dialogues. All participants accepted PACEE involvement, though their preferred levels and modes of engagement varied. Some parents demonstrated a high degree of acceptance (P5: “My kid really like it(to receive emotional coaching from PACEE). I think I am very receptive.”), whereas others felt that PACEE was not strictly necessary, especially in familiar emotional scenarios (P16: “We often discuss this topic so my child no longer need emotional guidance in this situation.”). Parents recognized the system’s capacity to deliver valuable, professional, and targeted recommendations. For example, P8 highlighted that such suggestions made parent–child communication more effective, “He felt I understood him better, so he answered every question loudly and confidently.
When discussing future interactions with PACEE, participants articulated clear expectations about PACEE’s role in parent–child dialogues, though they differed markedly on its ideal interlocutor. Some parents (P2, P16) envisioned AI forging a peer-like relationship with their child, simultaneously acting as an educator. Others (P8) preferred that PACEE interact only with the parent, acting as a co-educator, and avoid direct communication with the child. A third group of participants (P10, P13, and P15) believed that AI should adopt a neutral stance as an observer—providing limited, objective, data-driven suggestions to parents and offering age-appropriate guidance to children, to support emotional development, rather than engaging directly with either parent or child. For instance, P10 argued that “It can serve as a mediator (in parent–child emotional education).” P13 and P15 prefer PACEE to offer suggestions from a third-party perspective for both parents and children.
Crucially, parents asserted firm boundaries around AI intervention. Some parents rejected features that impose targets on children (P4: “For example, setting goals for children… I think it is up to parents to do it.”), and they were adamant that PACEE should not encroach on sensitive family matters (P1: “Don’t press further… for example, on topics that parents find taboo and children find sensitive. ”), or strictly require compliance with its advice (P11: “If it(PACEE) tells me I have to do something, I definitely won’t accept it.”). Moreover, participants cautioned that AI should not replace emotional intimacy and empathetic support intrinsic to parent-child bonding and the cultivation of self-confidence, such as “(I hope she can) build her confidence by communicating with others.”(P2), “I prefer she get emotional companionship from people…”(P5).
7.3. RQ3. How Do Parents Perceive PACEE in Addressing Their Challenges in Family Emotion Education?
The subjective ratings provided by parents indicate that PACEE significantly outperformed Baseline in tackling the three main challenges of family emotion education (see figure 14), with statistical significance observed in each case ($p<0.01$).
Figure 14. Post-questionnaire data focusing on the extent to which PACEE or Baseline address the three key challenges. We applied the Wilcoxon signed-rank tests to analyze the significance of each metric. Statistically significant ratings are marked with an asterisk (*).
Interviews with parents further confirmed that PACEE plays a multifaceted role in addressing the three main challenges of family emotion education.
7.3.1. Enhancing Parental Expertise and Communication Confidence
In terms of $C_{1}$ limited parental expertise, the Parent-Child Adventure Game module significantly alleviated parents’ difficulties in situations where they “don’t know what to say” or “don’t know how to guide.” During the interviews, parents acknowledged that they find it difficult to perform some steps alone, including recalling an experience ($n=4$), labeling emotions ($n=8$), expressing empathy ($n=8$), reflecting on emotions ($n=5$), and finding positive solutions ($n=6$).
The system’s dialogue suggestions were praised for mitigating these expertise gaps. First, it provides a structured framework for emotion communication. Two parents specifically highlighted how the AI facilitated guiding emotional conversations with children in a gradual, structured manner. As P8 noted: “The systematic process — understand the child first, then ask about their feelings, then summarize, encourage, and finally solve the problem. This step-by-step guidance is great, creating remarkably smoother interactions.” Second, several parents appreciated how AI helped them adopt their child’s perspective. P10 explained: “Placing us parents in a conversational setting that feels more like talking to a friend actually encourages us to respond more actively and calmly. It’s because of this setup that I’m more willing to take the child’s perspective, understand them, and help them manage their emotions.” Lastly, the dialogue suggestions offer ideas to steer the dialogue stream effectively. P16 stated, “AI’s reminders can offer us more ideas. During the reflection stage, both the child and I are thinking, and it’s easy to get caught up in the moment. The AI’s prompts provide a clearer perspective, which helps make the conversation more in-depth.
Beyond providing suggestions, the module actively enhanced parent-child communication and boosted parental confidence during real interactions. Most parents reported increased initiative and comfort in emotional conversations after using the Parent-Child Adventure Game module. Ten participants confirmed that the system’s dialogue suggestions provided concrete, child-friendly advice for expressing emotions. As P5 explained, “It can serve as a reference tool. For example, if you have an idea but don’t know how to express it, the system helps you organize your thoughts and express them more clearly and reasonably.” Additionally, two parents mentioned that AI-provided prompts served as useful entry points, promoting more open communication between children and parents. P1 explained: “As a father, I don’t talk with my child that often. For example, instead of saying, ‘Do you remember your last kindergarten performance?’ it feels more natural to ask, ‘How was your last performance?’ So the detailed suggestions provided by the system are quite helpful for me as a reference.” Some parents also noted that the module facilitated more reciprocal exchanges, helping to enhance parent-child relationships.
While the benefits of AI assistance were frequently highlighted, some parents expressed concerns. A few parents felt that frequently checking the AI’s suggestions could slightly disrupt the conversational flow. One parent worried that the AI’s scripted suggestions might make children feel nervous or overly reliant. Another parent emphasized that AI should act as an aid rather than lead in communication, suggesting it should provide direction rather than take over the dialogue entirely. Furthermore, some saw the structured nature of AI prompts as potentially limiting the child’s freedom to express diverse experiences. There was also concern that excessive reliance on AI-generated expressions might hinder parents’ own communicative abilities.
7.3.2. Providing Trackable Feedback and Actionable Future Strategies
In terms of $C_{2}$ insufficient feedback, the Parental Feedback Report module effectively addressed pain points such as “lack of timely feedback” and “invisible outcomes”. Eight parents highlighted how AI-generated analyses shed light on previously unnoticed shortcomings in their parenting approaches, which prompted them to engage in significant self-reflection. For instance, P6 shared, “The system reminded me to avoid interrupting. I hadn’t realized I was doing that, but it made me see that I still need to be more patient in communication.” Furthermore, the module provided a structured review by organizing conversation highlights, which five parents found particularly beneficial in bridging feedback gaps and making abstract parenting efforts trackable. This transformation into visible progress was crucial for parents to assess their parenting strategies.
Beyond reviews on past interactions, the module also provides actionable future strategies. Eight parents appreciated the evidence-based nature of the strategies provided, while three were pleased with the introduction of new, beneficial perspectives. P12 noted, “For example, when a child loses control, besides suggesting enjoyable activities, the system also mentions deep breathing and meditation. I used to think those were only for adults, but this made me realize they might work for kids too.”. Also, P1 noted: “AI can indeed offer suggestions we wouldn’t normally think of, like encouraging the child to observe how others handle situations or simulate a scenario at home — these are quite practical.” Parents praised that these suggestions provide a clear foundation for future communication.
The module also incorporated practical features that increased its utility. The “review and scoring” design was highlighted by four parents as a useful tool for evaluating the effectiveness of their interactions, as it allowed them to pinpoint areas needing improvement. P15 noted: “I think the scoring in the review session is useful. It helps me understand which parts of my communication need improvement. Everyone is busy, so this kind of tool is really practical.”. Additionally, motivational elements such as the “badge achievement system” were noted to enhance parental engagement and motivation.
Despite these positive aspects, some concerns remained. A few parents pointed out that a single session was not enough to assess real outcomes, and some found the suggestions too generic, which points to the need for the AI to better adapt to specific parental contexts.
7.3.3. Improving Parental Understanding and Grasp of Child Developmental Trajectories
In terms of $C_{3}$, a lack of understanding of children’s emotional states, the Child Emotional Modeling module helped parents break through cognitive blind spots by leveraging data-driven insights to reveal their child’s emotional traits and behavioral patterns. Twelve parents reported that the system’s scientifically grounded feedback provided “new knowledge and dimensions”, helping them overcome cognitive blind spots regarding their child’s emotions and behavior. As P12 explained: “The system’s questions prompted me to think about aspects I hadn’t paid attention to before. The conclusions helped me understand my child more systematically and comprehensively. It was very helpful.” Thirteen parents further emphasized how AI generated “new angles” or “reference dimensions”, saying that the system helped them move beyond existing subjective impressions to enable more objective and reflective observations. This objective analysis was particularly valued in understanding the causes of children’s emotions, especially in labeling children’s emotions ( $S_{2}$). P6 explained: “AI can analyze the causes of a child’s emotions more objectively, whereas we as parents often bring in our own biases or emotions, making it hard to view the situation in a scientific and objective way.
Parents also discussed the alignment between their views and the system’s analyses. Some parents said the system’s views were highly consistent with their own, enhancing their trust. Others felt that when discrepancies arose, the system helped them see where their perceptions diverged — even serving as a useful reference for resolving differences within the family.
Beyond individual insights, the module also contextualized children’s behavior within broader developmental frameworks. Four parents mentioned that the AI’s use of large-scale data gave them their first systematic understanding of “common behaviors at this age,” which helped them better assess their child’s traits. Additionally, several parents valued the forward-looking nature of the AI’s suggestions, which strengthened their overall grasp of their child’s developmental trajectory.
Despite these advantages, several limitations were noted concerning the AI’s understanding and accuracy. Seven parents were concerned that the limited interaction time during the trial meant the generated profiles might not be comprehensive. Four parents noted that the system cannot verify the truthfulness of the information entered, which may lead to biased analysis. P1 noted: “AI can only analyze based on what the child says, but children don’t always express their true feelings. Sometimes they even lie—for example, they might be very nervous but say they’re fine. This affects the AI’s judgment.” Two parents expressed discomfort with the scoring system and said they didn’t want to receive low scores from the system.
8. DISCUSSION
In our study, we designed PACEE from a parent-centered view. We based this approach on recognizing that the role of parents is essential in early childhood education and that the ability to interact with AI varies among children aged 3 to 6 years old. PACEE leverages multiple generative AI technologies (LLMs, text-to-image, text-to-speech) to address challenges for parents ( $C_{1}$- $C_{3}$). Our evaluation studies show that parents perceive PACEE to have successfully addressed these challenges, reducing their effort and boosting their satisfaction. Here, we further discuss the impact of generative AI on early childhood family education and the impact of PACEE on Gottman’s Five Steps of Emotion Coaching framework. We then propose broader design implications for parent-centered, AI-supported education systems.
8.1. PACEE Enhanced Children’s Learning through Improving Parents’ Performance
Prior work on AI-assisted child education has explored AI systems directly coaching the child (seo2024chacha; tang2024emoeden), AI systems for intervening with parents (entenberg2023ai), and systems that facilitate parent-child interaction simultaneously (zhang2022storybuddy; theofanopoulou2022exploring). Similarly, we identify the need for generative AI support in our work. However, our context is different in that children aged 3 to 6 often face challenges in interacting independently with AI (sun2024exploring; bush2025bridging; escobar2022guidelines; thill2022modelling), and early childhood emotion coaching requires more substantial support from parents (theofanopoulou2022exploring; zhu2025maternal; lee2023effects). Also, prior work has expressed concerns that AI alone may not be able to adequately respond to young children’s emotional needs (kurian2025ai; kurian2024no). Thus, we propose a parent-centered, AI-supported system for parent-child interaction. This parent-centered interaction pattern, where AI primarily assists parents rather than directly interacting with children, was preliminarily explored in therapeutic domains, such as for children with language delay (hwang2014talkbetter). We further extend this pattern to the domain of early-childhood education.
In our work, PACEE enhances children’s learning outcomes mainly by improving parents’ performance. By positioning itself as a strategic partner rather than merely a tool, PACEE has fostered more effective communication between parents and their children, including higher interaction engagement and more in-depth emotional dialogues. Consequently, child participants exhibited higher engagement levels and used more emotion-related words during the teaching sessions.
However, the potential impact of generative AI directly interacting with children aged 3-6 in emotional education needs further exploration. In our study, we emphasized parents’ participation and informed parents that PACEE was not intended for the child to engage solely with the AI. This decision is due to findings from our formative study, where all five participating parents preferred to maintain their primary role in emotional guidance. Regarding the extent of AI directly interacting with the children, parents showed varying acceptance levels in the interviews. We also found that children’s AI literacy significantly varied based on parents’ attitudes. Therefore, the future design of emotion education systems needs to customize AI interaction levels based on parental acceptance and the child’s AI literacy. Also, longitudinal studies are required in order to assess the impact of varying degrees of AI interaction on children’s emotional development over time.
Under our parent-centered interaction pattern, one challenge is to help parents better understand their children. To address this, we propose a framework for AI-parent collaborative modeling of children’s emotional profiles. In our framework, parents tend to contribute experiential knowledge of their children’s habits, while AI tends to evaluate developmental capabilities by analyzing parent-child interactions. This approach can also be extended to other areas of early childhood education, building upon prior work on personalized education systems. Prior systems either allow parents to manually set basic information and preferences (zhang2022storybuddy; fang2024edulive; chen2025characterizing) or automatically assess children’s learning progress (lee2024open). By combining the strengths of parental expertise and AI-driven insights, our framework fosters a dynamic process of mutual learning. Parents and AI can collaboratively refine their understanding of the child’s needs, enabling more effective and personalized educational strategies.
8.2. The Impact of PACEE on Five Steps Framework
Here, we analyze the impact of our approach on the FiveStepsFive\ Steps framework. To begin with, considering the developmental characteristics of 3-6-year-old children, PACEE adopts two variations: (1) the involvement of parents in recalling experiences or creating a virtual scenario, considering the limited memory and expression capabilities of 3-6-year-olds; (2) the integration of the PERMA model of positive psychology into the Five Steps, including reflection on positive emotions and finding positive solutions. In our evaluation studies, we found that the interaction patterns and dialogue analysis align with our adjustments.
We further perform an analysis of the implications of integrating generative AI into the FiveStepsFive\ Steps framework. Generally, PACEE supports each of the five steps from the parent’s perspective, consistent with prior work exploring design interventions to support emotional coaching based on this theory. However, we also highlight that the design needs further consideration in the unique developmental context of children aged 3-6: (1) Both parents and children should avoid overdependence on AI to preserve skills like free expression, problem-solving, and empathy. (2) Conflicts between parental opinions and AI recommendations must be resolved thoughtfully. (3) Advanced technical solutions are required to tailor AI advice more accurately to real-life individual contexts. Here is a detailed analysis of each step:
$S_{1}$
Help the child recall an experience or describe a virtual scenario. Previous work has designed chatbots to encourage children to share personal experiences and associated emotions (seo2024chacha; grove2021co). In contrast, our approach focuses on supporting parents who directly encourage their children to recall past experiences. Given the limited memory recall of 3- to 6-year-olds (xuelian2018developments), it’s more effective for parents to assist in recalling and narrating experiences than to let children do it alone. From the parents’ perspective, PACEE was seen as a valuable tool, offering helpful prompts and contextual cues that facilitated memory recall and fostered more open communication between parents and children. Additionally, the parent-child interaction duration for this step is significantly increased by PACEE, which, in turn, helps strengthen the parent-child bond.
However, several limitations were also identified. For instance, some parents reported that the scenarios provided by AI lacked authenticity or real-life relevance. Additionally, the structured nature of AI-generated prompts was seen by some as potentially restricting the child’s ability to freely express diverse experiences. These findings suggest the need for future work to focus on improving the flexibility of AI-generated prompts, ensuring they align more closely with real-life contexts and encourage creative expression.
$S_{2}$
Help the child label emotions and reason. Previous studies have designed systems that collect data from children to assess their emotions, such as storytelling chatbots (santos2020therapist) or audio-visual recognition models (nojavanasghari2016emoreact). Similarly, PACEE leverages LLMs to analyze real-time parent-child conversations, aiming to identify the emotions underlying children’s expressions. Differently, PACEE also provides precise vocabulary and child-friendly explanations for parents during this process —a feature participants found particularly valuable. Some parents also noted that AI provided objective insights into the causes of emotions, which facilitated deeper discussions.
However, our findings revealed that AI interpretations of children’s emotions occasionally contradicted parents’ own understandings. This highlights the need for future research to develop a third-party mediator between parents and AI, which could help bridge differing perspectives and promote collaborative decision-making. Additionally, there were concerns that excessive reliance on AI-generated expressions might hinder parents’ own communication abilities, which was also noted in prior studies on parental dependence on AI (szondy2025artificial). Future work should explore strategies to balance AI assistance with cultivating parents’ natural communicative abilities, such as offering optional training modules or limiting AI intervention in certain scenarios.
$S_{3}$
Express empathy to children. This step revealed a firm reliance on AI support because parents often neglect this step in home education. In the evaluation studies, this step had the shortest duration. Prior work has tried to enhance this step by encouraging parents to adopt the child’s perspective during moments of conflict (yoo2019mom). Similarly, in our work, PACEE was valued for providing reference expressions for parents that made children feel genuinely understood and assisting parents in adopting the child’s perspective.
Prior work has argued that AI tools should assist parents rather than entirely replace their role in areas such as storytelling (sun2024exploring). Similarly, our participants cautioned that AI cannot replace the emotional intimacy intrinsic to parent-child bonding. Expanding on this, we apply these insights to the context of emotion coaching, which involves navigating sensitive emotional exchanges. We emphasize the importance of parents providing additional support through emotional expressions, such as using body language (e.g., hugging). Future research should also investigate how varying levels of support impact children’s emotional development.
$S_{4}$
Help the child reflect on positive and negative emotions. PACEE significantly increased the interaction duration for this step. PACEE was regarded as helpful in supporting this stage by encouraging children to consider alternative outcomes and assisting parents in presenting new viewpoints. Several parents also credited PACEE with offering scientifically grounded guidance and simulating the child’s perspectives, which boosted their confidence in leading emotional reflections. This is similar to prior work regarding in-situ suggestions to help parents explore the various response strategies (choi2024aacesstalk). While PACEE demonstrated significant benefits, parents also expressed the desire for AI systems to provide more personalized guidance through repeated use and long-term data collection.
$S_{5}$
Help the child set boundaries and find positive solutions. Research has shown that collaborative problem-solving can significantly reduce children’s behavioral challenges and ease parenting stress (heath2020collaborative). Building on this, we adapted the approach to an emotion-sensitive context with AI support to fill cognitive gaps for parents. PACEE was praised for its clear and structured suggestions, which made setting boundaries and developing solutions more practical and achievable. Parents also noted that the AI broadened their perspectives by offering fresh approaches and precise language for communicating behavioral expectations. This step received the most positive feedback regarding AI assistance, with participants spending the longest time on it during the evaluation study.
However, some areas for improvement were noted. Parents expressed a need for more personalized, context-specific advice to better fit their family dynamics. They also opposed features that imposed rigid goals on children, feeling these could undermine the collaborative nature of the process. Lastly, they expressed concerns about children becoming overly reliant on AI for finding solutions, echoing previous research on children’s dependency on AI (solyst2024children). To address these concerns, future AI systems should promote children’s independent thinking by encouraging open discussions and fostering creative problem-solving.
8.3. Design Considerations for Parent-centered, AI-supported Education Systems
Parent interviews also revealed several challenges that need to be addressed. Combining our findings with prior work, we propose design considerations for future parent-centered, AI-supported education systems.
First, customize AI’s aiding modules based on parents’ profiles, in addition to their children’s. Some recent work has integrated children’s profiles or preferences to facilitate a more personalized coaching experience (clabaugh2019long; tang2024emoeden; zhang2022storybuddy). This study supplements prior research and demonstrates the necessity of customizing AI systems based on parental profiles. During our interviews, we found that parents expressed significantly varied opinions about the “ideal support” from PACEE. Concerning content difficulty, some experienced parents criticized that AI’s guidance overlapped heavily with their own approaches, while a few parents mentioned that some suggestions were too complex to understand. Concerning the frequency of suggestions, some parents felt that frequently checking the AI’s advice could slightly disrupt the conversational flow. In contrast, others believed that PACEE could offer more real-time interactions.
These findings highlight that parent-centered educational systems should accommodate such differences in parents’ needs. For example, parents’ preferences regarding the appropriateness of early-childhood digital play, which may influence the quality of AI-assisted activities at home (isikoglu2019parents), should be collected beforehand. Furthermore, various parental inputs can be captured during parent-child interactions, besides audio data from dialogues. These inputs may include whether parental feedback during interactions is positive or negative (robinson1981dyadic), behavioral differences between fathers and mothers (cutler2020fathers), and parents’ emotional regulation abilities (bariola2011child). By incorporating diverse data sources, the system can better tailor its functions to the specific needs of each family.
Second, avoid excessive reliance on AI capabilities and promote transparent and negotiable collaboration. In the evaluation studies, parents expressed worries that the system cannot verify the truthfulness of the information entered, which may lead to biased analysis. They also noted that AI’s interpretation of a child’s emotions may contradict the parents’ understanding. These concerns align with the communication aspect of cooperative AI (dafoe2020open). By explicitly sharing information, AI agents and parents can gain insights into each other’s behaviors and intentions more effectively.
Here, we recommend leveraging research findings from explainable AI to enhance transparency in collaboration (liao2020questioning). Parents need to understand the rationale behind AI-generated assessments or suggestions, thus enhancing trust in AI’s decisions (kim2024establishing). In designing future systems, it is crucial to consider how to balance the parents’ cognitive burden of processing information with the goal of enhancing trust by providing a comprehensive explanation, which may be related to parents’ diverse decision-making styles (van2025selective).
Considering the potential conflicts between parents and AI systems, we recommend the introduction of negotiation tools (crandall2018cooperating; jennings2001automated), as the current context lacks a formal agreement protocol. In the context of emotion education, it is essential to develop interactive negotiation mechanisms to address varying interpretations of children’s emotional states. Furthermore, AI can serve as a mediator to reconcile differing perspectives on children’s emotions among family members, a role that some parents expressed interest in during the interviews.
Third, respect parents’ boundaries regarding AI intervention and provide them with sufficient autonomy. Prior work has explored the area of parents’ boundaries, such as parental concerns with AI-driven technologies in their child’s healthcare (sisk2020parental), showing that parents value shared decision-making but may resist interventions that undermine their sense of control. We supplement prior research and expand this area to include both emotion-sensitive contexts and early childhood education domains.
In our context, children’s emotional education may involve sensitive family matters. In the interviews, parents expressed that they do not want AI to discuss family privacy or parents’ professional information. Future work could look into value-sensitive AI design frameworks (umbrello2021mapping) and provide information sensitivity detection techniques that align with diverse parental values. Moreover, emotional education may involve delicate feelings or negative topics that could affect children’s mood. To address this issue, PACEE incorporated the PERMA positive psychology model into its design to reduce the emphasis on negative content. Future work could investigate how to further reduce emotional overwhelm that may be perceived as harmful by parents. For example, adaptive AI systems could use real-time sentiment analysis to detect emotional intensity and compare it with predefined emotional boundaries set by parents.
As for education for children aged 3 to 6, parents consistently viewed themselves as the primary decision-makers in their children’s education. In the interviews, parents expressed concerns about the possibility of AI imposing rigid goals on children and compelling parents to follow prescribed steps. Future systems should prioritize bidirectional discussions with parents, rather than simply providing prescriptions. Furthermore, they should ensure parents maintain authority over critical decision-making in parent-AI collaboration.
8.4. Limitations and Future Work
This study has several limitations. First, the participants’ cultural background was limited to Chinese parents and children, which may constrain the cultural generalizability of our findings. Second, most of the parent participants were mothers. This reflects the broader caregiving patterns noted in prior research (mcbride1993comparison; schoppe2013comparisons; hernawati2020differences), and likely represents the primary user group for a system designed to support caregivers of 3-6-year-olds in the real world. However, this imbalance may limit our ability to capture the full range of parental perspectives. Future work should recruit across more diverse cultural and gender contexts, and systematically examine how fathers and mothers may differ in their collaboration with AI to support children.
Third, a limitation of the current study design is the inability to fully isolate the foundation model variable. While our study’s primary focus was on the interaction framework rather than raw model output, as supported by our findings and clarified in Section 6.2, we acknowledge that this does not fully eliminate the potential bias inherent in the two models. Therefore, future work could employ a 2x2 factorial design — (Interface: Baseline vs. PACEE) x (Model: Doubao vs. GPT-4o) — to further isolate the effect of the model.
Besides, PACEE currently supports audio input and fixed UI interactions. Future developments could, with parental consent, enable the collection of children’s gestures and facial expressions to better assess their emotional and behavioral states during interactions.
Finally, our study was conducted in lab-based settings using a touchscreen laptop, which may not fully reflect the long-term usage experience at home. In addition, first-time exposure to PACEE’s interactive features may have temporarily increased participants’ engagement, reflecting a “novelty effect” commonly faced by HCI research (stromel2024narrating; soler2024arcadia; davis2025sketchai). To address this issue, we provided parents with a practice session to familiarize them with the procedure. While our analyses did not reveal significant differences in comparative outcomes between early and later interactions, future work should explore how such novelty influences parents and children during longer interactions. Future work should also implement longitudinal in-home deployments to better understand how parent-AI collaborative emotion education unfolds in natural family environments.
9. CONCLUSION
We introduced PACEE, an assistant designed to support family emotion education through parent-AI collaborative emotional profiling and guidance. Drawing on interviews with kindergarten teachers and parents, PACEE leverages generative AI to address the key challenges identified in family emotion education. Our evaluation study with 16 pairs of parents and children demonstrated that PACEE enhanced parent-child engagement, fostering deeper emotional exchanges. Furthermore, PACEE significantly alleviated parental challenges, resulting in greater satisfaction and less effort from parents. We hope our work will advance emotion coaching theory in both family and AI-assisted contexts. Additionally, we hope to inspire the development of parent-centered, AI-supported education systems that empower families to promote children’s well-being.
Appendix A Demographic Details of Participants in the Formative Study
Table 1. Demographic details of kindergarten teachers in the interviews.
Alias Age Gender
K1 31 Male
K2 29 Female
K3 30 Male
K4 48 Female
K5 31 Female
Table 2. Demographic details of parents in the interviews.
Alias Age Gender
G1 37 Female
G2 45 Female
G3 41 Female
G4 38 Male
G5 44 Female
Table 3. Demographic details of participants in the observational experiments.
Alias Parent Child
Age Gender Age Gender
O1 37 Male 3 Boy
O2 37 Female 6 Boy
O3 41 Female 4 Boy
O4 43 Female 6 Boy
O5 33 Female 5 Girl
Appendix B Common Scenarios of Negative Emotions in Children Aged 3-6
See Table 4.
Table 4. Our formative study identified common scenarios of negative emotions experienced by children aged 3-6.
Type Example Scenarios Common Emotions
Separation
Leaving parents for school;
missing parents; waking up alone
Anxiety, sadness
Peer conflict
Disputes over toys/turns; sibling
quarrels; having belongings taken
Anger, frustration
Social setbacks
Being ignored by peers;
unable to join new groups
Loneliness, sadness
Physical discomfort
Injury, illness, falls, or
grogginess upon waking
Distress, sadness
Autonomy violation
Forced task termination (e.g., iPad
taken away); not winning first place
Anger, disappointment
Negative feedback
Teacher corrections;
critical remarks from peers
Sadness, shame
Stressful challenges
Challenging sports/games;
performing on stage
Anxiety, fear
Appendix C Metrics in User Evaluation
Here, we elaborate on the metrics applied in our user evaluation.
Post-Questionnaire. We used a questionnaire consisting of three sections. The first section consisted of NASA-TLX (hart1988development) metrics to measure the mental demand, effort, and frustration of parents when using AI systems to provide emotional guidance to children. The second section used the Parenting Sense of Competence Scale (PSOC) (gibaud1978development) to measure parents’ confidence, satisfaction, and self-efficacy. The last part focuses on the extent to which PACEE or Baseline address the three key challenges ( $C_{1}$- $C_{3}$) in family emotion education. All ratings were done on a 7-point Likert scale.
Semi-structured interview. In the semi-structured interview, we asked parents about the impact of the AI system on emotion teaching between parents and children, its role in overcoming challenges, and its effect on the FiveStepsFive\ Steps. We discussed their experiences with each module of PACEE and broader topics like AI’s strengths in collaborative teaching and the essential roles of parents. The interviews were audio recorded, transcribed, and then analyzed thematically (clarke2017thematic) by four experienced HCI researchers. To ensure consistency with annotation standards, each part of the analysis was reviewed by at least two researchers.
Appendix D Quantitative Data Analysis
We transcribed the audio recordings from the teaching sessions of both the PACEE and Baseline systems into text and then conducted automated quantitative analyses. The identification of emotion-related words was implemented using the NTUSD dictionary (ku2007mining). To better analyze the transcript data, we applied descriptive statistical methods, calculating the mean and standard deviation for each of the four dimensions among all 16 groups of participants. Additionally, we employed significance testing methods to evaluate the statistical significance of the collected data. We used the Paired Sample t-test for interaction duration data and the Wilcoxon signed-rank test for the number of dialogue turns, word counts, and emotion-related word counts, based on the distribution characteristics of the data. Furthermore, we calculated the mean and standard deviation of dialogue durations across the FiveStepsFive\ Steps.
Appendix E More Findings in User Evaluation
See Table 5 and Table 6.
Table 5. Parents and PACEE contribute different themes in children’s emotional profiles in the evaluation studies.
Source Theme Sub-theme (Partial) Example
Parent
Knowledge
and Events
Emotions learned at school;
Proficiency in describing
certain emotions
C2: Complex emotions, such as fatigue, are typically
guided appropriately by kindergarten teachers in class,
so they should be distinguishable.
Contextual
Responses
Emotional reactions or
regulation strategies for
different scenarios
C11: When he fails in a game, it tends to trigger his
specific feelings of frustration.
Behavioral
Habits
Habits of mood adjustment;
Seeking external help;
Tendency to go to parents
C6: At present, she cannot calmly face negative
emotions; she still seeks help from his mother or
those she feels close to.
Growth
Trends
Comparison with early
childhood
C12: When he is happy, he directly expresses his joy.
When he was younger, he would jump around and
even scream and shout.
PACEE
Specific
Emotional
Capacity
Understanding, expressing,
or positively regulating
specific emotions
C15: When facing possible failure or competition, she
considers responding to challenges through effort and
learning, showing her progress in stressful situations.
Macro
Emotional
Capacity
Ability to identify emotion
origins or others’ emotions
C2: …and can recognize the emergence of these
emotions in specific situations (such as speaking
on stage). He can feel the impact of the audience’s
reactions on himself. This indicates he can identify
the source and influence of emotions.
Capacity
Development
Trends
Emotional capacities
being developed
C14: He demonstrates a certain awareness of strategies
for emotional regulation, indicating he is developing
the ability to regulate emotions and can attempt to
use different methods to cope with emotions.
Languaguistic
Characteristics
Directness of language
expression; Use of non-verbal
communication methods
C13: He can directly express his emotional state but
may lack a more extensive and nuanced vocabulary to
describe his feelings.
Table 6. Types of accepted real-time advice during the studies.
Type Number Description Example
Open-ended
Questioning
22
Asking open-ended questions to
encourage children to describe their
emotions, thoughts, or experiences.
Do you have any special feelings
when redeeming your points? Do
you feel happy or a bit unsure?
Scenario
Simulation
16
Setting up or recalling specific
scenarios to encourage role-playing
or imagination.
Imagine yourself as a brave little
hero, like My Little Pony—you can
show your courage and talent on stage.
Concrete
Suggestions
7
Providing clear, specific strategies
to help children manage emotions
or enhance positive feelings.
We can go to the small playground
on the rooftop for a run, or maybe
find something delicious to eat.
Positive
Encouragement
9
Using praise and positive feedback
to reinforce children’s positive
behaviors and emotional experiences.
You did a great job jumping rope—
210 jumps is really impressive!
Empathy and
Acceptance
16
Expressing empathy and acceptance
towards children’s emotions, making
them feel understood and supported.
I understand you might feel nervous
on such a big stage. Dad has felt like
that before, too.
Collaborative
Problem-solving
31
Discussing solutions together with
children.
Let’s think together about what we can
do to perform even better in the next
competition.

### v2

@@
 Emotion education is a crucial lesson for children aged 3 to 6.
-However, existing technologies primarily focus on promoting emotion education from the child’s perspective, often neglecting the central role of parents in guiding early childhood emotion development.
+However, existing technologies primarily focus on promoting emotion education from the child’s perspective, often neglecting the central role of parents in guiding early childhood emotion development at home.
 In this work, we conducted co-design sessions with five experienced kindergarten teachers and five parents to identify parental challenges and the roles that AI can play in family emotion education.
 Guided by these insights, we developed PACEE, an assistant for supporting parent-AI collaborative emotion education.
-PACEE enables parents to engage in emotional dialogues about common scenarios, with multiple forms of support provided by generative AI.
-It combines insights from parents and AI to model children’s emotional states and collaboratively delivers personalized, parent-mediated guidance.
+PACEE enables parents to engage in conversations about common emotional scenarios, with multiple forms of AI support to address parents’ challenges.
+It combines insights from parents and AI to model children’s emotional states and delivers personalized, parent-mediated guidance.
 In a user study involving 16 families, we found that PACEE significantly enhances parent-child engagement, encourages more in-depth emotional communication, and improves the parental experience.
-Our findings advance emotion coaching theory in both family settings and LLM-assisted contexts, offering valuable insights for designing AI-supported, parent-centered family education systems.
+Our findings advance emotion coaching guidelines for family education in the era of generative AI, offering valuable insights for designing AI-supported, parent-centered family education systems.
 Parent-child Interaction, Emotion Education, Large Language Models
-journalyear: 2025copyright: ccconference: ; ; $c$s: Human-centered computing Collaborative interaction
+$c$s: Human-centered computing Collaborative interaction
 Figure 1.
@@
 Emotion education plays a critical role in the developmental years of children aged 3 to 6.
-According to developmental psychology, during this period, children begin to understand various emotions — such as anger, fear, and sadness — and learn to express and regulate these feelings (bartlett1982emotional; saarni1999development; wulandari2021development).
+According to developmental psychology, during this period, children begin to understand various emotions, such as anger, fear, and sadness, and learn to express and regulate these feelings (Bartlett et al., 1982; Saarni, 1999; Wulandari, 2021).
 Recent studies have explored AI-based systems that help children recognize, express, and manage their emotions.
-However, these systems mainly target children aged 8 and above, as they rely on independent conversations with LLM-based agents — for example, expressing emotions through dialogue (seo2024chacha; tang2024emoeden; shen2025easel).
-Such designs are not well-suited for younger children aged 3 to 6, since at this stage, parents play a foundational role in shaping children’s emotional development by guiding their early emotional learning at home (fadlillah2022analysis; alegre2011parenting).
-Unlike older children, those aged 3 to 6 require parental support to engage with technology (toran2024parent; morawska2023managing) and rely on their parents for emotional security and co-regulation (corwin2012secure; brock2016interparental).
-However, providing effective emotion education can be challenging for parents, as it requires knowledge of child psychology and strong communication skills — expertise most parents lack (sanders1999triple; saarni1999development; cole1994development).
-In practice, parents often encounter difficulties like misreading children’s ambiguous emotional signals (castro2015parents), or failing to support emotional regulation through dialogue (meng2020effects).
+However, these systems primarily target children aged 8 and above, as they rely on independent conversations with LLM-based agents — for example, expressing emotions through conversations (Seo et al., 2024; Tang et al., 2024; Shen et al., 2025).
+Such designs are not well-suited for younger children aged 3 to 6, since at this stage, parents play a foundational role in shaping children’s emotional development by guiding their early emotional learning at home (Fadlillah and Fauziah, 2022; Alegre, 2011).
+Unlike older children, those aged 3 to 6 require parental support to engage with technology (Toran et al., 2024; Morawska et al., 2023) and rely on their parents for emotional security and co-regulation (Corwin, 2012; Brock and Kochanska, 2016).
+However, providing effective emotion education can be challenging for parents, as most parents lack expertise in child psychology and strong communication skills (Sanders, 1999; Saarni, 1999; Cole et al., 1994).
+In practice, parents often encounter difficulties, such as misreading children’s ambiguous emotional signals (Castro et al., 2015) or failing to support emotional regulation through conversations (Meng et al., 2020).
 To gain a more systematic understanding of these difficulties, we conducted interviews with five kindergarten teachers and five parents and identified three key challenges: ($C_{1}$) limited parental expertise in child psychology and communication techniques; ($C_{2}$) insufficient feedback on the effectiveness of emotion education for children; ($C_{3}$) a lack of comprehensive understanding of children’s emotional states.
@@
 We developed PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education for children aged 3 to 6.
-PACEE enables parents to engage in emotion education with their children through interactive adventure games.
-PACEE consists of three key modules, each addressing one of the aforementioned challenges: (1) Parent-Child Adventure Game.
-In each adventure game, parents and their children engage in five-step emotional dialogues based on Gottman’s Emotion Coaching Theory (gable1999heart).
+PACEE helps parents engage in emotion education with their children through interactive adventure games.
+PACEE consists of three key modules, each addressing one of the challenges mentioned above: (1) Parent-Child Adventure Game.
+In each adventure game, parents and their children engage in a five-stage conversation that operationalizes Gottman’s Five Steps of Emotion Coaching guideline (Gottman and DeClaire, 1997).
 The AI system provides parents with real-time guidance, such as actionable suggestions and progress tracking, thus solving $C_{1}$.
@@
 (3) Child Emotional Modeling.
-This module creates an emotional profile for each child by combining two views: long-term understanding provided by parents and AI’s analysis of parent-child dialogue data in the adventure game.
-The emotional profile is then used to offer tailored guidance to parents during the adventure game.
-After each adventure game, it also provides parents with a comparison between their own understanding of their child’s emotional states and AI observations, thereby improving parents’ understanding of their child and effectively addressing $C_{3}$.
-To evaluate PACEE’s effectiveness in supporting family emotion education, we conducted a user study with 16 pairs of parent and child.
+This module creates an emotional profile for each child from two sources: the primary source being parents’ long-term understanding, elicited through an AI interviewer, supplemented by text-based analysis of parent-child conversation data.
+The emotional profile is then used to provide tailored guidance to parents during the adventure game.
+After each adventure game, it also offers a comparison between parents’ understanding of their child’s emotions and the AI’s analysis, thereby effectively addressing $C_{3}$.
+To evaluate PACEE’s effectiveness in supporting family emotion education, we conducted a user study with 16 pairs of parents and children.
 Compared to a baseline (LLM-assisted) method, PACEE significantly fosters meaningful parent-child interactions and in-depth emotion education.
@@
 The contributions of our research are as follows:
-  • Through a formative study, we identified three key challenges in family emotion education, AI’s role in supporting this process, and principles for AI-assisted emotion education from experts’ perspectives.
+  • Through a formative study, we identified three key challenges in family emotion education, the role of AI in supporting this process, and principles for AI-assisted emotion education from the experts’ perspectives.
   • We present PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education.
-PACEE collaboratively builds children’s emotional profiles with parents and provides personalized, parent-mediated guidance.
-  • Through a user study with 16 families, we validated PACEE’s effectiveness in enhancing family emotion education and advanced emotion coaching theory in both family and LLM-assisted contexts, offering valuable insights for designing AI-supported, parent-centered family education systems.
+PACEE collaborates with parents to build children’s emotional profiles and provides personalized, parent-mediated guidance.
+  • Through a user study with 16 families, we validated PACEE’s effectiveness in enhancing family emotion education.
+Our work advanced emotion coaching theory in both family and LLM-assisted contexts, offering valuable insights for designing AI-supported, parent-centered family education systems.
 2. RELATED WORK
 2.1. Emotion Coaching Techniques and Theories
-Emotion coaching is a supportive parenting technique that helps children understand and regulate emotions and has been widely studied in parenting and child development (fu2022social; havighurst2021tuning).
-A key framework widely applied in HCI (havighurst2021tuning) is the “Five Steps of Emotion Coaching” framework by John Gottman (gottman2005five), which includes (1) attending to children’s emotional signals, (2) viewing emotional moments as opportunities for connection and learning, (3) helping children identify and articulate emotions, (4) expressing empathy and validating emotions, and (5) providing guidance and boundaries while collaboratively exploring solutions.
-This model is often complemented by principles from positive psychology (baker2017positive; fan2018emostory), such as the PERMA model (seligman2010flourish).
-PERMA model outlines five elements of well-being (Positive Emotion, Engagement, Relationships, Meaning, and Achievement) and has been applied in HCI areas such as well-being-centered design (kovich2023application), persuasive technologies (nebrida2018m), and digital mental health (chisale2022perma).
+Emotion coaching is a supportive parenting technique that helps children understand and regulate their emotions, and has been widely studied in the fields of parenting and child development (Fu et al., 2022; Havighurst and Kehoe, 2021).
+A key framework widely applied in HCI (Havighurst and Kehoe, 2021) is Gottman’s Five Steps of Emotion Coaching guideline (Gottman and DeClaire, 1997; Gottman, 2005), which includes: (1) Be aware of the child’s emotion, (2) Recognize the child’s expression of emotion as a perfect moment for intimacy and teaching, (3) Listen with empathy and validate the child’s feelings, (4) Help the child learn to label their emotions with words, and (5) Set limits when helping the child to solve problems.
+This guideline is often complemented by principles from positive psychology (Baker et al., 2017; Fan et al., 2018), such as the PERMA model (Seligman, 2010).
+The PERMA model outlines five elements of well-being (Positive Emotion, Engagement, Relationships, Meaning, and Achievement) and has been applied in HCI areas such as well-being-centered design (Kovich et al., 2023), persuasive technologies (Nebrida and Dullas, 2018), and digital mental health (Chisale and Phiri, 2022).
 Current HCI implementations of emotion coaching primarily target older children or specific clinical populations.
-Research has explored using AI to enhance emotional expression and feedback for school-aged children and adolescents (seo2024chacha; kim2018can; ibrahim2025designing).
-Meanwhile, other research focuses on providing specialized learning support for children with clinical needs, such as autism (tang2024emoeden; ghiotti2023prototyping; sharma2016promoting) or trauma (grassl2024coding).
+Research has explored using AI to enhance emotional expression and feedback for school-aged children and adolescents (Seo et al., 2024; Kim et al., 2018; Ibrahim et al., 2025).
+Meanwhile, other research focuses on providing specialized learning support for children with clinical needs, such as autism (Tang et al., 2024; Ghiotti et al., 2023; Sharma et al., 2016) or trauma (Graßl and Fraser, 2024).
 However, insufficient attention is given to children aged 3 to 6, highlighting a critical gap in HCI-driven emotional education.
-Compared to older children, 3- to 6-year-olds experience emotions more immediately, intensely, and situationally while having limited verbal and self-regulation skills (sun2024exploring).
-These traits make parental guidance central for fostering emotional competence (gottman1996parental; hightower2019exploring).
-However, many parents face barriers such as time constraints and a lack of knowledge in addressing their children’s emotional needs (ahmadpour2023understanding; quan2025parents; BROWN2022126).
-Existing tools frequently overlook these developmental traits, assuming children can engage with AI independently and failing to empower parents in their crucial coaching role (sharma2025robot).
+Compared to older children, 3- to 6-year-olds experience emotions more intensely and situationally, while having limited verbal and self-regulation skills (Sun et al., 2024).
+These traits make parental guidance central for fostering emotional competence (Gottman et al., 1996; Hightower et al., 2019).
+However, many parents face barriers such as time constraints and a lack of knowledge in addressing their children’s emotional needs (Ahmadpour et al., 2023; Quan et al., 2025; Brown et al., 2022).
+Existing tools frequently overlook these developmental traits, assuming children can engage with AI independently and failing to empower parents in their crucial coaching role (Sharma et al., 2025).
 To address this gap, we propose a parent-centered, theory-grounded approach designed to empower parents as the primary emotion coaches for their 3- to 6-year-old children.
@@
 Based on how AI participates in parent-child interactions, existing research can be categorized into three types: AI-centered, parent-centered, and parent-AI collaborative systems.
-AI-centered systems position AI as a “teaching tool” that interacts directly with children, offering guidance to help children (wang2024chaitok; wang2022don; karunaratna2024evaluation).
-For example, systems may generate personalized learning content (tang2024emoeden), support children in learning math language via story-based interaction (xu2023mathkingdom), co-create stories to boost vocabulary and feedback quality (zhang2024mathemyths), or generate context-based prompts to aid language expression (seo2024chacha).
+AI-centered systems position AI as a “teaching tool” that interacts directly with children, offering guidance to help them (Wang et al., 2024, 2022; Karunaratna et al., 2024).
+For example, systems may generate personalized learning content (Tang et al., 2024), support children in learning math language via story-based interaction (Xu et al., 2023b), co-create stories to boost vocabulary and feedback quality (Zhang et al., 2024), or generate context-based prompts to aid language expression (Seo et al., 2024).
 These systems emphasize direct AI-child interaction.
 Parent-centered systems emphasize parents as primary educators, with AI as a supportive tool that does not interact directly with the child.
-These studies often use interviews or co-design methods to explore parents’ needs and challenges (dumaru2023after; cingel2017parents; nikkhah2021helping; hiniker2016not; mazmanian2017okay; wang2021protection).
-For example, research shows how parents co-use digital media with children (yu2024parent) and welcome daily prompts for reflection (ibrahim2025designing).
-However, few systems support real-world implementation, especially regarding sustained engagement and feedback mechanisms.
-Parent-AI collaborative systems integrate the strengths of both parties through a division of labor.
-Parents can act as information providers, content configurators, facilitators, and evaluators (slovak2016scaffolding; chen2020new; yarosh2013almost).
-Meanwhile, AI is used to facilitate communication and storytelling (choi2024aacesstalk; zhang2022storybuddy), structure collaborative tasks and routines (jonas2023supporting; xiao2012supporting), and integrate smart devices into daily life (theofanopoulou2022exploring; chan2017wakey).
+These studies often use interviews or co-design methods to explore parents’ needs and challenges (Dumaru and Al-Ameen, 2023; Cingel and Piper, 2017; Nikkhah et al., 2021; Hiniker et al., 2016; Mazmanian and Lanette, 2017; Wang et al., 2021).
+For example, research shows how parents co-use digital media with children (Yu et al., 2024) and welcome daily prompts for reflection (Ibrahim et al., 2025).
+However, few systems support real-world implementation, particularly in terms of sustained engagement and feedback mechanisms.
+Parent-AI collaborative systems integrate the strengths of both parties through a division of labor, leveraging the complementary capabilities of each.
+Parents can act as information providers, content configurators, facilitators, and evaluators (Slovák et al., 2016; Chen et al., 2020; Yarosh et al., 2013).
+Meanwhile, AI is used to facilitate communication and storytelling (Choi et al., 2024; Zhang et al., 2022), structure collaborative tasks and routines (Jonas et al., 2023; Xiao and Martin, 2012), and integrate smart devices into daily life (Theofanopoulou and Slovak, 2022; Chan et al., 2017).
 These systems establish a three-way interaction model among the parent, child, and AI.
@@
 2.3. Modeling Children’s Emotional States
-Personalized instruction is crucial in modern education and children’s emotional development (bernacki2021systematic; dewey2024democracy).
+Personalized instruction is crucial in modern education and children’s emotional development (Bernacki et al., 2021; Dewey, 2024).
 Effective personalization requires modeling a child’s emotional state to enable targeted interventions.
 We examine the challenges of such modeling through two aspects: content design and assessment methods.
-Regarding content design, a classic assessment for children’s emotional development is the Test of Emotion Comprehension (TEC) (pons2004emotion; albanese2006children).
-TEC identifies nine key elements for evaluating children’s emotional competence, which have also been widely acknowledged in prior research, including: emotion recognition (bullock1985further), emotional expression (barden1980children), emotion external cause understanding (harris1989young), emotion–belief link (fonagy1997relationship), emotion reminder effect (harris1983children), emotional masking (gardner1988japanese), mixed emotional understanding (arsenio1995children), emotional morality (harter1989developmental), and emotion regulation (altshuler1989developmental).
-Regarding assessment methods, traditional approaches like parent-reported or teacher-reported questionnaires offer high ecological validity but suffer from subjectivity and inefficiency (denham1997parental; rothbart2004temperament).
-In contrast, self-report instruments like TEIQue-CF (mavroveli2008investigation) provide direct insight into children’s self-perceived emotional competence.
-However, for young children aged 3 to 6, limited cognitive and expressive abilities (wellman2004scaling) can lead to unreliable results.
-More recently, emerging AI systems using chatbots or games show promise for automated observation (santos2020therapist; nicolaidou2022gamified).
-However, these systems often struggle to interpret subtle and developmentally specific cues, especially when the data types available are limited due to privacy concerns (calvo2010affect).
+Regarding content design, a classic assessment for children’s emotional development is the Test of Emotion Comprehension (TEC) (Pons et al., 2004; Albanese et al., 2006).
+TEC identifies nine key elements for evaluating children’s emotional competence, which have also been widely acknowledged in prior research, including: emotion recognition (Bullock and Russell, 1985), emotional expression (Barden et al., 1980), emotion external cause understanding (Harris et al., 1989), emotion-belief link (Fonagy et al., 1997), emotion reminder effect (Harris, 1983), emotional masking (Gardner et al., 1988), mixed emotional understanding (Arsenio and Lover, 1995), emotional morality (Harter and Whitesell, 1989), and emotion regulation (Altshuler and Ruble, 1989).
+Regarding assessment methods, traditional approaches, such as parent-reported or teacher-reported questionnaires, offer high ecological validity but suffer from subjectivity and inefficiency (Denham et al., 1997; Rothbart et al., 2004).
+In contrast, self-report instruments like TEIQue-CF (Mavroveli et al., 2008) provide direct insight into children’s self-perceived emotional competence.
+However, for young children aged 3 to 6, limited cognitive and expressive abilities (Wellman and Liu, 2004) can lead to unreliable results.
+More recently, emerging AI systems, such as those using chatbots or games, show promise for automated observation (Santos et al., 2020; Nicolaidou et al., 2022).
+However, these systems often struggle to interpret subtle and developmentally specific cues, especially when the data types available are limited due to privacy concerns (Calvo and D’Mello, 2010).
 To address the limitations of existing assessment methods, we propose a human-AI collaborative framework that integrates the real-world understanding of parents with the multimodal analysis of AI, enabling more comprehensive emotional modeling.
@@
 To mitigate these challenges, we developed three core modules of PACEE: Parent-Child Adventure Game ($C_{1}$), Parental Feedback Report ($C_{2}$), and Child Emotional Modeling ($C_{3}$).
-PACEE follows Gottman’s Five Steps of Emotion Coaching framework, the PERMA model, and the design details advised from formative study interviews.
+PACEE operationalizes Gottman’s Five Steps of Emotion Coaching guideline (Gottman and DeClaire, 1997) and follows the design details advised from formative study interviews.
 Section 5 describes the design and implementation details.
@@
 To gain a deeper understanding of family emotion education and explore how AI can assist in this process, we conducted three formative studies: an interview session with 5 experienced kindergarten teachers (denoted as K1-K5), an interview session with 5 parents of children aged 3-6 (denoted as G1-G5), and an observational experiment focusing on emotion education involving 5 children (denoted as O1-O5).
-4.1. Procedure and Analysis
-4.1.1. Need-finding Interview with Experienced Kindergarten Teachers
-We interviewed 5 experienced kindergarten teachers (3 female, 2 male; $M=9.40$ years of experience, $SD=6.34$) recruited via social media, each receiving a \$10 stipend.
-Each interview session lasted about one hour and was conducted through video conferencing, moderated by two researchers.
-First, we interviewed the teachers on several key topics, including common negative emotions in children aged 3-6, triggers, and teachers’ strategies.
-Next, we explored the potential role of AI in supporting emotion education.
-Specifically, we introduced the design objectives and scenarios of parents providing regular emotional guidance to their children at home while the AI offers additional support throughout the process.
-Further, we co-designed an AI-assisted system based on participants’ visions and expertise, using conceptual design images to illustrate our design objectives (see figure 3(b)).
-4.1.2. Need-finding Interview with Parents of Children Aged 3-6
-We also interviewed 5 parents (4 mothers, 1 father; $M=41.0$ years old, $SD=3.16$) of children aged 3-6, recruited via social media, with \$10 stipends.
-Each hour-long video session was moderated by two researchers.
-First, we interviewed the parents about their observations of children’s emotional behaviors, challenges in guiding emotions, and understanding of emotional regulation.
-Next, our design scenarios explored the parents’ perspectives on AI-assisted tools, preferred features, and concerns.
-Together with the kindergarten teachers’ sessions, these interview sessions were audio-recorded, transcribed, and analyzed using a coding process by three HCI experts.
-4.1.3. Observational Experiment on Children’s Emotion Education
-To additionally understand the context of how parents and children communicate and interact during emotion education conversations, we observed emotion education interactions in 5 families recruited via social media, with \$10 compensation.
-The participants included four mothers, one father ($M=38.2$ years old, $SD=3.49$), and their children (4 boys, 1 girl; $M=4.8$ years old, $SD=1.17$).
-Each family received \$10 or an equivalent gift as compensation.
-Conducted in a lab over 30 minutes, parents guided their children based on a predefined script adapted from Gottman’s Five Steps of Emotion Coaching theory (gable1999heart).
-Parents followed a script and selected one of three negative emotion scenarios provided.
-Sessions were video-recorded, analyzed, and annotated by three HCI experts.
-Finally, we conducted a brief interview to ask parents about their coaching experience and difficulties.
-(a) Observational experiment settings
-(b) A conceptual design image in the interview session
-Figure 3.
-Photos taken in the formative studies, including: (a) the setup of an observational experiment, where parents guide their children following a predefined script, and (b) a conceptual design image presented during the interview session to illustrate our design objectives.
-4.2. Findings
-4.2.1. Key Challenges in Family Emotion Education
-Through interviews with 5 kindergarten teachers (K1-K5) and 5 parents (G1-G5), we identified common scenarios where children aged 3-6 struggle with negative emotions.
-These emotions are often difficult to manage and require repeated guidance for early childhood.
-The findings serve as initial data for our system’s scenario database (see Table 4 in Appendix B).
-Additionally, parental interviews revealed current practices and challenges in emotional communication and guidance.
-Combined with insights from observational experiments, we categorized the key challenges in family emotion education into three dimensions:
-$C_{1}$
-.
-Limited parental expertise.
-All five interviewed parents (5/5) acknowledged lacking professional-level skills in child emotion education, which manifests in three key aspects.
-First, parents demonstrated limited psychological and educational knowledge and struggled to adapt general theoretical guidelines to specific situations.
-For example, G2 remarked: “I feel like I don’t have access to expert advice or reliable sources, and the examples in books don’t always apply to my kid’s situation.” We also observed similar patterns in the observational study, as parents also reported that they “don’t often apply every step before”(O3) and “don’t know what to say” about certain steps (O1).
-Second, parents reported insufficient communication techniques, emphasizing the need for “concrete, easy-to-implement language and techniques” (G1, G2, G5) in daily practice.
-Third, parents experience occasional impatience during emotional communication with their children.
-G3 observed that “Sometimes my husband gets impatient when he and our child disagree, and it makes our child hesitant to open up and share feelings with him”, illustrating how parental emotional instability can negatively impact communication effectiveness.
-$C_{2}$
-.
-Insufficient feedback.
-Parents frequently lack objective, professional feedback when guiding children’s emotional development.
-As G3 noted, “When he has strong emotions, I explain and comfort him in the moment, but I can’t tell if he really understood or whether my approach was effective”.
-In the observational experiment, parents reported similar uncertainties (O4: “I just followed the guidelines step-by-step, but I’m not sure each step worked.
-Next time this happens, he might forget what we discussed today.”).
-$C_{3}$
-.
-A lack of understanding of children’s emotional states.
-Most parents reported difficulties in accurately recognizing their children’s emotional states.
-This stems partially from limited awareness of school experiences, as G4 noted “not knowing what happens at kindergarten since children don’t always share things at school.” Parents may struggle to identify the root causes of emotions (G1: “Once he cried because he was just too tired, but I initially thought it was because of something that happened at kindergarten.”).
-Additionally, parents acknowledged their subjective biases when assessing their children (G5: “We naturally think our child is the best, so this bias colors our evaluations.
-We need something objective to show how our child compares with peers.”).
-4.2.2. Roles and Design Principles of AI in Family Emotion Education
-During the interview sessions, parents expressed their vision for AI’s potential roles in emotional guidance, while experts highlighted the cognitive and emotional characteristics of young children that should guide system design.
-All five participating parents expressed a preference for maintaining their primary role in emotional guidance while leveraging AI to supplement their limitations.
-They also envisioned specific roles for AI, including serving as a dialogue facilitator, an expert resource, and a generative tool.
-4/5 parents anticipated AI’s involvement as a dialogue facilitator.
-For instance, G3 suggested that AI could dynamically adjust its level of participation based on the parent’s engagement: “When parents take the lead, AI should speak less; when parents struggle, AI can step in to enrich the conversation and reduce communication barriers.” G2 further highlighted AI’s potential to assist less experienced parents: “If parents are unsure how to respond, AI could intervene.
-It could also detect and correct misguided parental guidance in real time.
-All five parents expected AI to serve as an expert resource.
-By integrating relevant knowledge into conversations, as G4 noted: “Providing real-time pedagogical insights or case studies, such as developmental benchmarks for 5-6-year-olds.” Additionally, it could offer post-conversation reports to improve parents’ long-term educational skills and provide objective assessments of children’s behavior to counteract parental biases, with G4 adding: “I hope AI can help me observe my child’s development more objectively and comprehensively.
-Parents also proposed practical applications for AI functioning as a generative tool.
-AI could reduce communication burdens by “generating images or videos to clarify scenarios for children” or “suggesting real-time conversation strategies”.
-It could also produce post-interaction reports with performance feedback, as G5 mentioned: “It could analyze which guidance techniques I applied successfully and how well I executed them.
-The interview sessions with kindergarten teachers highlighted key cognitive characteristics of 3-6-year-old children, informing the design of AI-assisted emotion education systems.
-First, the system design must align with children’s developmental stage.
-Considering children’s tendency toward concrete and visual thinking, the system should incorporate visual aids, such as pictorial representations, to help them better grasp emotional concepts.
-Additionally, given their limited language skills, the interface should guide parents on how to use age-appropriate language during emotional coaching.
-Second, the system should provide personalized instruction to accommodate individual differences in emotional and cognitive development.
-Experts emphasized that developmental abilities vary significantly among young children.
-Thus, it is crucial to consider each child’s specific cognitive traits and emotional states to deliver tailored learning strategies.
-An AI system could effectively track behavioral patterns and provide developmental benchmarks, enabling personalized adjustments to the learning process.
-It is important to note that these design principles enable us to delve deeper into some of the main features of the core modules, which address three primary challenges in AI-assisted family emotion education.
-Since these principles are crucial design details that enhance the system’s functionality and user experience, they are not evaluated separately in subsequent studies.
-5. DESIGN AND IMPLEMENTATION OF PACEE
-5.1. User Interface and Design of System Modules
-Based on the insights from the formative studies, we designed and implemented PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education.
-PACEE consists of 3 modules—Parent-Child Adventure Game, Parental Feedback Report, Child Emotional Modeling— each addressing a key challenge in family emotion education.
-Figure 4.
-Main screens of PACEE.
-PACEE consists of 3 modules: (a) Parent-Child Adventure Game.
-Each adventure game facilitates parent-guided emotional conversations centered around a specific emotional scenario, promoting meaningful parent-child interactions.
-(b) Parental Feedback Report, which provides feedback on the adventure game and practical recommendations for parents.
-(c) Child Emotional Modeling collaboratively develops a child’s emotional profile with parental input.
-Parents contribute long-term observational data about their child, gathered through interactions with an AI interviewer.
-5.1.1. Parent-Child Adventure Game
-To address the challenge of $C_{1}$ limited parental expertise, we designed the Parent-Child Adventure Game module.
-Parents and children can choose scenarios from the system’s scenario database, which was initialized using formative study data to engage in conversations (e.g., performing on stage).
-During these dialogues, parents are guided through five steps derived from Gottman’s Five Steps of Emotion Coaching (gable1999heart).
-We made specific adaptations for children aged 3 to 6 based on formative study findings: (1) These children often have limited memory recall, so if they cannot remember relevant experiences, they can create a virtual scenario with their parents during the first step; (2) These children respond well to positive reinforcement, so the fourth step includes reflection on both negative and positive emotions, and the fifth step focuses on finding positive solutions.
-PACEE’s dialogue ultimately includes the following five stages (referred to as the FiveStepsFive\ Steps):
-$S_{1}$
-. Help the child recall an experience or describe a virtual scenario.
-$S_{2}$
-. Help the child label emotions and reason.
-$S_{3}$
-. Express empathy to children.
-$S_{4}$
-. Help the child reflect on positive and negative emotions.
-$S_{5}$
-. Help the child set boundaries and find positive solutions.
-Within this module, PACEE offers a range of support for parents.
-As a dialogue facilitator, PACEE can join the conversation when prompted by the parent, which helps reduce the communication burden.
-As an expert, PACEE provides professional advice for each of the 5 steps, considering the child’s emotional profile.
-It also periodically updates real-time advice, offering communication strategies suited to the cognitive characteristics of 3 to 6-year-old children.
-As a tool, PACEE can generate images of the current scenario for parents to show to their children, catering to children’s preference for visual thinking.
-Additionally, PACEE monitors the dialogue’s progress and stage, providing useful references for parents to decide when to progress to the next stage.
-After completing a scenario, PACEE generates rewards like stars and medals for the children, supporting their preference for encouragement-based learning.
-Our design also aligns with the PERMA model (seligman2011flourish), where PACEE encompasses Positive Emotion, Engagement, Relationship, and Achievement, contributing to a holistic approach to emotion education.
-5.1.2. Parental Feedback Report
-To address the challenge of $C_{2}$ insufficient feedback, we designed the Parental Feedback Report module.
-PPACEE evaluates parents’ performance during the Parent-Child Adventure Game, and offers actionable guidance for future interactions.
-Over extended use, parents have the opportunity to unlock achievement badges, enhancing their sense of accomplishment.
-For the evaluation of parental performance, PACEE analyzes real-time audio of parent-child conversations based on Emotional Intelligence Theory (goleman1996emotional) and Attachment Theory (howe2012attachment).
-This analysis provides insights into how parents effectively navigate the emotional education process and establish secure attachments with their children through responsive and empathetic communication.
-For future suggestions, PACEE offers detailed, personalized guidance tailored to individual differences between parents and children.
-Additionally, when formulating suggestions, PACEE considers the cognitive characteristics of children aged 3-6 to ensure the recommendations are practical and implementable.
-5.1.3. Child Emotional Modeling
-To address the challenge of $C_{3}$ a lack of understanding of children’s emotional states, we designed the Child Emotional Modeling module.
-This module combines insights from both parents and AI to overcome the limitations of singular perspectives, ensuring that the system provides personalized guidance in emotional and cognitive development.
-On the AI interviewer page, PACEE conducts semi-structured interviews with parents about their children’s emotional states, capturing long-term insights from the parents.
-Parents can respond verbally and have the option to skip questions if they find them challenging to answer.
-PACEE presents a comparative analysis of children’s emotional states on the child modeling page based on both parents’ and AI’s perceptions.
-The AI’s understanding is derived from objective, expert-level analysis of children’s speech during the Parent-Child Adventure Game.
-We draw inspiration from the Test of Emotion Comprehension (pons2004emotion; albanese2006children), and model children’s emotions across the following three dimensions:
-  • Emotional understanding, including emotion recognition, emotion elicitors, emotion-related beliefs, the impact of memory on emotions, and mixed emotions.
-  • Emotional expression, including emotional expression and emotional masking.
-  • Emotional regulation, including emotion regulation abilities and moral emotions.
-5.2. LLM-based Agent Workflow Design
-Figure 5. Agent workflow of PACEE.
-In this section, we introduce the agent workflow design of PACEE (see figure 5).
-In the Parent-Child Adventure Game, PACEE applies rapid and reflective reasoning to real-time dialogue information.
-Inspired by how human systems of “thinking fast and slow” (kahneman2011thinking) can be applied to generative agents (christakopoulou2024agents; lin2023swiftsage), we implement rapid and intuitive reasoning to generate real-time coaching support for parents, while employing reflective and deliberative reasoning steps to understand the emotional characteristics of the child progressively.
-PACEE transcribes the real-time parent-child dialogues and monitors explicit transitions across the FiveStepsFive\ Steps by mouse clicks.
-The dialogue analyzer uses this information to assess the completion level of the current dialogue stage and update the node status in the dialogue graph.
-The content generator module, combining dialogue information and the child’s emotional profile, applies rapid reasoning to provide timely support for parents: (1) phase advice updated at each stage, (2) real-time advice updated every half minute, (3) agent dialogue converted into voice mode, and (4) scenarios converted into image mode by text-to-image models.
-In addition, the content generator module integrates all dialogue information from the latest Parent-Child Adventure Game and the child’s emotional profile to generate the Parental Feedback Report for the parents.
-PACEE also applies reflective reasoning within the parent-child dialogue to gradually build the child’s emotional profile based on real-time dialogue information.
-At certain intervals (empirically set to every five dialogue turns), the cognitive analyzer is tasked with multi-step reasoning to update the child’s emotional profile (see Section 5.3).
-5.3. Modeling Children’s Emotional Profiles
-Figure 6.
-Mechanism of PACEE modeling children’s emotional profiles collaboratively with parents.
-It functions in two scenarios: (1) individual parent interviews conducted by an AI Interviewer and (2) AI-facilitated parent-child emotion dialogues monitored by a Dialog Agent.
-The Cognitive Analyzer processes inputs from both scenarios, analyzing and synthesizing emotional profile data.
-The resulting Child’s Emotional Profile includes three sub-dimensions: emotion understanding, emotion expression, and emotion regulation.
-Additionally, it compares the parent’s understanding of the child’s emotions with observations made by the AI, offering a nuanced perspective on the child’s emotional development.
-In the Child Emotional Modeling module, PACEE enables parents to contribute their long-term understanding and empathetic observations of their children in the physical world.
-To facilitate this, PACEE employs an AI interviewer to simulate a semi-structured interview (scenario 1).
-The process begins by selecting a question from a fixed list, and based on the parent’s responses, the interviewer decides whether to ask follow-up questions, adjusting the phrasing as necessary.
-If no further probing is required, the interviewer proceeds to the next question on the list.
-This list is generated with reference to the Test of Emotion Comprehension (pons2004emotion; albanese2006children).
-The AI provides detailed and objective analyses of parent-child interactions (scenario 2), addressing limitations such as human attentional constraints and subjective biases.
-Conversation data from both scenarios are analyzed by the cognitive analyzer.
-Initially, the profile extractor identifies incremental updates across the dimensions of emotional understanding, emotional expression, and emotional regulation.
-Subsequently, the profile integrator merges these incremental updates into the existing profile, minimizing redundancy.
-Furthermore, the profile integrator attempts to compare and organize entries from the AI and parents within the same dimension, presenting potential cognitive gaps to the parents.
-5.4. Implementation
-We implemented the core agent workflow in Python on a Flask server.
-The PACEE interface, written in Vue, interacts with the server by exchanging messages via WebSocket.
-Children’s profiles and dialogue graph information are stored in separate JSON files.
-For most text generation tasks, we leveraged the GPT-4o API to balance speed and output quality.
-To minimize latency in agent dialogues and real-time advice delivery, the faster GPT-4o-mini model was adopted.
-Speech-to-text processing utilized the TINGWU API, with audio sliced and compressed on the frontend before transcription.
-Text-to-speech conversion relied on Alibaba Cloud’s streaming synthesis API, while image generation integrated the Stable Diffusion XL (SDXL) model.
-6. USER EVALUATION
-We conducted a within-subject study to evaluate how PACEE supports family emotion education collaboratively with parents. We aimed to answer the following questions:
-RQ1. How does PACEE enhance family emotion education?
-RQ2. How do parents collaborate with PACEE to develop children’s emotional profiles and offer guidance to children?
-RQ3. How do parents perceive PACEE in addressing their challenges in family emotion education?
-We included a baseline condition to compare the effectiveness of PACEE in supporting family emotion education with a generic LLM chatbot.
-Each pair of parent and child participants was tasked with two common emotional scenarios from PACEE’s scenario database, one under each condition.
-6.1. Participants
-We recruited 16 pairs of parents and children from social media platforms.
-Parent participants (P1-16) included four fathers and twelve mothers, with an average age of 37.69 years ($SD=3.31$).
-Child participants (C1-16) included seven girls and nine boys, with ages ranging from 3 to 6 ($M=5.06,SD=0.66$).
-We carefully took ethical measures for participants in our user study.
-First, we obtained verbal consent from parent and child participants, confirming their willingness to participate.
-Second, we informed the participants that they had the right to withdraw from the experiment at any stage without facing any consequences.
-Third, parents signed a written consent form before the experiment, which detailed the complete process, the types of information we collected (video data, audio data, screen activity), and the anonymization methods.
-All 16 parent participants agreed to sign the consent form.
-Lastly, we conducted remote real-time manual reviews on the server during the experiment.
-These reviews included examining text and images to ensure their content appropriateness and to prevent displaying any harmful content to the children.
-6.2. Baseline Condition
-To compare PACEE with generic LLM chatbots, we designed a Baseline condition.
-In this condition, parents are provided with a brief guide to FiveStepsFive\ Steps and a standalone LLM (see Figure 7).
-They are required to guide children through the assigned scenario by following the FiveStepsFive\ Steps and can request assistance from the LLM via audio when needed.
-We do not limit the number of times parents can use the LLM during the experiment.
-For the LLM in the Baseline, we chose the Doubao model  [^1] https://www.doubao.com/chat/ to ensure a realistic and robust comparison.
-Our selection was guided by three reasons: (1) As one of the most popular free AI chatbots in Chinese families (newsDoubao1; newsDoubao2; yu2025suitability), it provides an ecologically valid baseline that participants might encounter in daily lives.
-(2) It has strong support for real-time Chinese audio conversations, which aligns with PACEE’s hands-free interaction flow.
-(3) To ensure our comparison focused on the interaction framework rather than model capability, we selected a model with competitive performance in the context of our study.
-In the Chinese context, Doubao-Seed-1.6-thinking-250715 model outperforms the ChatGPT-4o-latest model in the SuperCLUE Benchmark  [^2] https://www.superclueai.com (xu2023superclue).
-Figure 7. Illustration of Baseline condition. In the Baseline condition, participants followed a brief guide to the FiveStepsFive\ Steps and an LLM-chatbot attached to Doubao model.
-6.3. Evaluation Design
-We randomly chose two scenarios from our system’s scenario database (see Table 4), which was initialized during the formative study: (1) not winning first place (referred to as Not First), where children often encounter anger and disappointment, and (2) performing on stage (referred to as Up Stage), where children often encounter anxiety and fear.
-As illustrated in Figure 8, we used a counterbalanced design where each parent and child pair was randomly assigned to one of four groups.
-These groups were determined by the order of using the two conditions (Baseline vs.
-PACEE) and learning about the two emotional scenarios (Not First vs.
-Up Stage).
-Figure 8. Design of the evaluation study. We followed a counterbalanced design where each pair of parent and child was randomly assigned to one of four groups.
-6.4. Procedure
-Each pair of parent and child went through a 2-hour study in our lab (see figure 8). We provided a touch-screen computer for each family and recorded the entire evaluation process.
-Introduction session.
-Before each emotion education session, we introduced the parent participants to the designated system.
-We demonstrated how to interact with each system and provided tips to avoid misuse.
-We also provided parents with a practice session using several example scenarios to familiarize them with the procedure.
-In addition, we informed parents that the session required joint participation from both parents and children, and it was not intended for the child to engage solely with the AI.
-Emotion Education Session Using PACEE or Baseline.
-In each emotion education session, we asked parents to freely use the system to learn about the specified emotional scenario with their children.
-Considering children aged 3 to 6 may be shy around strangers, the researcher left the room to allow parents and children to interact freely.
-We monitored the session via live video feed and informed parents they could contact us anytime if they encountered any issues.
-Post-Questionnaire, Parent Interview, and Child Drawing.
-Parent participants completed a post-questionnaire to evaluate their educational experience after completing all education sessions.
-Then, one researcher interviewed the parents about their user experience and perceptions, while another researcher asked the child participants to use emoji stickers to represent their feelings during conversations, the scenarios discussed, and the image of PACEE.
-Finally, the children drew a picture to depict the scenarios discussed with PACEE.
-(a) Evaluation experiment settings using PACEE.
-(b) Children create drawings after the experiment.
-Figure 9.
-Photos taken in the evaluation studies, including: (a) the setup of an evaluation experiment session, where parents and child engage in emotion education using PACEE, and (b) a drawing created by a child after the experiment session, depicting the Not First scenario discussed with PACEE.
-6.5. Data Analysis
-For qualitative analysis, we transcribed all parent interviews.
-Five HCI researchers with qualitative coding experience conducted a thematic analysis (braun2006using; terry2017thematic; braun2019reflecting).
-The process began with a data familiarization phase, where all researchers went through the transcripts to gain a comprehensive understanding of the content.
-The leading author then generated initial code themes, which were reviewed and refined with feedback from another researcher.
-The team then collaboratively developed a shared codebook by coding data from one participant (P1), with discussions for consensus.
-The remaining data were then independently coded, and disagreements were resolved through discussion.
-Finally, the team manually and iteratively clustered the codes to identify recurring and meaningful patterns.
-Details of the statistical analysis are presented in Appendix D.
-7. FINDINGS
-7.1. RQ1. How Does PACEE Enhance Family Emotion Education?
-7.1.1. Increased Parent-Child Interaction
-Figure 10.
-User interaction data from the PACEE and Baseline systems during the teaching sessions.
-(a-d) represent data on four dimensions: interaction duration, number of dialogue turns, word count, and the count of emotion-related words.
-(e) Reflect the interaction duration in Five Steps.
-The x-axis indicates participant identity for (a-d) and teaching steps for (e), with error bars representing standard deviation.
-Significance analysis: (a, e) uses the Paired Sample t-test, while (b-d) are based on the Wilcoxon signed-rank test.
-Significance results are marked with asterisks (***: p ¡ .001, **: p ¡ .01, *: p ¡ .05).
-To compare user engagement and interaction patterns between the PACEE and Baseline systems, we conducted descriptive statistics and significance testing across four key dimensions: interaction duration, number of dialogue turns, word count, and the count of emotion-related words.
-The results show that PACEE consistently outperforms the Baseline, with statistically significant advantages across all four areas.
-For PACEE, the average interaction duration was 892.37 seconds ($SD=310.96$), with an average of 60 dialogue turns ($SD=46.979$).
-In contrast, the average interaction duration for the Baseline system was 560.17 seconds ($SD=194.61$), with 38 average dialogue turns ($SD=25.3$).
-Considering that both systems were matched in workload and participants were given flexibility in session completion, the results suggest that PACEE leads to higher user engagement, with significant differences in interaction duration ($p<0.01$) and dialogue turns ($p<0.05$).
-This difference was evident in both parents and children.
-Parents using PACEE produced an average of 1,597 words ($SD=678.56$), compared to 994 words ($SD=468.36$) with the Baseline ($p<0.05$).
-Children expressed an average of 370 words ($SD=240.43$) with PACEE and 165 words ($SD=102.25$) with the Baseline ($p<0.001$), as shown in Figure 10.
-To further examine conversation quality alongside user engagement, we compared the number of emotion-related words used by parents and children.
-Under PACEE, parents used an average of 111 emotion-related words (53 positive, 58 negative), while children used 19 (8 positive, 11 negative).
-In contrast, under the Baseline, parents used an average of 70 emotion-related words (31 positive, 39 negative), and children used 10 (5 positive, 5 negative).
-Both parents ($p<0.05$) and children ($p<0.01$) are more willing to express their emotions using PACEE.
-These findings suggest that PACEE fosters more meaningful emotional exchanges, which are essential for supporting children’s early development of language and emotional competence (brushe2025causal; reschke2024links).
-To gain a more intuitive understanding of participants’ engagement across the FiveStepsFive\ Steps of the teaching session, we calculated the average time spent at each step for both systems, as shown in Figure 10 (e).
-Participants’ engagement was generally higher in PACEE at all five steps, particularly in $S_{1}$ ($p<0.05$) and $S_{4}$ ($p<0.05$).
-Considering the limited memory capacity of preschool children (scarf2013have; roman2014assessment), the higher participation time in $S_{1}$ suggests that PACEE effectively supported participants in recalling scene details (PACEE $M=183.38$ vsvs Baseline $M=123.50$; $p<0.05$).
-Meanwhile, in $S_{4}$, PACEE participants engaged in significantly more reflection on positive and negative emotions (PACEE $M=163.86$ vsvs Baseline $M=85.33$; $p<0.05$).
-7.1.2. Sustained Participation Over Time
-To investigate the potential influence of the novelty effect, we conducted a temporal analysis by dividing each interaction session within the PACEE system into two equal halves by time (see figure 11).
-Our analysis revealed no statistical differences between the first and second halves of the session regarding user dialogue turns, number of expression words, or number of emotion-related words over time.
-This stability suggests that the high level of user participation is a consistent feature within PACEE.
-Figure 11.
-Temporal analysis of user interaction data within the PACEE system, comparing interaction during the first and second temporal halves of the PACEE teaching sessions across three metrics: (a) number of dialogue turns, (b) user word count, and (c) the count of emotion-related words.
-Error bars represent standard deviation.
-No significant differences were found between the two periods based on Wilcoxon signed-rank tests.
-7.1.3. Improved Parental Experience and Children’s Engagement Respectively
-Figure 12.
-Post-questionnaire data on parental experience, including metrics from the NASA-TLX and the Parenting Sense of Competence Scale.
-We applied the Wilcoxon signed-rank tests to analyze the significance of each metric.
-Statistically significant ratings are marked with an asterisk (*).
-Figure 12 shows parent participants’ responses to the post-questionnaire on their parental experience.
-Compared to Baseline, PACEE improves parents’ immediate interaction experience, with higher satisfaction ($p<0.05$), reduced effort ($p<0.01$), and less frustration ($p<0.05$).
-However, intrinsic psychological factors (e.g., parenting confidence or effort with the child) show no significant change during the study, likely because these factors need long-term exposure and reinforcement to yield measurable differences (bandura1997self; sandler2011long).
-In addition, we analyzed children’s post-study drawings to explore their perceptions of PACEE.
-Thirteen of the 16 children were able to accurately reconstruct key interaction scenes, such as the stage, parents, and audience, demonstrating strong memory retention.
-Their illustrations often conveyed positive emotions through visual cues, including smiley faces, hand-holding gestures, and a dominance of warm color palettes.
-For example, C1’s drawing depicted the stage with smiling characters rendered in warm hues.
-Similarly, C5’s drawing was colorful, featuring balloons, smiles, and hand-holding gestures.
-Beyond scene recollection, the drawings revealed nuanced perceptions of PACEE itself.
-Thirteen of the 16 children portrayed PACEE as a friendly and supportive figure.
-Common representations included anthropomorphic characters with open arms and smiling faces (C1, C8, C12, C16), friendly cartoon animals like cats with smiles (C5, C6, C10), or simplified smiley-face icons (C3, C7, C13).
-Personalized elements, such as C9’s drawing of PACEE resting on the child’s own head, further implied emotional attachment.
-In contrast, three of the 16 children (C2, C4, C11) depicted PACEE as a machine-like figure with neutral or absent facial expressions.
-These children also had difficulty conveying scene details and emotions in their drawings.
-These findings suggest that perceiving PACEE as impersonal may correlate with lower engagement.
-Figure 13. Drawings created by children (C1, C5, C10) after the experiment.
-7.2. RQ2. How Do Parents Collaborate with PACEE to Develop Children’s Emotional Profiles and Offer Guidance to Children?
-7.2.1. Parents and PACEE Contribute Different Factors in Construction of Children’s Emotional Profiles
-According to experts’ annotations, parents and PACEE play distinct roles in constructing children’s emotional profiles.
-Regarding data sources, PACEE derives its data from the content of parent-child communication.
-At the same time, parents tend to provide information based on firsthand observations of events, third-party feedback, school progress, and comparisons to their children’s childhood experiences.
-Furthermore, Table 5 (see Appendix E) illustrates the cognition-level themes in children’s emotional profiles identified in our evaluation studies.
-We found that parents prefer to extract insights from various life events, focusing on experiential knowledge of their children’s habits.
-In contrast, PACEE evaluates developmental capabilities by analyzing parent-child interactions, emphasizing expert assessments of abilities and developmental characteristics.
-7.2.2. PACEE Provides New Insights on Children’s Profile for Parental Reflection
-During the interviews, parents acknowledged that PACEE offers a more objective assessment of children’s emotional capacities.
-Some parents realized they had underestimated their child’s emotional understanding after reviewing the AI’s evaluation.
-As P2 noted, “Sometimes I didn’t realize how well my child could understand emotions, but the AI, speaking with the tone of an expert, showed me that he is truly growing.” On the other hand, some parents discovered they had overestimated their child’s abilities.
-As P4 shared, “Parents often view their children through a biased lens, believing they’re performing well.
-However, the AI objectively highlighted certain issues, and despite its evaluation diverging from my own perspective, I found the AI more rational and unbiased.” Additionally, AI can help parents understand the underlying reasons behind their child’s emotional responses.
-As P8 explained, “When my child participated in an English competition, he was very reluctant to go on stage.
-I assumed he was nervous and anxious because he was afraid of the stage, but the AI suggested that the anxiety might have stemmed from the absence of his mother’s support.
-In a broader view, parents express opinions on AI’s advantages in modeling children’s emotional profiles, including more comprehensive summaries, more objective perspectives, and more educational insights.
-Also, AI has potential disadvantages such as lacking long-term data, lacking specificity, and making incorrect inferences.
-These factors can lead to cognitive differences between parents and AI regarding children’s emotional states.
-We further inquired how parents perceive these differences.
-14/16 parents considered these differences between humans and AI as valuable new insights.
-They acknowledged that the individual suggestions made by AI are derived from broader patterns observed in children of similar age, allowing for more long-term and holistic reflections on a child’s emotional tendencies.
-As P16 explained, “I can only see the emotional issues my child is currently experiencing and how he copes with them, I have no experience dealing with newly emerging emotions.” In contrast, AI offers a more rational point of view that can help parents break free from fixed thought patterns.
-P8 remarked, “Because our communication with the child is usually smooth, we’ve fallen into a comfortable but fixed mode of interaction.
-While this is pleasant for daily life, it might hinder development in certain areas.
-The involvement of AI could reduce this limitation.
-However, 2/16 parents expressed concerns about the accuracy of AI-generated insights, viewing discrepancies from their own understanding as potential system errors.
-P15 commented, “The AI detected a sense of compromise in my child during the conversation, but I doubt he has the capacity to express such emotion yet.
-The term ’compromise’ seems imprecise.” Some parents also believed that AI could not detect when a child was lying, whereas their familiarity with their child enabled them to make more accurate judgments.
-As P11 noted, “AI can’t tell whether what the child says is sincere or not —it records everything regardless of accuracy.
-So if the child lies, it might offer the parent an incorrect solution.
-7.2.3. Parents Appreciate Open and Positive Advice from PACEE
-Parent participants received a total of 513 updates of real-time advice during the studies.
-Two experienced HCI researchers reviewed system logs and video recordings to determine whether the parents adopted the suggestions.
-They then iteratively developed classification criteria for the accepted advice, selecting the most prominent type in cases where advice fell into multiple categories.
-Table 6 (see Appendix E) summarizes the types of accepted real-time advice.
-We found that the types of advice most frequently adopted by parents were collaborative problem-solving and open-ended questioning.
-Parents also mentioned the usefulness of open-ended questioning in the interview, as P2 stated, “Real-time advice helps me realize some issues I hadn’t thought about, such as prompting my child to provide more details.
-Sometimes I think it’s okay when a child says one word, but the advice reminds me to ask for more details.” Furthermore, advice on scenario simulation, concrete suggestions, positive encouragement, and empathy and acceptance also helped some parents.
-7.2.4. Parents Accept PACEE in Collaborative Guidance, but with Concerns
-Parents shared their perspectives on integrating PACEE into the emotional guidance process during the interviews.
-Participants consistently viewed themselves as the primary agents in their children’s education.
-Nine obviously reported making the greatest contribution during the experimental sessions, whereas only P10 regarded the PACEE as having had a larger impact.
-Furthermore, P5, P8, and P13 explained that their main role was to evaluate and select which of the PACEE recommendations to implement, while P6, P9, P10, and P13 credited the role of the system in providing the specialized professional expertise necessary to complete the tasks.
-In addition, parents expressed a positive attitude toward incorporating PACEE into parent–child dialogues.
-All participants accepted PACEE involvement, though their preferred levels and modes of engagement varied.
-Some parents demonstrated a high degree of acceptance (P5: “My kid really like it(to receive emotional coaching from PACEE).
-I think I am very receptive.”), whereas others felt that PACEE was not strictly necessary, especially in familiar emotional scenarios (P16: “We often discuss this topic so my child no longer need emotional guidance in this situation.”).
-Parents recognized the system’s capacity to deliver valuable, professional, and targeted recommendations.
-For example, P8 highlighted that such suggestions made parent–child communication more effective, “He felt I understood him better, so he answered every question loudly and confidently.
-When discussing future interactions with PACEE, participants articulated clear expectations about PACEE’s role in parent–child dialogues, though they differed markedly on its ideal interlocutor.
-Some parents (P2, P16) envisioned AI forging a peer-like relationship with their child, simultaneously acting as an educator.
-Others (P8) preferred that PACEE interact only with the parent, acting as a co-educator, and avoid direct communication with the child.
-A third group of participants (P10, P13, and P15) believed that AI should adopt a neutral stance as an observer—providing limited, objective, data-driven suggestions to parents and offering age-appropriate guidance to children, to support emotional development, rather than engaging directly with either parent or child.
-For instance, P10 argued that “It can serve as a mediator (in parent–child emotional education).” P13 and P15 prefer PACEE to offer suggestions from a third-party perspective for both parents and children.
-Crucially, parents asserted firm boundaries around AI intervention.
-Some parents rejected features that impose targets on children (P4: “For example, setting goals for children… I think it is up to parents to do it.”), and they were adamant that PACEE should not encroach on sensitive family matters (P1: “Don’t press further… for example, on topics that parents find taboo and children find sensitive.
-”), or strictly require compliance with its advice (P11: “If it(PACEE) tells me I have to do something, I definitely won’t accept it.”).
-Moreover, participants cautioned that AI should not replace emotional intimacy and empathetic support intrinsic to parent-child bonding and the cultivation of self-confidence, such as “(I hope she can) build her confidence by communicating with others.”(P2), “I prefer she get emotional companionship from people…”(P5).
-7.3. RQ3. How Do Parents Perceive PACEE in Addressing Their Challenges in Family Emotion Education?
-The subjective ratings provided by parents indicate that PACEE significantly outperformed Baseline in tackling the three main challenges of family emotion education (see figure 14), with statistical significance observed in each case ($p<0.01$).
-Figure 14.
-Post-questionnaire data focusing on the extent to which PACEE or Baseline address the three key challenges.
-We applied the Wilcoxon signed-rank tests to analyze the significance of each metric.
-Statistically significant ratings are marked with an asterisk (*).
-Interviews with parents further confirmed that PACEE plays a multifaceted role in addressing the three main challenges of family emotion education.
-7.3.1. Enhancing Parental Expertise and Communication Confidence
-In terms of $C_{1}$ limited parental expertise, the Parent-Child Adventure Game module significantly alleviated parents’ difficulties in situations where they “don’t know what to say” or “don’t know how to guide.” During the interviews, parents acknowledged that they find it difficult to perform some steps alone, including recalling an experience ($n=4$), labeling emotions ($n=8$), expressing empathy ($n=8$), reflecting on emotions ($n=5$), and finding positive solutions ($n=6$).
-The system’s dialogue suggestions were praised for mitigating these expertise gaps.
-First, it provides a structured framework for emotion communication.
-Two parents specifically highlighted how the AI facilitated guiding emotional conversations with children in a gradual, structured manner.
-As P8 noted: “The systematic process — understand the child first, then ask about their feelings, then summarize, encourage, and finally solve the problem.
-This step-by-step guidance is great, creating remarkably smoother interactions.” Second, several parents appreciated how AI helped them adopt their child’s perspective.
-P10 explained: “Placing us parents in a conversational setting that feels more like talking to a friend actually encourages us to respond more actively and calmly.
-It’s because of this setup that I’m more willing to take the child’s perspective, understand them, and help them manage their emotions.” Lastly, the dialogue suggestions offer ideas to steer the dialogue stream effectively.
-P16 stated, “AI’s reminders can offer us more ideas.
-During the reflection stage, both the child and I are thinking, and it’s easy to get caught up in the moment.
-The AI’s prompts provide a clearer perspective, which helps make the conversation more in-depth.
-Beyond providing suggestions, the module actively enhanced parent-child communication and boosted parental confidence during real interactions.
-Most parents reported increased initiative and comfort in emotional conversations after using the Parent-Child Adventure Game module.
-Ten participants confirmed that the system’s dialogue suggestions provided concrete, child-friendly advice for expressing emotions.
-As P5 explained, “It can serve as a reference tool.
-For example, if you have an idea but don’t know how to express it, the system helps you organize your thoughts and express them more clearly and reasonably.” Additionally, two parents mentioned that AI-provided prompts served as useful entry points, promoting more open communication between children and parents.
-P1 explained: “As a father, I don’t talk with my child that often.
-For example, instead of saying, ‘Do you remember your last kindergarten performance?’ it feels more natural to ask, ‘How was your last performance?’ So the detailed suggestions provided by the system are quite helpful for me as a reference.” Some parents also noted that the module facilitated more reciprocal exchanges, helping to enhance parent-child relationships.
-While the benefits of AI assistance were frequently highlighted, some parents expressed concerns.
-A few parents felt that frequently checking the AI’s suggestions could slightly disrupt the conversational flow.
-One parent worried that the AI’s scripted suggestions might make children feel nervous or overly reliant.
-Another parent emphasized that AI should act as an aid rather than lead in communication, suggesting it should provide direction rather than take over the dialogue entirely.
-Furthermore, some saw the structured nature of AI prompts as potentially limiting the child’s freedom to express diverse experiences.
-There was also concern that excessive reliance on AI-generated expressions might hinder parents’ own communicative abilities.
-7.3.2. Providing Trackable Feedback and Actionable Future Strategies
-In terms of $C_{2}$ insufficient feedback, the Parental Feedback Report module effectively addressed pain points such as “lack of timely feedback” and “invisible outcomes”.
-Eight parents highlighted how AI-generated analyses shed light on previously unnoticed shortcomings in their parenting approaches, which prompted them to engage in significant self-reflection.
-For instance, P6 shared, “The system reminded me to avoid interrupting.
-I hadn’t realized I was doing that, but it made me see that I still need to be more patient in communication.” Furthermore, the module provided a structured review by organizing conversation highlights, which five parents found particularly beneficial in bridging feedback gaps and making abstract parenting efforts trackable.
-This transformation into visible progress was crucial for parents to assess their parenting strategies.
-Beyond reviews on past interactions, the module also provides actionable future strategies.
-Eight parents appreciated the evidence-based nature of the strategies provided, while three were pleased with the introduction of new, beneficial perspectives.
-P12 noted, “For example, when a child loses control, besides suggesting enjoyable activities, the system also mentions deep breathing and meditation.
-I used to think those were only for adults, but this made me realize they might work for kids too.”.
-Also, P1 noted: “AI can indeed offer suggestions we wouldn’t normally think of, like encouraging the child to observe how others handle situations or simulate a scenario at home — these are quite practical.” Parents praised that these suggestions provide a clear foundation for future communication.
-The module also incorporated practical features that increased its utility.
-The “review and scoring” design was highlighted by four parents as a useful tool for evaluating the effectiveness of their interactions, as it allowed them to pinpoint areas needing improvement.
-P15 noted: “I think the scoring in the review session is useful.
-It helps me understand which parts of my communication need improvement.
-Everyone is busy, so this kind of tool is really practical.”.
-Additionally, motivational elements such as the “badge achievement system” were noted to enhance parental engagement and motivation.
-Despite these positive aspects, some concerns remained.
-A few parents pointed out that a single session was not enough to assess real outcomes, and some found the suggestions too generic, which points to the need for the AI to better adapt to specific parental contexts.
-7.3.3. Improving Parental Understanding and Grasp of Child Developmental Trajectories
-In terms of $C_{3}$, a lack of understanding of children’s emotional states, the Child Emotional Modeling module helped parents break through cognitive blind spots by leveraging data-driven insights to reveal their child’s emotional traits and behavioral patterns.
-Twelve parents reported that the system’s scientifically grounded feedback provided “new knowledge and dimensions”, helping them overcome cognitive blind spots regarding their child’s emotions and behavior.
-As P12 explained: “The system’s questions prompted me to think about aspects I hadn’t paid attention to before.
-The conclusions helped me understand my child more systematically and comprehensively.
-It was very helpful.” Thirteen parents further emphasized how AI generated “new angles” or “reference dimensions”, saying that the system helped them move beyond existing subjective impressions to enable more objective and reflective observations.
-This objective analysis was particularly valued in understanding the causes of children’s emotions, especially in labeling children’s emotions ( $S_{2}$).
-P6 explained: “AI can analyze the causes of a child’s emotions more objectively, whereas we as parents often bring in our own biases or emotions, making it hard to view the situation in a scientific and objective way.
-Parents also discussed the alignment between their views and the system’s analyses.
-Some parents said the system’s views were highly consistent with their own, enhancing their trust.
-Others felt that when discrepancies arose, the system helped them see where their perceptions diverged — even serving as a useful reference for resolving differences within the family.
-Beyond individual insights, the module also contextualized children’s behavior within broader developmental frameworks.
-Four parents mentioned that the AI’s use of large-scale data gave them their first systematic understanding of “common behaviors at this age,” which helped them better assess their child’s traits.
-Additionally, several parents valued the forward-looking nature of the AI’s suggestions, which strengthened their overall grasp of their child’s developmental trajectory.
-Despite these advantages, several limitations were noted concerning the AI’s understanding and accuracy.
-Seven parents were concerned that the limited interaction time during the trial meant the generated profiles might not be comprehensive.
-Four parents noted that the system cannot verify the truthfulness of the information entered, which may lead to biased analysis.
-P1 noted: “AI can only analyze based on what the child says, but children don’t always express their true feelings.
-Sometimes they even lie—for example, they might be very nervous but say they’re fine.
-This affects the AI’s judgment.” Two parents expressed discomfort with the scoring system and said they didn’t want to receive low scores from the system.
-8. DISCUSSION
-In our study, we designed PACEE from a parent-centered view.
-We based this approach on recognizing that the role of parents is essential in early childhood education and that the ability to interact with AI varies among children aged 3 to 6 years old.
-PACEE leverages multiple generative AI technologies (LLMs, text-to-image, text-to-speech) to address challenges for parents ( $C_{1}$- $C_{3}$).
-Our evaluation studies show that parents perceive PACEE to have successfully addressed these challenges, reducing their effort and boosting their satisfaction.
-Here, we further discuss the impact of generative AI on early childhood family education and the impact of PACEE on Gottman’s Five Steps of Emotion Coaching framework.
-We then propose broader design implications for parent-centered, AI-supported education systems.
-8.1. PACEE Enhanced Children’s Learning through Improving Parents’ Performance
-Prior work on AI-assisted child education has explored AI systems directly coaching the child (seo2024chacha; tang2024emoeden), AI systems for intervening with parents (entenberg2023ai), and systems that facilitate parent-child interaction simultaneously (zhang2022storybuddy; theofanopoulou2022exploring).
-Similarly, we identify the need for generative AI support in our work.
-However, our context is different in that children aged 3 to 6 often face challenges in interacting independently with AI (sun2024exploring; bush2025bridging; escobar2022guidelines; thill2022modelling), and early childhood emotion coaching requires more substantial support from parents (theofanopoulou2022exploring; zhu2025maternal; lee2023effects).
-Also, prior work has expressed concerns that AI alone may not be able to adequately respond to young children’s emotional needs (kurian2025ai; kurian2024no).
-Thus, we propose a parent-centered, AI-supported system for parent-child interaction.
-This parent-centered interaction pattern, where AI primarily assists parents rather than directly interacting with children, was preliminarily explored in therapeutic domains, such as for children with language delay (hwang2014talkbetter).
-We further extend this pattern to the domain of early-childhood education.
-In our work, PACEE enhances children’s learning outcomes mainly by improving parents’ performance.
-By positioning itself as a strategic partner rather than merely a tool, PACEE has fostered more effective communication between parents and their children, including higher interaction engagement and more in-depth emotional dialogues.
-Consequently, child participants exhibited higher engagement levels and used more emotion-related words during the teaching sessions.
-However, the potential impact of generative AI directly interacting with children aged 3-6 in emotional education needs further exploration.
-In our study, we emphasized parents’ participation and informed parents that PACEE was not intended for the child to engage solely with the AI.
-This decision is due to findings from our formative study, where all five participating parents preferred to maintain their primary role in emotional guidance.
-Regarding the extent of AI directly interacting with the children, parents showed varying acceptance levels in the interviews.
-We also found that children’s AI literacy significantly varied based on parents’ attitudes.
-Therefore, the future design of emotion education systems needs to customize AI interaction levels based on parental acceptance and the child’s AI literacy.
-Also, longitudinal studies are required in order to assess the impact of varying degrees of AI interaction on children’s emotional development over time.
-Under our parent-centered interaction pattern, one challenge is to help parents better understand their children.
-To address this, we propose a framework for AI-parent collaborative modeling of children’s emotional profiles.
-In our framework, parents tend to contribute experiential knowledge of their children’s habits, while AI tends to evaluate developmental capabilities by analyzing parent-child interactions.
-This approach can also be extended to other areas of early childhood education, building upon prior work on personalized education systems.
-Prior systems either allow parents to manually set basic information and preferences (zhang2022storybuddy; fang2024edulive; chen2025characterizing) or automatically assess children’s learning progress (lee2024open).
-By combining the strengths of parental expertise and AI-driven insights, our framework fosters a dynamic process of mutual learning.
-Parents and AI can collaboratively refine their understanding of the child’s needs, enabling more effective and personalized educational strategies.
-8.2. The Impact of PACEE on Five Steps Framework
-Here, we analyze the impact of our approach on the FiveStepsFive\ Steps framework.
-To begin with, considering the developmental characteristics of 3-6-year-old children, PACEE adopts two variations: (1) the involvement of parents in recalling experiences or creating a virtual scenario, considering the limited memory and expression capabilities of 3-6-year-olds; (2) the integration of the PERMA model of positive psychology into the Five Steps, including reflection on positive emotions and finding positive solutions.
-In our evaluation studies, we found that the interaction patterns and dialogue analysis align with our adjustments.
-We further perform an analysis of the implications of integrating generative AI into the FiveStepsFive\ Steps framework.
-Generally, PACEE supports each of the five steps from the parent’s perspective, consistent with prior work exploring design interventions to support emotional coaching based on this theory.
-However, we also highlight that the design needs further consideration in the unique developmental context of children aged 3-6: (1) Both parents and children should avoid overdependence on AI to preserve skills like free expression, problem-solving, and empathy.
-(2) Conflicts between parental opinions and AI recommendations must be resolved thoughtfully.
-(3) Advanced technical solutions are required to tailor AI advice more accurately to real-life individual contexts.
-Here is a detailed analysis of each step:
-$S_{1}$
-Help the child recall an experience or describe a virtual scenario.
-Previous work has designed chatbots to encourage children to share personal experiences and associated emotions (seo2024chacha; grove2021co).
-In contrast, our approach focuses on supporting parents who directly encourage their children to recall past experiences.
-Given the limited memory recall of 3- to 6-year-olds (xuelian2018developments), it’s more effective for parents to assist in recalling and narrating experiences than to let children do it alone.
-From the parents’ perspective, PACEE was seen as a valuable tool, offering helpful prompts and contextual cues that facilitated memory recall and fostered more open communication between parents and children.
-Additionally, the parent-child interaction duration for this step is significantly increased by PACEE, which, in turn, helps strengthen the parent-child bond.
-However, several limitations were also identified.
-For instance, some parents reported that the scenarios provided by AI lacked authenticity or real-life relevance.
-Additionally, the structured nature of AI-generated prompts was seen by some as potentially restricting the child’s ability to freely express diverse experiences.
-These findings suggest the need for future work to focus on improving the flexibility of AI-generated prompts, ensuring they align more closely with real-life contexts and encourage creative expression.
-$S_{2}$
-Help the child label emotions and reason.
-Previous studies have designed systems that collect data from children to assess their emotions, such as storytelling chatbots (santos2020therapist) or audio-visual recognition models (nojavanasghari2016emoreact).
-Similarly, PACEE leverages LLMs to analyze real-time parent-child conversations, aiming to identify the emotions underlying children’s expressions.
-Differently, PACEE also provides precise vocabulary and child-friendly explanations for parents during this process —a feature participants found particularly valuable.
-Some parents also noted that AI provided objective insights into the causes of emotions, which facilitated deeper discussions.
-However, our findings revealed that AI interpretations of children’s emotions occasionally contradicted parents’ own understandings.
-This highlights the need for future research to develop a third-party mediator between parents and AI, which could help bridge differing perspectives and promote collaborative decision-making.
-Additionally, there were concerns that excessive reliance on AI-generated expressions might hinder parents’ own communication abilities, which was also noted in prior studies on parental dependence on AI (szondy2025artificial).
-Future work should explore strategies to balance AI assistance with cultivating parents’ natural communicative abilities, such as offering optional training modules or limiting AI intervention in certain scenarios.
-$S_{3}$
-Express empathy to children.
-This step revealed a firm reliance on AI support because parents often neglect this step in home education.
-In the evaluation studies, this step had the shortest duration.
-Prior work has tried to enhance this step by encouraging parents to adopt the child’s perspective during moments of conflict (yoo2019mom).
-Similarly, in our work, PACEE was valued for providing reference expressions for parents that made children feel genuinely understood and assisting parents in adopting the child’s perspective.
-Prior work has argued that AI tools should assist parents rather than entirely replace their role in areas such as storytelling (sun2024exploring).
-Similarly, our participants cautioned that AI cannot replace the emotional intimacy intrinsic to parent-child bonding.
-Expanding on this, we apply these insights to the context of emotion coaching, which involves navigating sensitive emotional exchanges.
-We emphasize the importance of parents providing additional support through emotional expressions, such as using body language (e.g., hugging).
-Future research should also investigate how varying levels of support impact children’s emotional development.
-$S_{4}$
-Help the child reflect on positive and negative emotions.
-PACEE significantly increased the interaction duration for this step.
-PACEE was regarded as helpful in supporting this stage by encouraging children to consider alternative outcomes and assisting parents in presenting new viewpoints.
-Several parents also credited PACEE with offering scientifically grounded guidance and simulating the child’s perspectives, which boosted their confidence in leading emotional reflections.
-This is similar to prior work regarding in-situ suggestions to help parents explore the various response strategies (choi2024aacesstalk).
-While PACEE demonstrated significant benefits, parents also expressed the desire for AI systems to provide more personalized guidance through repeated use and long-term data collection.
-$S_{5}$
-Help the child set boundaries and find positive solutions.
-Research has shown that collaborative problem-solving can significantly reduce children’s behavioral challenges and ease parenting stress (heath2020collaborative).
-Building on this, we adapted the approach to an emotion-sensitive context with AI support to fill cognitive gaps for parents.
-PACEE was praised for its clear and structured suggestions, which made setting boundaries and developing solutions more practical and achievable.
-Parents also noted that the AI broadened their perspectives by offering fresh approaches and precise language for communicating behavioral expectations.
-This step received the most positive feedback regarding AI assistance, with participants spending the longest time on it during the evaluation study.
-However, some areas for improvement were noted.
-Parents expressed a need for more personalized, context-specific advice to better fit their family dynamics.
-They also opposed features that imposed rigid goals on children, feeling these could undermine the collaborative nature of the process.
-Lastly, they expressed concerns about children becoming overly reliant on AI for finding solutions, echoing previous research on children’s dependency on AI (solyst2024children).
-To address these concerns, future AI systems should promote children’s independent thinking by encouraging open discussions and fostering creative problem-solving.
-8.3. Design Considerations for Parent-centered, AI-supported Education Systems
-Parent interviews also revealed several challenges that need to be addressed.
-Combining our findings with prior work, we propose design considerations for future parent-centered, AI-supported education systems.
-First, customize AI’s aiding modules based on parents’ profiles, in addition to their children’s.
-Some recent work has integrated children’s profiles or preferences to facilitate a more personalized coaching experience (clabaugh2019long; tang2024emoeden; zhang2022storybuddy).
-This study supplements prior research and demonstrates the necessity of customizing AI systems based on parental profiles.
-During our interviews, we found that parents expressed significantly varied opinions about the “ideal support” from PACEE.
-Concerning content difficulty, some experienced parents criticized that AI’s guidance overlapped heavily with their own approaches, while a few parents mentioned that some suggestions were too complex to understand.
-Concerning the frequency of suggestions, some parents felt that frequently checking the AI’s advice could slightly disrupt the conversational flow.
-In contrast, others believed that PACEE could offer more real-time interactions.
-These findings highlight that parent-centered educational systems should accommodate such differences in parents’ needs.
-For example, parents’ preferences regarding the appropriateness of early-childhood digital play, which may influence the quality of AI-assisted activities at home (isikoglu2019parents), should be collected beforehand.
-Furthermore, various parental inputs can be captured during parent-child interactions, besides audio data from dialogues.
-These inputs may include whether parental feedback during interactions is positive or negative (robinson1981dyadic), behavioral differences between fathers and mothers (cutler2020fathers), and parents’ emotional regulation abilities (bariola2011child).
-By incorporating diverse data sources, the system can better tailor its functions to the specific needs of each family.
-Second, avoid excessive reliance on AI capabilities and promote transparent and negotiable collaboration.
-In the evaluation studies, parents expressed worries that the system cannot verify the truthfulness of the information entered, which may lead to biased analysis.
-They also noted that AI’s interpretation of a child’s emotions may contradict the parents’ understanding.
-These concerns align with the communication aspect of cooperative AI (dafoe2020open).
-By explicitly sharing information, AI agents and parents can gain insights into each other’s behaviors and intentions more effectively.
-Here, we recommend leveraging research findings from explainable AI to enhance transparency in collaboration (liao2020questioning).
-Parents need to understand the rationale behind AI-generated assessments or suggestions, thus enhancing trust in AI’s decisions (kim2024establishing).
-In designing future systems, it is crucial to consider how to balance the parents’ cognitive burden of processing information with the goal of enhancing trust by providing a comprehensive explanation, which may be related to parents’ diverse decision-making styles (van2025selective).
-Considering the potential conflicts between parents and AI systems, we recommend the introduction of negotiation tools (crandall2018cooperating; jennings2001automated), as the current context lacks a formal agreement protocol.
-In the context of emotion education, it is essential to develop interactive negotiation mechanisms to address varying interpretations of children’s emotional states.
-Furthermore, AI can serve as a mediator to reconcile differing perspectives on children’s emotions among family members, a role that some parents expressed interest in during the interviews.
-Third, respect parents’ boundaries regarding AI intervention and provide them with sufficient autonomy.
-Prior work has explored the area of parents’ boundaries, such as parental concerns with AI-driven technologies in their child’s healthcare (sisk2020parental), showing that parents value shared decision-making but may resist interventions that undermine their sense of control.
-We supplement prior research and expand this area to include both emotion-sensitive contexts and early childhood education domains.
-In our context, children’s emotional education may involve sensitive family matters.
-In the interviews, parents expressed that they do not want AI to discuss family privacy or parents’ professional information.
-Future work could look into value-sensitive AI design frameworks (umbrello2021mapping) and provide information sensitivity detection techniques that align with diverse parental values.
-Moreover, emotional education may involve delicate feelings or negative topics that could affect children’s mood.
-To address this issue, PACEE incorporated the PERMA positive psychology model into its design to reduce the emphasis on negative content.
-Future work could investigate how to further reduce emotional overwhelm that may be perceived as harmful by parents.
-For example, adaptive AI systems could use real-time sentiment analysis to detect emotional intensity and compare it with predefined emotional boundaries set by parents.
-As for education for children aged 3 to 6, parents consistently viewed themselves as the primary decision-makers in their children’s education.
-In the interviews, parents expressed concerns about the possibility of AI imposing rigid goals on children and compelling parents to follow prescribed steps.
-Future systems should prioritize bidirectional discussions with parents, rather than simply providing prescriptions.
-Furthermore, they should ensure parents maintain authority over critical decision-making in parent-AI collaboration.
-8.4. Limitations and Future Work
-This study has several limitations.
-First, the participants’ cultural background was limited to Chinese parents and children, which may constrain the cultural generalizability of our findings.
-Second, most of the parent participants were mothers.
-This reflects the broader caregiving patterns noted in prior research (mcbride1993comparison; schoppe2013comparisons; hernawati2020differences), and likely represents the primary user group for a system designed to support caregivers of 3-6-year-olds in the real world.
-However, this imbalance may limit our ability to capture the full range of parental perspectives.
-Future work should recruit across more diverse cultural and gender contexts, and systematically examine how fathers and mothers may differ in their collaboration with AI to support children.
-Third, a limitation of the current study design is the inability to fully isolate the foundation model variable.
-While our study’s primary focus was on the interaction framework rather than raw model output, as supported by our findings and clarified in Section 6.2, we acknowledge that this does not fully eliminate the potential bias inherent in the two models.
-Therefore, future work could employ a 2x2 factorial design — (Interface: Baseline vs.
-PACEE) x (Model: Doubao vs.
-GPT-4o) — to further isolate the effect of the model.
-Besides, PACEE currently supports audio input and fixed UI interactions.
-Future developments could, with parental consent, enable the collection of children’s gestures and facial expressions to better assess their emotional and behavioral states during interactions.
-Finally, our study was conducted in lab-based settings using a touchscreen laptop, which may not fully reflect the long-term usage experience at home.
-In addition, first-time exposure to PACEE’s interactive features may have temporarily increased participants’ engagement, reflecting a “novelty effect” commonly faced by HCI research (stromel2024narrating; soler2024arcadia; davis2025sketchai).
-To address this issue, we provided parents with a practice session to familiarize them with the procedure.
-While our analyses did not reveal significant differences in comparative outcomes between early and later interactions, future work should explore how such novelty influences parents and children during longer interactions.
-Future work should also implement longitudinal in-home deployments to better understand how parent-AI collaborative emotion education unfolds in natural family environments.
-9. CONCLUSION
-We introduced PACEE, an assistant designed to support family emotion education through parent-AI collaborative emotional profiling and guidance.
-Drawing on interviews with kindergarten teachers and parents, PACEE leverages generative AI to address the key challenges identified in family emotion education.
-Our evaluation study with 16 pairs of parents and children demonstrated that PACEE enhanced parent-child engagement, fostering deeper emotional exchanges.
-Furthermore, PACEE significantly alleviated parental challenges, resulting in greater satisfaction and less effort from parents.
-We hope our work will advance emotion coaching theory in both family and AI-assisted contexts.
-Additionally, we hope to inspire the development of parent-centered, AI-supported education systems that empower families to promote children’s well-being.
-Appendix A Demographic Details of Participants in the Formative Study
-Table 1. Demographic details of kindergarten teachers in the interviews.
+Table 1. Demographic details of kindergarten teachers in the need-finding interviews.
 Alias Age Gender
@@
 K5 31 Female
-Table 2. Demographic details of parents in the interviews.
+Table 2. Demographic details of parents in the need-finding interviews.
 Alias Age Gender
@@
 O5 33 Female 5 Girl
-Appendix B Common Scenarios of Negative Emotions in Children Aged 3-6
-See Table 4.
-Table 4. Our formative study identified common scenarios of negative emotions experienced by children aged 3-6.
+(a) Observational experiment settings
+(b) A conceptual design image in the interview session
+Figure 3.
+Photos taken in the formative studies, including: (a) the setup of an observational experiment, where parents guide their children following a predefined script, and (b) a conceptual design image presented during the interview session to illustrate our design objectives.
+4.1. Procedure and Analysis
+4.1.1. Need-finding Interview with Experienced Kindergarten Teachers
+We interviewed 5 experienced kindergarten teachers (3 female, 2 male; $M=9.40$ years of experience, $SD=6.34$) recruited via social media, each receiving a \$10 stipend.
+Each interview session lasted about one hour and was conducted through video conferencing, moderated by two researchers.
+First, we interviewed the teachers on several key topics, including common negative emotions in children aged 3-6, triggers, and the strategies employed by teachers.
+Next, we explored the potential role of AI in supporting emotion education.
+Specifically, we introduced the design objectives and scenarios of parents providing regular emotional guidance to their children at home, while the AI offers additional support throughout the process.
+Furthermore, we co-designed an AI-assisted system based on participants’ visions and expertise, utilizing conceptual design images to illustrate our design objectives (see Figure 3(b)).
+4.1.2. Need-finding Interview with Parents of Children Aged 3-6
+We also interviewed 5 parents (4 mothers, 1 father; $M=41.0$ years old, $SD=3.16$) of children aged 3-6, recruited via social media, with \$10 stipends.
+Each hour-long video session was moderated by two researchers.
+First, we interviewed the parents about their observations of children’s emotional behaviors, challenges in guiding children’s emotions, and their understanding of emotional regulation.
+Next, our design scenarios explored the parents’ perspectives on AI-assisted tools, preferred features, and concerns.
+Together with the kindergarten teachers’ sessions, these interview sessions were audio-recorded, transcribed, and analyzed using a coding process by three HCI experts.
+4.1.3. Observational Experiment on Children’s Emotion Education
+To additionally understand the context of how parents and children communicate and interact during emotion education conversations, we observed emotion education interactions in 5 families recruited via social media, with \$10 compensation.
+The participants included four mothers, one father ($M=38.2$ years old, $SD=3.49$), and their children (4 boys, 1 girl; $M=4.8$ years old, $SD=1.17$).
+Each family received \$10 or an equivalent gift as compensation.
+Conducted in a lab over 30 minutes, parents guided their children based on a predefined script operationalizing Gottman’s Five Steps of Emotion Coaching guideline (Gottman and DeClaire, 1997).
+Parents followed a script and selected one of three negative emotion scenarios provided.
+Sessions were video-recorded, analyzed, and annotated by three HCI experts.
+Finally, we conducted a brief interview to ask parents about their experiences with coaching and any difficulties they encountered.
+4.2. Findings
+4.2.1. Key Challenges in Family Emotion Education
+Through interviews with 5 kindergarten teachers (K1-K5) and 5 parents (G1-G5), we identified common scenarios where children aged 3-6 struggle with negative emotions.
+These emotions can be difficult to manage and often require repeated guidance in early childhood.
+The findings serve as initial data for our system’s scenario database (see Table 4).
+Table 4. Our need-finding interviews identified common scenarios of negative emotions experienced by children aged 3-6.
 Type Example Scenarios Common Emotions
@@
 Anxiety, fear
-Appendix C Metrics in User Evaluation
-Here, we elaborate on the metrics applied in our user evaluation.
-Post-Questionnaire.
+Additionally, parental interviews revealed current practices and challenges in emotional communication and guidance.
+Combined with insights from observational experiments, we categorized the key challenges in family emotion education into three dimensions:
+$C_{1}$
+.
+Limited parental expertise.
+All five interviewed parents (5/5) acknowledged lacking professional-level skills in child emotion education, which manifests in three key aspects.
+First, parents demonstrated limited psychological and educational knowledge, struggling to apply general theoretical guidelines to specific situations.
+For example, G2 remarked: “I feel like I don’t have access to expert advice or reliable sources, and the examples in books don’t always apply to my kid’s situation.” We also observed similar patterns in the observational study, as parents also reported that they “don’t often apply every step before”(O3) and “don’t know what to say” about certain steps (O1).
+Second, parents reported insufficient communication techniques, emphasizing the need for “concrete, easy-to-implement language and techniques” (G1, G2, G5) in daily practice.
+Third, parents experience occasional impatience during emotional communication with their children.
+G3 observed that “Sometimes my husband gets impatient when he and our child disagree, and it makes our child hesitant to open up and share feelings with him”, illustrating how parental emotional instability can negatively impact communication effectiveness.
+$C_{2}$
+.
+Insufficient feedback.
+Parents frequently lack objective, professional feedback when guiding children’s emotional development.
+As G3 noted, “When he has strong emotions, I explain and comfort him in the moment, but I can’t tell if he really understood or whether my approach was effective”.
+In the observational experiment, parents reported similar uncertainties (O4: “I just followed the guidelines step-by-step, but I’m not sure each step worked.
+Next time this happens, he might forget what we discussed today.”).
+$C_{3}$
+.
+A lack of understanding of children’s emotional states.
+Most parents reported difficulties in accurately recognizing their children’s emotional states.
+This stems partially from limited awareness of school experiences, as G4 noted “not knowing what happens at kindergarten since children don’t always share things at school.” Parents may struggle to identify the root causes of emotions (G1: “Once he cried because he was just too tired, but I initially thought it was because of something that happened at kindergarten.”).
+Additionally, parents acknowledged their subjective biases when assessing their children (G5: “We naturally think our child is the best, so this bias colors our evaluations.
+We need something objective to show how our child compares with peers.”).
+4.2.2. Roles and Design Principles of AI in Family Emotion Education
+During the interview sessions, parents expressed their vision for AI’s potential roles in emotional guidance, while experts highlighted the cognitive and emotional characteristics of young children that should guide system design.
+All five participating parents expressed a preference for maintaining their primary role in emotional guidance while leveraging AI to supplement their limitations.
+They also envisioned specific roles for AI, including serving as a conversation facilitator, an expert resource, and a generative tool.
+4/5 parents anticipated AI’s involvement as a conversation facilitator.
+For instance, G3 suggested that AI could dynamically adjust its level of participation based on the parent’s engagement: “When parents take the lead, AI should speak less; when parents struggle, AI can step in to enrich the conversation and reduce communication barriers.” G2 further highlighted AI’s potential to assist less experienced parents: “If parents are unsure how to respond, AI could intervene.
+It could also detect and correct misguided parental guidance in real time.
+All five parents expected AI to serve as an expert resource.
+By integrating relevant knowledge into conversations, as G4 noted: “Providing real-time pedagogical insights or case studies, such as developmental benchmarks for 5-6-year-olds.” Additionally, it could offer post-conversation reports to improve parents’ long-term educational skills and provide objective assessments of children’s behavior to counteract parental biases, with G4 adding: “I hope AI can help me observe my child’s development more objectively and comprehensively.
+Parents also proposed practical applications for AI functioning as a generative tool.
+AI could reduce communication burdens by “generating images or videos to clarify scenarios for children” or “suggesting real-time conversation strategies”.
+It could also produce post-interaction reports with performance feedback, as G5 mentioned: “It could analyze which guidance techniques I applied successfully and how well I executed them.
+The interview sessions with kindergarten teachers highlighted the key cognitive characteristics of 3- to 6-year-old children, informing the design of AI-assisted emotion education systems.
+First, the system design must align with children’s developmental stage.
+Considering children’s tendency toward concrete and visual thinking, the system should incorporate visual aids, such as pictorial representations, to help them better grasp emotional concepts.
+Additionally, given their limited language skills, the interface should guide parents on how to use age-appropriate language during emotional coaching.
+Second, the system should provide personalized instruction to accommodate individual differences in emotional and cognitive development.
+Experts emphasized that developmental abilities vary significantly among young children.
+Thus, it is crucial to consider each child’s specific cognitive traits and emotional states to deliver tailored learning strategies.
+An AI system could effectively track behavioral patterns and provide developmental benchmarks, enabling personalized adjustments to the learning process.
+It is important to note that these design principles enable us to delve deeper into some of the main features of the core modules, which address three primary challenges in AI-assisted family emotion education.
+Since these principles are crucial design details that enhance the system’s functionality and user experience, they are not evaluated separately in subsequent studies.
+5. DESIGN AND IMPLEMENTATION OF PACEE
+5.1. User Interface and Design of System Modules
+Based on the insights from the formative studies, we designed and implemented PACEE, an LLM-based assistant for supporting parent-AI collaborative emotion education.
+PACEE consists of 3 modules—Parent-Child Adventure Game, Parental Feedback Report, and Child Emotional Modeling— each addressing a key challenge in family emotion education.
+Figure 4.
+Main screens of PACEE.
+PACEE consists of 3 modules: (a) Parent-Child Adventure Game.
+Each adventure game facilitates parent-guided emotional conversations centered on a specific emotional scenario, promoting meaningful interactions between parents and children.
+(b) Parental Feedback Report, which provides feedback on the adventure game and practical recommendations for parents.
+(c) Child Emotional Modeling collaborates with parents to develop a child’s emotional profile.
+Parents contribute long-term observational data about their child, gathered through interactions with an AI interviewer.
+5.1.1. Parent-Child Adventure Game
+To address the challenge of $C_{1}$ limited parental expertise, we designed the Parent-Child Adventure Game module.
+Parents and children can choose a scenario (e.g., performing on stage) from the system’s scenario database, which was initialized using formative study results.
+PACEE operationalized Gottman’s Five Steps of Emotion Coaching guideline (Gottman and DeClaire, 1997) into a five-stage conversational workflow (see Table 5).
+While maintaining the theoretical core, PACEE applied specific design adaptations to align with the developmental needs of children aged 3-6, as identified through our formative study interviews.
+The five-stage conversational workflow ( $S_{1}$ - $S_{5}$) are later referred to as the FiveStagesFive\ Stages in this paper.
+Table 5. How we operationalized Gottman’s Five Steps of Emotion Coaching guideline into PACEE’s conversation stages.
+Tag Gottman’s Guideline (Gottman and DeClaire, 1997)
+Developmental Consideration
+of Children Aged 3-6
+PACEE’s Conversation Stage
+$S_{1}$
+Be aware of the child’s
+emotion. Recognize the
+expression of emotions
+Have limited memory recall.
+Need for concrete context.
+Help the child recall an
+experience or describe a
+virtual scenario
+$S_{2}$
+Help the child learn to
+label their emotions
+Need scaffolding to articulate
+feelings
+Help the child label
+emotions and reason
+$S_{3}$ Listen with empathy
+Respond well to positive
+reinforcement (e.g., PERMA
+model (Seligman, 2010))
+Express empathy to children
+$S_{4}$
+Validate the child’s
+feelings
+Help the child reflect on
+positive and negative emotions
+$S_{5}$
+Set limits when helping
+the child to solve problem
+Need help or collaboration
+in problem solving
+Help the child set boundaries
+and find positive solutions
+During the five-stage conversation, PACEE offers a range of support for parents.
+As a conversation facilitator, PACEE can join the conversation when prompted by the parent, which helps reduce the communication burden.
+As an expert, PACEE provides professional advice for each of the 5 steps, considering the child’s emotional profile.
+It also periodically updates real-time advice, offering communication strategies tailored to the cognitive characteristics of children aged 3 to 6 years old.
+As a tool, PACEE can generate images of the current scenario for parents to show to their children, catering to children’s preference for visual thinking.
+Additionally, PACEE monitors the conversation’s progress and stage, providing useful references for parents to decide when to progress to the next stage.
+After completing a scenario, PACEE generates rewards, such as stars and medals, for the children, supporting their preference for encouragement-based learning.
+5.1.2. Parental Feedback Report
+To address the challenge of $C_{2}$ insufficient feedback, we designed the Parental Feedback Report module.
+PACEE evaluates parents’ performance during the Parent-Child Adventure Game, and offers actionable guidance for future interactions.
+Over extended use, parents have the opportunity to unlock achievement badges, enhancing their sense of accomplishment.
+For evaluating parental performance, PACEE analyzes real-time transcripts of parent-child conversations.
+This analysis provides insights into how parents effectively navigate the emotional education process and establish secure attachments with their children.
+PACEE also delivers future suggestions grounded in established theories (e.g., Gottman’s guideline (Gottman and DeClaire, 1997), PERMA model (Seligman, 2010)).
+By tailoring these insights to the early learning and development guidelines for children aged 3 to 6 years (of Education of the People’s Republic of China, 2012), the system ensures the guidance is both practical and developmentally appropriate.
+5.1.3. Child Emotional Modeling
+To address the challenge of $C_{3}$ a lack of understanding of children’s emotional states, we designed the Child Emotional Modeling module.
+Grounded in the Test of Emotion Comprehension (Pons et al., 2004; Albanese et al., 2006), we model children’s emotions across the following three dimensions:
+  • Emotional understanding, including emotion recognition, emotion elicitors, emotion-related beliefs, the impact of memory on emotions, and mixed emotions.
+  • Emotional expression, including emotional expression and emotional masking.
+  • Emotional regulation, including emotion regulation abilities and moral emotions.
+Figure 5.
+Mechanism of PACEE modeling children’s emotional profiles collaboratively with parents.
+It functions in two scenarios: (1) individual parent interviews conducted by an AI Interviewer and (2) AI-facilitated parent-child emotion conversations monitored by a Conversation Agent.
+The Cognitive Analyzer processes inputs from both scenarios, analyzing and synthesizing emotional profile data.
+The resulting Child’s Emotional Profile includes three sub-dimensions: emotion understanding, emotion expression, and emotion regulation.
+Additionally, it compares the parent’s understanding of the child’s emotions with conversational analysis by the AI, offering a nuanced perspective on the child’s emotional development.
+Figure 6. Agent workflow of PACEE.
+The primary data source for modeling is the parent’s long-term understanding of the child in the physical world (Scenario 1 in Figure 5).
+This information is elicited through an AI interviewer, which follows a semi-structured interview workflow.
+The process begins by selecting a question from a fixed list, and based on the parent’s verbal responses, the interviewer decides whether to ask follow-up questions, adjusting the phrasing as necessary.
+If no further probing is required, the interviewer proceeds to the next question on the list.
+This question list is also derived from the Test of Emotion Comprehension (Pons et al., 2004; Albanese et al., 2006).
+Concurrently, PACEE parses the parent’s responses to construct the child’s initial emotional profile.
+The complementary data source is PACEE’s text-based analysis of parent-child conversation (Scenario 2 in Figure 5).
+During the Parent-Child Adventure Game, PACEE segments, transcribes, and analyzes conversations to identify observations that enrich the child’s emotional profile.
+This process helps address limitations such as human attentional constraints and subjective biases.
+Both data sources are analyzed by the cognitive analyzer.
+First, the profile extractor identifies incremental updates across the three dimensions of emotional understanding, emotional expression, and emotional regulation.
+Next, the profile integrator merges these updates into the existing profile while minimizing redundancy.
+The profile integrator also compares and organizes entries contributed by the parent and the system within each dimension, revealing potential cognitive gaps.
+Finally, PACEE presents a comparative analysis of children’s emotional states on the Child Emotional Modeling interface (see Figure 4 (c)).
+5.2. LLM-based Agent Workflow Design
+In this section, we introduce the agent workflow design of PACEE (see figure 6).
+In the Parent-Child Adventure Game, PACEE applies rapid and reflective reasoning to real-time conversation information.
+Inspired by how human systems of “thinking fast and slow” (Kahneman, 2011) can be applied to generative agents (Christakopoulou et al., 2024; Lin et al., 2023), we implement rapid and intuitive reasoning to generate real-time coaching support for parents, while employing reflective and deliberative reasoning steps to understand the emotional characteristics of the child progressively.
+PACEE transcribes the real-time parent-child conversations and monitors explicit transitions across the FiveStagesFive\ Stages by mouse clicks.
+The conversation analyzer uses this information to assess the completion level of the current conversation stage and update the node status in the conversation graph.
+The content generator module, combining conversation information and the child’s emotional profile, applies rapid reasoning to provide timely support for parents: (1) phase advice updated at each stage, (2) real-time advice updated every half minute, (3) agent conversation converted into voice mode, and (4) scenarios converted into image mode by text-to-image models.
+In addition, the content generator module integrates all conversation information from the latest Parent-Child Adventure Game and the child’s emotional profile to generate the Parental Feedback Report for the parents.
+PACEE also applies reflective reasoning within the parent-child conversation to gradually build the child’s emotional profile based on real-time conversation information.
+At certain intervals (empirically set to every five conversation turns), the cognitive analyzer is tasked with multi-step reasoning to update the child’s emotional profile (see Section 5.1.3).
+5.3. Implementation
+We implemented the core agent workflow in Python on a Flask server.
+The PACEE interface, written in Vue, interacts with the server by exchanging messages via WebSocket.
+Children’s emotional profile and conversation graph information are stored in separate JSON files.
+For most text generation tasks, we leveraged the GPT-4o API to balance speed and output quality.
+To minimize latency in agent conversations and real-time advice delivery, the faster GPT-4o-mini model was adopted.
+Speech-to-text processing utilized the TINGWU API, with audio sliced and compressed on the frontend before transcription.
+Text-to-speech conversion relied on Alibaba Cloud’s streaming synthesis API, while image generation integrated the Stable Diffusion XL (SDXL) model.
+6. USER EVALUATION
+We conducted a within-subject study to evaluate how PACEE supports family emotion education collaboratively with parents. We aimed to answer the following questions:
+RQ1. How does PACEE enhance family emotion education?
+RQ2. How do parents collaborate with PACEE to develop children’s emotional profiles and offer guidance to children?
+RQ3. How do parents perceive PACEE in addressing their challenges in family emotion education?
+We included a baseline condition to validate the effectiveness of PACEE’s scaffolding features.
+Each pair of parent and child participants was tasked with two common emotional scenarios from PACEE’s scenario database, one under each condition.
+Table 6. Participant demographic information in the user evaluation.
+Child ID Child Gender Child Age Parent ID Parent Gender Parent Age
+C1 Female 5 P1 Male 41
+C2 Male 5 P2 Male 36
+C3 Female 4 P3 Female 37
+C4 Male 5 P4 Female 43
+C5 Female 5 P5 Female 38
+C6 Female 4 P6 Female 43
+C7 Female 5 P7 Female 34
+C8 Male 6 P8 Female 36
+C9 Female 5 P9 Female 39
+C10 Female 4 P10 Female 29
+C11 Male 5 P11 Female 39
+C12 Male 5 P12 Female 39
+C13 Male 6 P13 Male 39
+C14 Male 5 P14 Female 36
+C15 Male 6 P15 Male 36
+C16 Male 6 P16 Female 38
+6.1. Participants
+We recruited 16 pairs of parents and children from social media platforms.
+Parent participants (P1-16) included four fathers and twelve mothers, with an average age of 37.69 years ($SD=3.31$).
+Child participants (C1-16) included seven girls and nine boys, with ages ranging from 3 to 6 ($M=5.06,SD=0.66$).
+Table 6 summarizes the demographic information of participants.
+We carefully took ethical measures for participants in our user study.
+First, we obtained verbal consent from the parents and child participants, confirming their willingness to participate in the study.
+Second, we informed the participants that they had the right to withdraw from the experiment at any stage without facing any consequences.
+Third, parents signed a written consent form before the experiment, which detailed the complete process, the types of information we collected (video data, audio data, screen activity), and the anonymization methods.
+All 16 parent participants agreed to sign the consent form.
+Lastly, we conducted remote real-time manual reviews on the server during the experiment.
+These reviews included examining text and images to ensure their content appropriateness and to prevent displaying any harmful content to the children.
+6.2. Baseline Condition
+Figure 7.
+Illustration of Baseline condition.
+Parents engaged with a generic LLM chatbot (Doubao) via a voice interface.
+To maintain conversational structural consistency with PACEE’s experimental condition, parents followed FiveStagesFive\ Stages conversation guide displayed on-screen.
+In contrast, the Baseline lacked PACEE’s key parental scaffolding features: Parent-Child Adventure Game, Parental Feedback Report, and Child Emotional Modeling.
+To validate the effectiveness of PACEE’s scaffolding features, we designed a Baseline condition in which parents engaged with a standard LLM chatbot via a web-based voice interface (see Figure 7, left).
+Parents can request assistance from the chatbot when needed, with no restrictions on usage frequency.
+We selected the Doubao chatbot  [^1] https://www.doubao.com/chat/ to ensure a realistic and robust comparison.
+Our selection was guided by three reasons: (1) As one of the most popular free AI chatbots in Chinese families (Jiang, 2024; Ahmed, 2025; Yu et al., 2025), it provides an ecologically valid baseline that participants might encounter in daily lives.
+(2) It has strong support for real-time Chinese audio conversations, which aligns with PACEE’s hands-free interaction flow.
+(3) To ensure our comparison focused on the interaction framework rather than model capability, we selected a model with competitive performance in the context of our study.
+In the Chinese context, Doubao-Seed-1.6-thinking-250715 model outperforms the ChatGPT-4o-latest model in the SuperCLUE Benchmark  [^2] https://www.superclueai.com (Xu et al., 2023a).
+To maintain conversational structural consistency with PACEE’s experimental condition, parents were provided with an on-screen guide outlining the FiveStagesFive\ Stages conversational workflow (see Figure 7, right).
+While the workflow stages were prescribed to ensure comparability, the specific conversational content within each stage remained open-ended, allowing parents and children to engage in a flexible conversation.
+6.3. Evaluation Design
+We randomly chose two scenarios from our system’s scenario database (see Table 4), which was initialized during the formative study: (1) not winning first place (referred to as Not First), where children often encounter anger and disappointment, and (2) performing on stage (referred to as Up Stage), where children often encounter anxiety and fear.
+As illustrated in Figure 8, we employed a counterbalanced design, where each parent-child pair was randomly assigned to one of four groups.
+These groups were determined by the order in which the two conditions (Baseline vs.
+PACEE) and the two emotional scenarios (Not First vs.
+Up Stage) were presented.
+Figure 8. Design of the evaluation study. We followed a counterbalanced design where each pair of parent and child was randomly assigned to one of four groups.
+(a) Evaluation experiment settings using PACEE.
+(b) Children create drawings after the experiment.
+Figure 9.
+Photos taken in the evaluation studies, including: (a) the setup of an evaluation experiment session, where parents and child engage in emotion education using PACEE, and (b) a drawing created by a child after the experiment session, depicting the Not First scenario discussed with PACEE.
+6.4. Procedure
+Each pair of parent and child underwent a 2-hour study in our lab (see Figure 8). We provided a touch-screen computer for each family and recorded the entire evaluation process.
+Introduction session.
+Before each emotion education session, we introduced the parent participants to the designated system.
+We demonstrated how to interact with each system and provided tips to avoid misuse.
+We also provided parents with a practice session using several example scenarios to familiarize them with the procedure.
+In addition, we informed parents that the session required joint participation from both parents and children, and it was not intended for the child to engage solely with the AI.
+Emotion Education Session Using PACEE or Baseline.
+In each emotion education session, we asked parents to freely use the system to learn about the specified emotional scenario with their children.
+Considering children aged 3 to 6 may be shy around strangers, the researcher left the room to allow parents and children to interact freely.
+We monitored the session via a live video feed and informed parents that they could contact us at any time if they encountered any issues.
+Post-Questionnaire Filling.
+Parent participants completed a post-questionnaire to evaluate their educational experience after completing all education sessions.
 We used a questionnaire consisting of three sections.
-The first section consisted of NASA-TLX (hart1988development) metrics to measure the mental demand, effort, and frustration of parents when using AI systems to provide emotional guidance to children.
-The second section used the Parenting Sense of Competence Scale (PSOC) (gibaud1978development) to measure parents’ confidence, satisfaction, and self-efficacy.
-The last part focuses on the extent to which PACEE or Baseline address the three key challenges ( $C_{1}$- $C_{3}$) in family emotion education.
+The first section consisted of NASA-TLX (Hart and Staveland, 1988) metrics to measure the mental demand, effort, and frustration experienced by parents when using AI systems to provide emotional guidance to their children.
+The second section used the Parenting Sense of Competence Scale (PSOC) (Gibaud-Wallston and Wandersmann, 1978) to measure parents’ confidence, satisfaction, and self-efficacy.
+The last part focuses on the extent to which PACEE or Baseline addresses the three key challenges ( $C_{1}$- $C_{3}$) in family emotion education.
 All ratings were done on a 7-point Likert scale.
-Semi-structured interview.
-In the semi-structured interview, we asked parents about the impact of the AI system on emotion teaching between parents and children, its role in overcoming challenges, and its effect on the FiveStepsFive\ Steps.
-We discussed their experiences with each module of PACEE and broader topics like AI’s strengths in collaborative teaching and the essential roles of parents.
-The interviews were audio recorded, transcribed, and then analyzed thematically (clarke2017thematic) by four experienced HCI researchers.
-To ensure consistency with annotation standards, each part of the analysis was reviewed by at least two researchers.
-Appendix D Quantitative Data Analysis
-We transcribed the audio recordings from the teaching sessions of both the PACEE and Baseline systems into text and then conducted automated quantitative analyses.
-The identification of emotion-related words was implemented using the NTUSD dictionary (ku2007mining).
+Parent Interview and Child Drawing.
+One researcher interviewed the parents about their user experience and perceptions.
+At the same time, another researcher asked the child participants to use emoji stickers to represent their feelings during conversations, the scenarios discussed, and the image of PACEE.
+Finally, the children drew a picture to depict the scenarios discussed with PACEE.
+6.5. Data Analysis
+For qualitative analysis, we transcribed all parent interviews.
+Five HCI researchers with qualitative coding experience conducted a thematic analysis (Braun and Clarke, 2006; Terry et al., 2017; Braun and Clarke, 2019).
+The process began with a data familiarization phase, during which all researchers reviewed the transcripts to gain a comprehensive understanding of the content.
+The leading author then generated initial code themes, which were reviewed and refined with feedback from another researcher.
+The team then collaboratively developed a shared codebook by coding data from one participant (P1), with discussions for consensus.
+The remaining data were then independently coded, and disagreements were resolved through discussion.
+Finally, the team manually and iteratively clustered the codes to identify recurring and meaningful patterns.
+For statistical analysis, we transcribed the audio recordings from the teaching sessions of both the PACEE and Baseline systems into text and then conducted automated quantitative analyses.
+The identification of emotion-related words was implemented using the NTUSD dictionary (Ku and Chen, 2007).
 To better analyze the transcript data, we applied descriptive statistical methods, calculating the mean and standard deviation for each of the four dimensions among all 16 groups of participants.
 Additionally, we employed significance testing methods to evaluate the statistical significance of the collected data.
-We used the Paired Sample t-test for interaction duration data and the Wilcoxon signed-rank test for the number of dialogue turns, word counts, and emotion-related word counts, based on the distribution characteristics of the data.
-Furthermore, we calculated the mean and standard deviation of dialogue durations across the FiveStepsFive\ Steps.
-Appendix E More Findings in User Evaluation
-See Table 5 and Table 6.
+We used the Paired Sample t-test for interaction duration data and the Wilcoxon signed-rank test for the number of conversation turns, word counts, and emotion-related word counts, based on the distribution characteristics of the data.
+Furthermore, we calculated the mean and standard deviation of conversation durations across the FiveStagesFive\ Stages.
+7. FINDINGS
+7.1. RQ1. How Does PACEE Enhance Family Emotion Education?
+7.1.1. Increased Parent-Child Interaction
+Figure 10.
+User interaction data from the PACEE and Baseline systems during parent-child conversation sessions.
+The x-axis indicates participant identity for (a-d) and the conversation stage for (e), with error bars representing standard deviation.
+Significance analysis: (a, e) uses the Paired Sample t-test, while (b-d) are based on the Wilcoxon signed-rank test.
+Significance results are marked with asterisks (***: p ¡ .001, **: p ¡ .01, *: p ¡ .05).
+To compare user engagement and interaction patterns between the PACEE and Baseline systems, we conducted descriptive statistics and significance testing across four key dimensions: interaction duration, number of dialogue turns, word count, and the count of emotion-related words.
+The results show that PACEE consistently outperforms the Baseline, with statistically significant advantages across all four areas.
+For PACEE, the average interaction duration was 892.37 seconds ($SD=310.96$), with an average of 60 dialogue turns ($SD=46.979$).
+In contrast, the average interaction duration for the Baseline system was 560.17 seconds ($SD=194.61$), with an average of 38 dialogue turns ($SD=25.3$).
+Considering that both systems were matched in workload and participants were given flexibility in session completion, the results suggest that PACEE leads to higher user engagement, with significant differences in interaction duration ($p<0.01$) and dialogue turns ($p<0.05$).
+This difference was evident in both parents and children.
+Parents using PACEE produced an average of 1,597 words ($SD=678.56$), compared to 994 words ($SD=468.36$) with the Baseline ($p<0.05$).
+Children expressed an average of 370 words ($SD=240.43$) with PACEE and 165 words ($SD=102.25$) with the Baseline ($p<0.001$), as shown in Figure 10.
+To further examine conversation quality alongside user engagement, we compared the number of emotion-related words used by parents and children.
+Under PACEE, parents used an average of 111 emotion-related words (53 positive, 58 negative), while children used 19 (8 positive, 11 negative).
+In contrast, under the Baseline, parents used an average of 70 emotion-related words (31 positive, 39 negative), and children used 10 (5 positive, 5 negative).
+Both parents ($p<0.05$) and children ($p<0.01$) are more willing to express their emotions using PACEE.
+These findings suggest that PACEE fosters more meaningful emotional exchanges, which are essential for supporting children’s early development of language and emotional competence (Brushe et al., 2025; Reschke et al., 2024).
+To gain a more intuitive understanding of participants’ engagement across the FiveStagesFive\ Stages of the teaching session, we calculated the average time spent at each step for both systems, as shown in Figure 10 (e).
+Participants’ engagement was generally higher in PACEE at all five steps, particularly in $S_{1}$ ($p<0.05$) and $S_{4}$ ($p<0.05$).
+Considering the limited memory capacity of preschool children (Scarf et al., 2013; Roman et al., 2014), the higher participation time in $S_{1}$ suggests that PACEE effectively supported participants in recalling scene details (PACEE $M=183.38$ vsvs Baseline $M=123.50$; $p<0.05$).
+Meanwhile, in $S_{4}$, PACEE participants engaged in significantly more reflection on positive and negative emotions (PACEE $M=163.86$ vsvs Baseline $M=85.33$; $p<0.05$).
+7.1.2. Sustained Participation Over Time
+To investigate the potential influence of the novelty effect, we conducted a temporal analysis by dividing each interaction session within the PACEE system into two equal halves by time (see figure 11).
+Our analysis revealed no statistically significant differences between the first and second halves of the session regarding user dialogue turns, the number of expression words, or the number of emotion-related words over time.
+This stability suggests that high user participation is a consistent feature within PACEE.
+Figure 11.
+Temporal analysis of user interaction data within the PACEE system, comparing interaction during the first and second temporal halves of parent-child conversation sessions across three metrics: (a) number of dialogue turns, (b) user word count, and (c) the count of emotion-related words.
+Error bars represent standard deviation.
+No significant differences were found between the two periods based on Wilcoxon signed-rank tests.
+7.1.3. Improved Parental Experience and Children’s Engagement Respectively
+Figure 12.
+Post-questionnaire data on parental experience, including metrics from the NASA-TLX and the Parenting Sense of Competence Scale.
+We applied the Wilcoxon signed-rank tests to analyze the significance of each metric.
+Statistically significant ratings are marked with an asterisk (*).
+Figure 12 shows parent participants’ responses to the post-questionnaire on their parental experience.
+Compared to the Baseline, PACEE improves parents’ immediate interaction experience, with higher satisfaction ($p<0.05$), reduced effort ($p<0.01$), and less frustration ($p<0.05$).
+However, intrinsic psychological factors (e.g., parenting confidence or effort with the child) show no significant change during the study, likely because these factors need long-term exposure and reinforcement to yield measurable differences (Bandura, 1997; Sandler et al., 2011).
+Additionally, we analyzed the children’s post-study drawings to explore their perceptions of PACEE.
+Thirteen of the 16 children were able to accurately reconstruct key interaction scenes, such as the stage, parents, and audience, demonstrating strong memory retention.
+Their illustrations often conveyed positive emotions through visual cues, including smiley faces, hand-holding gestures, and a dominance of warm color palettes.
+For example, C1’s drawing depicted the stage with smiling characters rendered in warm hues.
+Similarly, C5’s drawing was colorful, featuring balloons, smiles, and hand-holding gestures.
+Beyond scene recollection, the drawings revealed nuanced perceptions of PACEE itself.
+Thirteen of the 16 children portrayed PACEE as a friendly and supportive figure.
+Common representations included anthropomorphic characters with open arms and smiling faces (C1, C8, C12, C16), friendly cartoon animals like cats with smiles (C5, C6, C10), or simplified smiley-face icons (C3, C7, C13).
+Personalized elements, such as C9’s drawing of PACEE resting on the child’s own head, further implied emotional attachment.
+In contrast, three of the 16 children (C2, C4, C11) depicted PACEE as a machine-like figure with neutral or absent facial expressions.
+These children also had difficulty conveying scene details and emotions in their drawings.
+These findings suggest that perceiving PACEE as impersonal may correlate with lower engagement.
+Figure 13. Drawings created by children (C1, C5, C10) after the experiment.
+7.2. RQ2. How Do Parents Collaborate with PACEE to Develop Children’s Emotional Profiles and Offer Guidance to Children?
+7.2.1. Parents and PACEE Contribute Different Factors in Construction of Children’s Emotional Profiles
+According to experts’ annotations, parents and PACEE play distinct roles in constructing children’s emotional profiles.
+Regarding data sources, PACEE derives its data from the content of parent-child communication.
+At the same time, parents tend to provide information based on firsthand observations of events, third-party feedback, school progress, and growth trends compared to earlier childhood.
 Table 7. Parents and PACEE contribute different themes in children’s emotional profiles in the evaluation studies.
@@
 describe his feelings.
+Furthermore, Table 7 illustrates the cognition-level themes in children’s emotional profiles identified in our evaluation studies.
+We found that parents prefer to extract insights from various life events, focusing on experiential knowledge of their children’s habits.
+In contrast, PACEE evaluates developmental capabilities by analyzing parent-child interactions, emphasizing expert assessments of abilities and developmental characteristics.
+7.2.2. PACEE Provides New Insights on Children’s Profile for Parental Reflection
+During the interviews, parents acknowledged that PACEE offers a more objective assessment of children’s emotional capacities.
+Some parents realized they had underestimated their child’s emotional understanding after reviewing the AI’s evaluation.
+As P2 noted, “Sometimes I didn’t realize how well my child could understand emotions, but the AI, speaking with the tone of an expert, showed me that he is truly growing.” On the other hand, some parents discovered they had overestimated their child’s abilities.
+As P4 shared, “Parents often view their children through a biased lens, believing they’re performing well.
+However, the AI objectively highlighted certain issues, and despite its evaluation diverging from my own perspective, I found the AI more rational and unbiased.” Additionally, AI can help parents understand the underlying reasons behind their child’s emotional responses.
+As P8 explained, “When my child participated in an English competition, he was very reluctant to go on stage.
+I assumed he was nervous and anxious because he was afraid of the stage, but the AI suggested that the anxiety might have stemmed from the absence of his mother’s support.
+In a broader view, parents express opinions on AI’s advantages in modeling children’s emotional profiles, including more comprehensive summaries, more objective perspectives, and more educational insights.
+Additionally, AI has potential disadvantages, including a lack of long-term data, insufficient specificity, and incorrect inferences.
+These factors can lead to cognitive differences between parents and AI regarding children’s emotional states.
+We further inquired how parents perceive these differences.
+14/16 parents considered these differences between humans and AI as valuable new insights.
+They acknowledged that the individual suggestions made by AI are derived from broader patterns observed in children of similar age, allowing for more long-term and holistic reflections on a child’s emotional tendencies.
+As P16 explained, “I can only see the emotional issues my child is currently experiencing and how he copes with them, I have no experience dealing with newly emerging emotions.” In contrast, AI offers a more rational point of view that can help parents break free from fixed thought patterns.
+P8 remarked, “Because our communication with the child is usually smooth, we’ve fallen into a comfortable but fixed mode of interaction.
+While this is pleasant for daily life, it might hinder development in certain areas.
+The involvement of AI could reduce this limitation.
+However, 2/16 parents expressed concerns about the accuracy of AI-generated insights, viewing discrepancies from their own understanding as potential system errors.
+P15 commented, “The AI detected a sense of compromise in my child during the conversation, but I doubt he has the capacity to express such emotion yet.
+The term ’compromise’ seems imprecise.” Some parents also believed that AI could not detect when a child was lying, whereas their familiarity with their child enabled them to make more accurate judgments.
+As P11 noted, “AI can’t tell whether what the child says is sincere or not —it records everything regardless of accuracy.
+So if the child lies, it might offer the parent an incorrect solution.
+7.2.3. Parents Appreciate Open and Positive Advice from PACEE
+Parent participants received a total of 513 updates of real-time advice during the studies.
+Two experienced HCI researchers reviewed system logs and video recordings to determine whether the parents adopted the suggestions.
+They then iteratively developed classification criteria for the accepted advice, selecting the most prominent type in cases where advice fell into multiple categories.
+Table 8 summarizes the types of accepted real-time advice.
+We found that the types of advice most frequently adopted by parents were collaborative problem-solving and open-ended questioning.
+Parents also mentioned the usefulness of open-ended questioning in the interview, as P2 stated, “Real-time advice helps me realize some issues I hadn’t thought about, such as prompting my child to provide more details.
+Sometimes I think it’s okay when a child says one word, but the advice reminds me to ask for more details.” Furthermore, advice on scenario simulation, concrete suggestions, positive encouragement, and empathy and acceptance also helped some parents.
 Table 8. Types of accepted real-time advice during the studies.
@@
 competition.
+7.2.4. Parents Accept PACEE in Collaborative Guidance, but with Concerns
+Parents shared their perspectives on integrating PACEE into the emotional guidance process during the interviews.
+Participants consistently viewed themselves as the primary agents in their children’s education.
+Nine obviously reported making the most significant contribution during the experimental sessions, whereas only P10 regarded the PACEE as having had a larger impact.
+Furthermore, P5, P8, and P13 explained that their primary role was to evaluate and select which of the PACEE recommendations to implement, while P6, P9, P10, and P13 credited the system with providing the specialized professional expertise necessary to complete the tasks.
+In addition, parents expressed a positive attitude toward incorporating PACEE into parent–child conversations.
+All participants accepted PACEE involvement, though their preferred levels and modes of engagement varied.
+Some parents demonstrated a high degree of acceptance (P5: “My kid really like it(to receive emotional coaching from PACEE).
+I think I am very receptive.”), whereas others felt that PACEE was not strictly necessary, especially in familiar emotional scenarios (P16: “We often discuss this topic so my child no longer need emotional guidance in this situation.”).
+Parents recognized the system’s capacity to deliver valuable, professional, and targeted recommendations.
+For example, P8 highlighted that such suggestions made parent–child communication more effective, “He felt I understood him better, so he answered every question loudly and confidently.
+When discussing future interactions with PACEE, participants articulated clear expectations about PACEE’s role in parent–child conversations, though they differed markedly on its ideal interlocutor.
+Some parents (P2, P16) envisioned AI forging a peer-like relationship with their child, simultaneously acting as an educator.
+Others (P8) preferred that PACEE interact only with the parent, acting as a co-educator, and avoid direct communication with the child.
+A third group of participants (P10, P13, and P15) believed that AI should adopt a neutral stance as an observer, providing limited, objective, data-driven suggestions to parents and offering age-appropriate guidance to children, rather than engaging directly with either parent or child.
+For instance, P10 argued that “It can serve as a mediator (in parent–child emotional education).” P13 and P15 prefer PACEE to offer suggestions from a third-party perspective for both parents and children.
+Crucially, parents asserted firm boundaries around AI intervention.
+Some parents rejected features that impose targets on children (P4: “For example, setting goals for children… I think it is up to parents to do it.”), and they were adamant that PACEE should not encroach on sensitive family matters (P1: “Don’t press further… for example, on topics that parents find taboo and children find sensitive.
+”), or strictly require compliance with its advice (P11: “If it(PACEE) tells me I have to do something, I definitely won’t accept it.”).
+Moreover, participants cautioned that AI should not replace emotional intimacy and empathetic support intrinsic to parent-child bonding and the cultivation of self-confidence, such as “(I hope she can) build her confidence by communicating with others.”(P2), “I prefer she get emotional companionship from people…”(P5).
+7.3. RQ3. How Do Parents Perceive PACEE in Addressing Their Challenges in Family Emotion Education?
+The subjective ratings provided by parents indicate that PACEE significantly outperformed the Baseline in tackling the three main challenges of family emotion education (see Figure 14), with statistical significance observed in each case ($p<0.01$).
+Figure 14.
+Post-questionnaire data focusing on the extent to which PACEE or Baseline addresses the three key challenges.
+We applied the Wilcoxon signed-rank tests to analyze the significance of each metric.
+Statistically significant ratings are marked with an asterisk (*).
+Interviews with parents further confirmed that PACEE plays a multifaceted role in addressing the three main challenges of family emotion education.
+7.3.1. Enhancing Parental Expertise and Communication Confidence
+In terms of $C_{1}$ limited parental expertise, the Parent-Child Adventure Game module significantly alleviated parents’ difficulties in situations where they “don’t know what to say” or “don’t know how to guide.” During the interviews, parents acknowledged that they find it difficult to perform some steps alone, including recalling an experience ($n=4$), labeling emotions ($n=8$), expressing empathy ($n=8$), reflecting on emotions ($n=5$), and finding positive solutions ($n=6$).
+The system’s conversation suggestions were praised for mitigating these expertise gaps.
+First, it provides a structured framework for emotional communication.
+Two parents specifically highlighted how the AI facilitated guiding emotional conversations with children in a gradual, structured manner.
+As P8 noted: “The systematic process — understand the child first, then ask about their feelings, then summarize, encourage, and finally solve the problem.
+This step-by-step guidance is great, creating remarkably smoother interactions.” Second, several parents appreciated how AI helped them adopt their child’s perspective.
+P10 explained: “Placing us parents in a conversational setting that feels more like talking to a friend actually encourages us to respond more actively and calmly.
+It’s because of this setup that I’m more willing to take the child’s perspective, understand them, and help them manage their emotions.” Lastly, the conversation suggestions offer ideas to steer the conversation stream effectively.
+P16 stated, “AI’s reminders can offer us more ideas.
+During the reflection stage, both the child and I are thinking, and it’s easy to get caught up in the moment.
+The AI’s prompts provide a clearer perspective, which helps make the conversation more in-depth.
+Beyond providing suggestions, the module actively enhanced parent-child communication and boosted parental confidence during real interactions.
+Most parents reported increased initiative and comfort in emotional conversations after using the Parent-Child Adventure Game module.
+Ten participants confirmed that the system’s conversation suggestions provided concrete, child-friendly advice for expressing emotions.
+As P5 explained, “It can serve as a reference tool.
+For example, if you have an idea but don’t know how to express it, the system helps you organize your thoughts and express them more clearly and reasonably.” Additionally, two parents mentioned that AI-provided prompts served as valid entry points, promoting more open communication between children and parents.
+P1 explained: “As a father, I don’t talk with my child that often.
+For example, instead of saying, ‘Do you remember your last kindergarten performance?’ it feels more natural to ask, ‘How was your last performance?’ So the detailed suggestions provided by the system are quite helpful for me as a reference.” Some parents also noted that the module facilitated more reciprocal exchanges, helping to enhance parent-child relationships.
+While the benefits of AI assistance were frequently highlighted, some parents expressed concerns.
+A few parents felt that frequently checking the AI’s suggestions could slightly disrupt the conversational flow.
+One parent worried that the AI’s scripted suggestions might make children feel nervous or overly reliant.
+Another parent emphasized that AI should act as an aid rather than lead in communication, suggesting it should provide direction rather than take over the conversation entirely.
+Furthermore, some saw the structured nature of AI prompts as potentially limiting the child’s freedom to express diverse experiences.
+There was also concern that excessive reliance on AI-generated expressions might hinder parents’ own communicative abilities.
+7.3.2. Providing Trackable Feedback and Actionable Future Strategies
+In terms of $C_{2}$ insufficient feedback, the Parental Feedback Report module effectively addressed pain points such as “lack of timely feedback” and “invisible outcomes”.
+Eight parents highlighted how AI-generated analyses shed light on previously unnoticed shortcomings in their parenting approaches, which prompted them to engage in significant self-reflection.
+For instance, P6 shared, “The system reminded me to avoid interrupting.
+I hadn’t realized I was doing that, but it made me see that I still need to be more patient in communication.” Furthermore, the module provided a structured review by organizing conversation highlights, which five parents found particularly beneficial in bridging feedback gaps and making abstract parenting efforts trackable.
+This transformation into visible progress was crucial for parents to assess their parenting strategies.
+Beyond reviews on past interactions, the module also provides actionable future strategies.
+Eight parents appreciated the evidence-based nature of the strategies provided, while three were pleased with the introduction of new, beneficial perspectives.
+P12 noted, “For example, when a child loses control, besides suggesting enjoyable activities, the system also mentions deep breathing and meditation.
+I used to think those were only for adults, but this made me realize they might work for kids too.”.
+Also, P1 noted: “AI can indeed offer suggestions we wouldn’t normally think of, like encouraging the child to observe how others handle situations or simulate a scenario at home — these are quite practical.” Parents praised that these suggestions provide a clear foundation for future communication.
+The module also incorporated practical features that increased its utility.
+The “review and scoring” design was highlighted by four parents as a useful tool for evaluating the effectiveness of their interactions, as it allowed them to pinpoint areas needing improvement.
+P15 noted: “I think the scoring in the review session is useful.
+It helps me understand which parts of my communication need improvement.
+Everyone is busy, so this kind of tool is really practical.”.
+Additionally, motivational elements, such as the “badge achievement system”, were noted to enhance parental engagement and motivation.
+Despite these positive aspects, some concerns remained.
+A few parents pointed out that a single session was insufficient to assess real outcomes, and some found the suggestions too generic, which highlights the need for the AI to better adapt to specific parental contexts.
+7.3.3. Improving Parental Understanding and Grasp of Child Developmental Trajectories
+In terms of $C_{3}$, a lack of understanding of children’s emotional states, the Child Emotional Modeling module helped parents break through cognitive blind spots by leveraging data-driven insights to reveal their child’s emotional traits and behavioral patterns.
+Twelve parents reported that the system’s scientifically grounded feedback provided “new knowledge and dimensions”, helping them overcome cognitive blind spots regarding their child’s emotions and behavior.
+As P12 explained: “The system’s questions prompted me to think about aspects I hadn’t paid attention to before.
+The conclusions helped me understand my child more systematically and comprehensively.
+It was very helpful.” Thirteen parents further emphasized how AI generated “new angles” or “reference dimensions”, saying that the system helped them move beyond existing subjective impressions to enable more objective and reflective observations.
+This objective analysis was particularly valued in understanding the causes of children’s emotions, especially in labeling children’s emotions ( $S_{2}$).
+P6 explained: “AI can analyze the causes of a child’s emotions more objectively, whereas we as parents often bring in our own biases or emotions, making it hard to view the situation in a scientific and objective way.
+Parents also discussed the alignment between their views and the system’s analyses.
+Some parents reported that the system’s views were highly consistent with their own, which enhanced their trust.
+Others felt that when discrepancies arose, the system helped them see where their perceptions diverged—even serving as a valuable reference for resolving differences within the family.
+Beyond individual insights, the module also contextualized children’s behavior within broader developmental frameworks.
+Four parents mentioned that the AI’s use of large-scale data provided them with their first systematic understanding of “common behaviors at this age,” which helped them better assess their child’s traits.
+Additionally, several parents appreciated the forward-looking nature of the AI’s suggestions, which helped strengthen their overall understanding of their child’s developmental trajectory.
+Despite these advantages, several limitations were noted concerning the AI’s understanding and accuracy.
+Seven parents were concerned that the limited interaction time during the trial meant the generated profiles might not be comprehensive.
+Four parents noted that the system cannot verify the truthfulness of the information entered, which may lead to biased analysis.
+P1 said: “AI can only analyze based on what the child says, but children don’t always express their true feelings.
+Sometimes they even lie—for example, they might be very nervous but say they’re fine.
+This affects the AI’s judgment.” Two parents expressed discomfort with the scoring system and said they didn’t want to receive low scores from the system.
+8. DISCUSSION
+In our study, we designed PACEE from a parent-centered view.
+We based this approach on recognizing that the role of parents is essential in early childhood education and that the ability to interact with AI varies among children aged 3 to 6 years old.
+PACEE leverages multiple generative AI technologies (LLMs, text-to-image, text-to-speech) to address challenges for parents ( $C_{1}$- $C_{3}$).
+Our evaluation studies show that parents perceive PACEE to have successfully addressed these challenges, reducing their effort and boosting their satisfaction.
+Here, we further discuss the impact of generative AI on early childhood family education and the impact of PACEE on Gottman’s Five Steps of Emotion Coaching guideline (Gottman and DeClaire, 1997).
+We then propose broader design implications for parent-centered, AI-supported education systems.
+8.1. PACEE Enhanced Children’s Learning through Improving Parents’ Performance
+Prior work on AI-assisted child education has explored AI systems directly coaching the child (Seo et al., 2024; Tang et al., 2024), AI systems for intervening with parents (Entenberg et al., 2023), and systems that facilitate parent-child interaction simultaneously (Zhang et al., 2022; Theofanopoulou and Slovak, 2022).
+Similarly, we identify the need for generative AI support in our work.
+However, our context is different in that children aged 3 to 6 often face challenges in interacting independently with AI (Sun et al., 2024; Bush and Alibakhshi, 2025; Escobar-Planas et al., 2022; Thill et al., 2022), and early childhood emotion coaching requires more substantial support from parents (Theofanopoulou and Slovak, 2022; Zhu et al., 2025; Lee and Kim, 2023).
+Also, prior work has expressed concerns that AI alone may not be able to adequately respond to young children’s emotional needs (Kurian, 2025, 2024).
+Thus, we propose a parent-centered, AI-supported system for parent-child interaction.
+This parent-centered interaction pattern, where AI primarily assists parents rather than directly interacting with children, was preliminarily explored in therapeutic domains, such as for children with language delay (Hwang et al., 2014).
+We further extend this pattern to the domain of early-childhood education.
+In our work, PACEE enhances children’s learning outcomes mainly by improving parents’ performance.
+By positioning itself as a strategic partner rather than merely a tool, PACEE has fostered more effective communication between parents and their children, including higher interaction engagement and more in-depth emotional conversations.
+Consequently, child participants exhibited higher engagement levels and used more emotion-related words during the teaching sessions.
+However, the potential impact of generative AI directly interacting with children aged 3-6 in emotional education needs further exploration.
+In our study, we emphasized parents’ participation and informed parents that PACEE was not intended for the child to engage solely with the AI.
+This decision is due to findings from our formative study, where all five participating parents preferred to maintain their primary role in emotional guidance.
+Regarding the extent of AI directly interacting with the children, parents showed varying acceptance levels in the interviews.
+We also found that children’s AI literacy significantly varied based on parents’ attitudes.
+Therefore, the future design of emotion education systems needs to customize AI interaction levels based on parental acceptance and the child’s AI literacy.
+Also, longitudinal studies are required in order to assess the impact of varying degrees of AI interaction on children’s emotional development over time.
+Under our parent-centered interaction pattern, one challenge is to help parents better understand their children.
+To address this, we propose a framework for AI-parent collaborative modeling of children’s emotional profiles.
+In our framework, parents tend to contribute experiential knowledge of their children’s habits, while AI tends to evaluate developmental capabilities by analyzing parent-child interactions.
+This approach can also be extended to other areas of early childhood education, building upon prior work on personalized education systems.
+Prior systems either allow parents to manually set basic information and preferences (Zhang et al., 2022; Fang et al., 2024; Chen et al., 2025) or automatically assess children’s learning progress (Lee et al., 2024).
+By combining the strengths of parental expertise and AI-driven insights, our framework fosters a dynamic process of mutual learning.
+Parents and AI can collaboratively refine their understanding of the child’s needs, enabling more effective and personalized educational strategies.
+8.2. Adapting Gottman’s Five Steps of Emotion Coaching Guideline in the Era of Generative AI for Family Education
+While previous systems have implemented Gottman’s Five Steps of Emotion Coaching guideline (Gottman and DeClaire, 1997) in specific contexts (Seo et al., 2024; Ibrahim et al., 2025; Tang et al., 2024), our work demonstrates that a direct translation of this model is insufficient for the family education of children aged 3 to 6.
+We discuss this adaptation across three critical interaction phases, highlighting how generative AI addresses specific developmental and parental challenges to bridge the gap between theory and practical home education.
+Noticing, Recognizing, and Labeling Emotions.
+Gottman’s initial steps assume the child can verbally articulate the source of their emotion.
+However, prior literature indicates that preschoolers often lack the memory recall and verbal proficiency to reconstruct emotional events independently (Channell and Barth, 2013; Liwag and Stein, 1995).
+To address this, we emphasize active parental participation.
+Here, generative AI acts not only as a source of communication advice but also as a narrative scaffold.
+By generating virtual scenarios that align with the child’s preference for visual thinking, PACEE externalizes the child’s vague memories into concrete artifacts.
+This situates our design within the broader framework of mediated memory (Han, 2023), transforming the interaction from “interrogating the child” to “co-exploring a scenario.” However, limitations were identified: some parents noted that AI-generated scenarios sometimes lacked real-life relevance, reducing immersion.
+This suggests a trade-off in implementing Gottman’s guideline: how to balance generative inspiration with biographical accuracy to maximize the emotional bonding experience.
+Listening with Empathy and Validating Feelings.
+Gottman emphasizes that parents should show empathy and validate children’s feelings.
+Yet, our interviews reveal that parents often neglect this step in home education and find it challenging to perform it themselves.
+Consistent with prior work (Yoo et al., 2019), PACEE encourages parents to adopt the child’s perspective during moments of conflict.
+Moreover, PACEE innovatively introduces a Child Emotional Modeling module, utilizing text-based analysis of parent-child conversation to provide parents with an objective “second opinion” on the child’s state.
+Our evaluation shows that parents prefer to extract insights from specific life events, while the AI focuses on evaluating developmental capabilities.
+This result provides an empirical framework of parent-AI collaborative understanding: Generative AI compensates for the parent’s subjective blind spots with developmental metrics, while the parent grounds the AI’s analysis in the real-world family context.
+Setting Limits and Problem Solving.
+Research suggests that children aged 3 to 6 respond well to positive reinforcement mechanisms (Hardy and McLeod, 2020; Sigler and Aamidor, 2005).
+So we adapted Gottman’s guideline by integrating the PERMA model (Seligman, 2010), prioritizing the discovery of positive emotions and solutions.
+In this way, the parent-child interaction focus shifts from “correction” to “positive resource building.” However, our study highlights a critical tension in applying generative AI to this step: the risk of dependency and rigidity.
+Parents expressed concerns that rigid, AI-imposed goals might undermine the collaborative nature of parenting, and that children might over-rely on AI for solutions.
+Therefore, we suggest future adaptations of Gottman’s guidelines with generative AI prioritize process-oriented scaffolding (e.g., stimulating child thinking) over solution-oriented outputs (e.g., resolving the immediate conflict).
+In summary, our work refines Gottman’s dyadic “parent-child” interaction framework into a triadic “AI-Parent-Child” framework.
+We demonstrate that generative AI plays a unique role in bridging the “theory-practice gap”—specifically by mitigating the developmental limitations of children aged 3 to 6 and the communication barriers faced by parents.
+On the other hand, we maintain that parental agency remains central; the parent needs to serve as the primary emotional anchor and final content gatekeeper, while the AI empowers the parent.
+8.3. Design Considerations for Parent-centered, AI-supported Education Systems
+Parent interviews also revealed several challenges that need to be addressed.
+Combining our findings with prior work, we propose design considerations for future parent-centered, AI-supported education systems.
+First, customize AI’s aiding modules based on parents’ profiles, in addition to their children’s.
+Some recent work has integrated children’s profiles or preferences to facilitate a more personalized coaching experience (Clabaugh et al., 2019; Tang et al., 2024; Zhang et al., 2022).
+This study supplements prior research and demonstrates the necessity of customizing AI systems based on parental profiles.
+During our interviews, we found that parents expressed significantly varied opinions about the “ideal support” from PACEE.
+Concerning content difficulty, some experienced parents criticized that AI’s guidance overlapped heavily with their own approaches, while a few parents mentioned that some suggestions were too complex to understand.
+Concerning the frequency of suggestions, some parents felt that frequently checking the AI’s advice could slightly disrupt the conversational flow.
+In contrast, others believed that PACEE could offer more real-time interactions.
+These findings highlight that parent-centered educational systems should accommodate such differences in parents’ needs.
+For example, parents’ preferences regarding the appropriateness of early-childhood digital play, which may influence the quality of AI-assisted activities at home (Isikoglu Erdogan et al., 2019), should be collected beforehand.
+Furthermore, various parental inputs can be captured during parent-child interactions, besides audio data from conversations.
+These inputs may include whether parental feedback during interactions is positive or negative (Robinson and Eyberg, 1981), behavioral differences between fathers and mothers (Cutler and Palkovitz, 2020), and parents’ emotional regulation abilities (Bariola et al., 2011).
+By incorporating diverse data sources, the system can better tailor its functions to the specific needs of each family.
+Second, avoid excessive reliance on AI capabilities and promote transparent and negotiable collaboration.
+In the evaluation studies, parents expressed worries that the system cannot verify the truthfulness of the information entered, which may lead to biased analysis.
+They also noted that AI’s interpretation of a child’s emotions may contradict the parents’ understanding.
+These concerns align with the communication aspect of cooperative AI (Dafoe et al., 2020).
+By explicitly sharing information, AI agents and parents can gain insights into each other’s behaviors and intentions more effectively.
+Here, we recommend leveraging research findings from explainable AI to enhance transparency in collaboration (Liao et al., 2020).
+Parents need to understand the rationale behind AI-generated assessments or suggestions, thus enhancing trust in AI’s decisions (Kim, 2024).
+In designing future systems, it is crucial to consider how to balance the parents’ cognitive burden of processing information with the goal of enhancing trust by providing a comprehensive explanation, which may be related to parents’ diverse decision-making styles (van Arum et al., 2025).
+Considering the potential conflicts between parents and AI systems, we recommend the introduction of negotiation tools (Crandall et al., 2018; Jennings et al., 2001), as the current context lacks a formal agreement protocol.
+In the context of emotion education, it is essential to develop interactive negotiation mechanisms to address varying interpretations of children’s emotional states.
+Furthermore, AI can serve as a mediator to reconcile differing perspectives on children’s emotions among family members, a role that some parents expressed interest in during the interviews.
+Third, respect parents’ boundaries regarding AI intervention and provide them with sufficient autonomy.
+Prior work has explored the area of parents’ boundaries, such as parental concerns with AI-driven technologies in their child’s healthcare (Sisk et al., 2020), showing that parents value shared decision-making but may resist interventions that undermine their sense of control.
+We supplement prior research and expand this area to include both emotion-sensitive contexts and early childhood education domains.
+In our context, children’s emotional education may involve sensitive family matters.
+In the interviews, parents expressed that they do not want AI to discuss family privacy or parents’ professional information.
+Future work could look into value-sensitive AI design frameworks (Umbrello and Van de Poel, 2021) and provide information sensitivity detection techniques that align with diverse parental values.
+Moreover, emotional education may involve delicate feelings or negative topics that could affect children’s mood.
+To address this issue, PACEE incorporated the PERMA positive psychology model into its design to reduce the emphasis on negative content.
+Future work could investigate how to further mitigate emotional overwhelm that may be perceived as harmful by parents.
+For example, adaptive AI systems could use real-time sentiment analysis to detect emotional intensity and compare it with predefined emotional boundaries set by parents.
+As for education for children aged 3 to 6, parents consistently viewed themselves as the primary decision-makers in their children’s education.
+In the interviews, parents expressed concerns about the possibility of AI imposing rigid goals on children and compelling parents to follow prescribed steps.
+Future systems should prioritize bidirectional discussions with parents, rather than simply providing prescriptions.
+Furthermore, they should ensure parents maintain authority over critical decision-making in parent-AI collaboration.
+8.4. Limitations and Future Work
+This study has several limitations.
+First, the participants’ cultural background was limited to Chinese parents and children, which may constrain the cultural generalizability of our findings.
+Second, most of the parent participants were mothers.
+This reflects the broader caregiving patterns noted in prior research (McBride and Mills, 1993; Schoppe-Sullivan et al., 2013; Hernawati and Herawati, 2020), and likely represents the primary user group for a system designed to support caregivers of 3-6-year-olds in the real world.
+However, this imbalance may limit our ability to capture the full range of parental perspectives.
+Future work should recruit across more diverse cultural and gender contexts, and systematically examine how fathers and mothers may differ in their collaboration with AI to support children.
+Third, a limitation of the current study design is the inability to fully isolate the foundation model variable.
+While our study’s primary focus was on the interaction framework rather than raw model output, as supported by our findings and clarified in Section 6.2, we acknowledge that this does not entirely eliminate the potential bias inherent in the two models.
+Therefore, future work could employ a 2x2 factorial design — (Interface: Baseline vs.
+PACEE) x (Model: Doubao vs.
+GPT-4o) — to further isolate the effect of the model.
+Currently, PACEE relies on text-based conversation analysis to assist parents in understanding their child’s emotional states.
+Future developments could, with parental consent, incorporate additional modalities, such as collecting children’s gestures and facial expressions.
+However, regardless of the modality used, it is important to note that PACEE’s role is to augment parents’ understanding, not replace the parents.
+Ultimately, parents remain the final decision-makers in deciding whether to adopt the system’s suggestions and how to interpret their child’s emotional states.
+Finally, our study was conducted in lab-based settings using a touchscreen laptop, which may not fully reflect the long-term usage experience at home.
+In addition, first-time exposure to PACEE’s interactive features may have temporarily increased participants’ engagement, reflecting a “novelty effect” commonly faced by HCI research (Strömel et al., 2024; Soler-Dominguez et al., 2024; Davis et al., 2025).
+To address this issue, we provided parents with a practice session to familiarize them with the procedure.
+While our analyses did not reveal significant differences in comparative outcomes between early and later interactions, future work should explore how such novelty influences parents and children during longer interactions.
+Future work should also implement longitudinal in-home deployments to better understand how parent-AI collaborative emotion education unfolds in natural family environments.
+9. CONCLUSION
+We introduced PACEE, an assistant designed to support family emotion education through parent-AI collaborative emotional profiling and guidance.
+Drawing on interviews with kindergarten teachers and parents, PACEE leverages generative AI to address the key challenges identified in family emotion education.
+Our evaluation study with 16 pairs of parents and children demonstrated that PACEE enhanced parent-child engagement, fostering deeper emotional exchanges.
+Furthermore, PACEE significantly alleviated parental challenges, resulting in greater satisfaction and less effort from parents.
+We hope our work will advance emotion coaching theory in both family and AI-assisted contexts.
+Additionally, we hope to inspire the development of parent-centered, AI-supported education systems that empower families to promote children’s well-being.